### v2

# Early Classification of Time Series: Taxonomy of Methods and Extensive Benchmark

Abstract
In many situations, the measurements of a studied phenomenon are provided sequentially, and the prediction of its class needs to be made as early as possible so as not to incur too high a time penalty, but not too early and risk paying the cost of misclassification. This problem has been particularly studied in the case of time series, and is known as Early Classification of Time Series (ECTS). Although it has been the subject of a growing body of literature, there is still a lack of a systematic, shared evaluation protocol to compare the relative merits of the various existing methods. This document begins by situating these methods within a principle-based taxonomy. It defines dimensions for organizing their evaluation, and then reports the results of a very extensive set of experiments along these dimensions involving nine state-of-the art ECTS algorithms. In addition, these and other experiments can be carried out using an open-source library in which most of the existing ECTS algorithms have been implemented (see https://github.com/ML-EDM/ml_edm).
1 Introduction
In hospital emergency rooms (?), in the control rooms of national or international power grids (?), in government councils assessing critical situations, there is a time pressure to make early decisions. On the one hand, the longer a decision is delayed, the lower the risk of making the wrong decision, as knowledge of the problem increases with time. On the other hand, late decisions are generally more costly, if only because early decisions allow one to be better prepared. For example, a cyber-attack that is not detected quickly enough gives hackers time to exploit the security flaw found.
A number of applications involve making decisions that optimizes a trade-off between accuracy of the prediction and its earliness. The problem is that favoring one usually works against the other. Greater accuracy comes at the price of waiting for more data.
Such a compromise between the Earliness and the Accuracy of decisions has been particularly studied in the field of Early Classification of Time Series (ECTS) (?), and introduced by ? (?). ECTS consists in finding the optimal time to trigger the class prediction of an input time series observed over time. Successive measurements provide more and more information about the incoming time series, and ECTS algorithms aim to optimize online the trade-off between two conflicting objectives, namely, the earliness and accuracy of class predictions. More formally, we have the following problem.
Problem statement:
In the ECTS problem, measurements of an input time series are observed over time. At time $t$, the incomplete time series ${\mathbf{x}}_{t}\,=\,\langle{x_{1}},\ldots,{x_{t}}\rangle$ is available where ${x_{i}}_{(1\leq i\leq t)}$ denotes the time indexed measurements. These measurements can be single or multi-valued. The input time series belongs to an unknown class $y\in{\cal Y}$. The task is to make a prediction $\hat{y}\in{\cal Y}$ about the class of the incoming time series, at a time $\hat{t}\in[1,T]$ before the deadline $T$ which corresponds to the length of the full time series. A misclassification cost is incurred when a prediction is made, denoted by $\mathrm{C}_{m}(\hat{y}|y)$. Furthermore, there exists a delay cost $\mathrm{C}_{d}(\hat{t})$ that expresses the time pressure and encourages early decisions (defined in Section 2.2). The choice of the best trigger time should optimize a compromise between these costs that move in opposite direction. The choice of the best triggering time must optimize a compromise between the two costs, which are moving in opposite directions.
To the best of our knowledge, ? (?) are the earliest explicitly mentioning “classification when only part of the series are presented to the classifier”. Since then, several researchers have continued their efforts in this direction and have published a large number of research articles. A recent and extensive review of the ETCS approaches can be found in the paper written by ? (?), including the applications that motivated the researchers to work in this area, and covering about fifty relevant papers selected from the $213$ papers found by search engines at the time of this writing.
As pointed out by ? (?), the ECTS problem is a special case of optimal stopping (?, ?), where the decision to be made concerns both: (i) when to stop receiving new measurements in order to (ii) predict the class of the incoming time series.
In the same paper, the ECTS problem has been extended into a more general one, which consists in optimizing the decision times of Machine Learning (ML) models in a wide range of settings where data is collected over time. The authors proposed a set of open research questions to the community, in order to widen the range of applications that are amenable to the ECTS framework (i.e. dealing with other learning tasks, other types of data, other application contexts etc.).
However, despite the growing interest in ECTS over the last twenty years, there still remains a need for a shared taxonomy of approaches and an agreed well-grounded evaluation methodology. Here, in particular, we list limits that hamper a fair comparison of ECTS methods and algorithms:
  1. 1.
    Costs taken into account for evaluating the performance of the proposed method are not always clearly stated. It seems natural to distinguish between the misclassification costs $\mathrm{C}_{m}(\hat{y}|y)$, and the delay cost $\mathrm{C}_{d}(t)$, and to add them in order to define the cost of making a decision at time $t$. More generally, the delay cost may depend on the true class $y$ and the predicted one $\hat{y}$, and a single cost function $\mathrm{C}_{m}(\hat{y}|y,t)$ integrating misclassification and delay costs should then be used. For the sake of clarity, we keep the simple notation which distinguishes both cost functions in the rest of this paper. But in all cases, it is essential to state the framework used and the associated evaluation metric.
  2. 2.
    The performance of the proposed methods should be evaluated against a range of possible types of cost functions. It is usual to evaluate “by default” the methods using a $\ell_{0-1}$ loss function that penalizes a wrong classification by a unity cost, and to consider a linear delay cost function: $\mathrm{C}_{d}(t)=\lambda\,t$, for a value $\lambda>0$. However, lots of applications rather involve unbalanced miss-classification costs, and possibly also non linear delay costs. This is the case, for instance, in maintenance applications where wrongly not recognizing a critical situation is much more costly than wrongly predicting a problem and taking steps to fix it, and where delay cost may rise as a an exponential function of time: $\mathrm{C}_{d}(t)=\lambda\,e^{t}$ It is therefore quite important to assess the adaptability of the methods to various representative problem settings.
  3. 3.
    The contributions of the various components of a ECTS algorithm should be clearly delineated. The predominant approach to ECTS is to have a decision component which is in charge of evaluating the best moment to make the prediction about the class of the incoming times series, and a classifier one which makes the prediction itself. In order to fairly compare the triggering methods, which are at the heart of ECTS, the classifier used should be the same. We call these methods “separable methods”.
    An alternative approach relies on having a system that classifies the incoming time series $\mathbf{x}_{t}$ at each time with a prediction $\hat{y}$ in the set {‘postpone decision’, $y_{1},\ldots,y_{N}$} where $N$ is the number of classes. Therefore, within this approach, a single system decides either to wait at least one more time step, or to predict a class and stops the process. In this case, no distinction can be made between a decision component and a classifier one, and the whole system is evaluated as such. In the spirit of deep neural networks, we call this type of methods “end-to-end” to underline the fact that a single learning system is in charge of all operations, here decision and classification. Of course, this precludes a comparison involving only the choice of the decision component with other methods.
  4. 4.
    As with other supervised learning tasks, performance should be compared with that of “baseline” algorithms. In the case of ECTS tasks, two naive baselines are: (1) make a prediction as soon as it is allowed, and (2) make a prediction at $T$, after the entire time series has been observed. In our experiments reported in Section 4, we have added a third baseline, less simple than the two afore-mentioned ones, but still too obvious so that, to our knowledge, it has never been published as an original method. This is a confidence-based method where a decision is triggered as soon as the confidence for the likeliest class given $\mathbf{x}_{t}$ is greater than a threshold. (Formally, let $\hat{y}=\operatorname*{arg\,max}_{y\in\mathcal{Y}}p(y|\mathbf{x}_{t})$, then a prediction is made (i.e. $\hat{y}$) as soon as $p(\hat{y}|\mathbf{x}_{t})\geq\varepsilon$, for some threshold $\varepsilon\in[0,1]$.)
  5. 5.
    Precautions should be taken when using datasets of training time series to ensure that no bias enters unwillingly into the training and evaluation process. A case in point, concerns the normalization often used in time series datasets. ? (?) have reported that 71% of the reference time series classification datasets used to evaluate ECTS methods are made up of z-normalized time series, i.e. with measurements independently modified on each complete series to obtain a mean of zero and a standard deviation equal to 1. Clearly, this setting is not applicable in practice, as z-normalization would require knowledge of the entire incoming time series. In a research context, previous work has used such training sets to test the proposed algorithms. As ? (?, ?) note, this preprocessing is irreversible and can generate a problem for ECTS by introducing a temporal leakage of information. In order to assess its impact, we report in Section B.5 of Appendix B a comparison of results for z-normalized and non-normalized time series.
Up until now, it has been difficult to conduct fair comparisons between competing methods. Often, published performances are based on choices concerning data sets, the precise performance measure used, hyperparameter values and evaluation protocols (e.g., the split between training and test sets) that are not entirely explicit or, in any case, are difficult to reproduce. This is why we have recoded all the methods, specified a shared evaluation protocol with variants that can be employed by everyone, and searched for a collection of data sets that can be widely used to test and compare new as well as existing methods. We hope this will be a useful resource for the scientific community working in this field.
As this community shows a growing interest in the ECTS problem, granted by the increasing number of applications that fall in its range, it is timely (1) to propose a framework into which to cast the various approaches and thus indicate avenues for future research, and (2) a well-grounded evaluation methodology. Specifically, this paper makes the following contributions:
  • A taxonomy is proposed in Section 2, classifying approaches in the literature according to their design choices.
  • Extensive experiments have been performed, meeting the above-mentioned shortcomings. (1) The experimental protocol in Section 4.1 explicitly defines the costs used during training and evaluation, and varies the balance between misclassification and delay costs by using a large range of cost values. (2) Experiments are performed repeatedly for several types of cost function, i.e. balanced or unbalanced misclassification cost, and linear or exponential delay cost (see Sections 4.2 and 4.3) and many intermediate results are available in the supplementary materials. (3) Ablation and substitution studies are conducted in Section 4.4 with the aim of evaluating the impact of methodological choices, such as the choice of classifier, its calibration, or even z-normalization of training time series. (4) The experiments include three baseline approaches, rarely considered in the literature, which often give surprisingly good results. (5) In addition to the reference data used in the ECTS field, a collection of some thirty non-z-normalized datasets is proposed and provided to the community.
  • An open source library is being made available [^1] https://github.com/ML-EDM/ml_edm in order to enable reproducible experiments, as well as facilitate the scientific community’s development of future approaches. Particular care has been taken to ensure the quality of the code, so that this library may be used to develop real-life applications.
The rest of this paper is organised as follows : The section 2 proposes and describes a new ECTS taxonomy, the different choices to be made in a well-founded way when designing an ECTS method and a set of four questions that need to be answered to make these choices.
Section 3 presents a comprehensive view of the ECTS field, along the suggested taxonomy and the four questions raised in Section 2. In Section 4 we present the pipeline developed in order to realize extensive experimentation and we report the main results obtained for different cost setting. This benchmark is supported by a library released as Open Source for dissemination and used in the ECTS research community. Finally Section 5 concludes this paper. Appendix A lists the datasets used for the experiments, and complementary results are provided in Appendix B.
2 Organizing the ECTS approaches: a taxonomy
The aim of this section is to outline in a principled way the various choices that need to be made when designing one ECTS method.
General form of an ECTS model
An ECTS approach aims at optimizing a trade-off between accuracy and earliness of the prediction, and thus must be evaluated on this ground. The correctness of the prediction is measured by the misclassification cost $\mathrm{C}_{m}(\hat{y}|y)$ where $\hat{y}$ is the prediction and $y$ is the true class. The time pressure is sanctioned by a delay cost $\mathrm{C}_{d}(t)$ that is assumed to be positive and, in most applications, an increasing function of time. We thus consider:
  • $\mathrm{C}_{m}(\hat{y}|y):{\cal Y}\times{\cal Y}\rightarrow\mathbb{R}$, that corresponds to the misclassification cost of predicting $\hat{y}$ when the true class is $y$.
  • $\mathrm{C}_{d}(t):\mathbb{R}^{+}\rightarrow\mathbb{R}$, the delay cost that, usually, is a non-decreasing function over time.
An ECTS function involves a predictor $\hat{y}(\mathbf{x}_{t})$, which predicts the class of an input time series $\mathbf{x}_{t}$ for any $t\in[1,T]$. The cost incurred when a prediction has been triggered at time $t$ is given by a loss function $\mathcal{L}(\hat{y}({\mathbf{x}}_{t}),y,t)=\mathrm{C}_{m}(\hat{y}({\mathbf{x}}_{t})|y)+\mathrm{C}_{d}(t)$. The best decision time $t^{*}$ is given by:
$t^{\star}=\operatorname*{arg\,min}_{t\in[1,T]}\mathcal{L}(\hat{y}({\mathbf{x}}_{t}),t,y).$ (1)
Let $s^{\star}\in\mathcal{S}$ an optimal ECTS function belonging to a class of functions $\mathcal{S}$, whose output at time $t$ when receiving $\mathbf{x}_{t}$ is:
$s^{\star}({\mathbf{x}}_{t})=\left\{\begin{array}[]{ll}\emptyset&\mbox{if extrameasures are queried;}\\ y^{\star}=\hat{y}(x_{t^{\star}})&\mbox{when prediction is triggered at $t=t^{\star}$;}\end{array}\right.$ (2)
ECTS is however an online optimization problem, where at each time step $t$ a function $s({\mathbf{x}}_{t})$ must decide to make a prediction or not. Equation 1 is thus no longer operational since it requires a complete knowledge of the time series. In practice, the function $s({\mathbf{x}}_{t})$ triggers a decision at $\hat{t}$, based on a partial description ${\mathbf{x}}_{\hat{t}}$ of the incoming time series ${\mathbf{x}}_{T}$ (with $t\leq T$). The goal of an ECTS system is to chose a triggering time $\hat{t}$ as close as possible to the optimal one $t^{*}$, at least in terms of cost, minimizing $\mathcal{L}(\hat{y}({\mathbf{x}}_{\hat{t}}),\hat{t},y)-\mathcal{L}(\hat{y}({\mathbf{x}}_{t^{\star}}),t^{\star},y)$ as much as possible.
From a machine learning point of view, the goal is to find a function $s\in{\cal S}$ that best optimizes the loss function $\mathcal{L}$, minimizing the true risk over all time series distributed according to the distribution222 Notice that the notation $\mathcal{X}$ is an abuse that we use use to simplify our purpose. In all mathematical rigor, the measurements observed successively constitute a family of time-indexed random variables $\mathbf{x}=(\mathbf{x}_{t})_{t\in[1,T]}$. This stochastic process $\mathbf{x}$ is not generated as commonly by a distribution, but by a filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[1,T]}$ which is defined as a collection of nested $\sigma$-algebras (?) allowing to consider time dependencies. Therefore, the distribution $\mathcal{X}$ should also be re-written as a filtration. $\mathbb{P}_{\mathcal{X}}$ that governs the time series in the application:
$\small\normalsize\operatorname*{arg\,min}_{s\in\mathcal{S}}\operatornamewithlimits{\mathbb{E}_{\>\mathbf{x}\sim\mathbb{P}_{\mathcal{X}}}^{\>}}\left[\mathcal{L}(\hat{y}({\mathbf{x}}_{\hat{t}}),\hat{t},y)\right]$ (3)
The questions then are:
  1. 1.
    Which form can take the function $s(\cdot)$? We will distinguish end-to-end architecture from separable ones.
  2. 2.
    How the criterion to be optimized accounts for the trade-off between accuracy and earliness? We will see that the costs implied, about misclassification and delay, are variously explicit in the existing methods.
  3. 3.
    How the when question of the stopping problem can be approached? This will lead us to distinguish between cost-informed and cost-uninformed ones on one hand, and between anticipation-based methods versus myopic ones, on the other hand.
  4. 4.
    How the prediction problem itself can be solved given that $\mathbf{x}_{t}$ belongs to a different input spaces at each time step $t$?
This set of questions and possible choices for their solution are illustrated in Figure 1. We turn successively to each one in the following.
Figure 1: Proposed ECTS taxonomy
2.1 The different forms of the function $s(\cdot)$
An ECTS function must solve both the question of (i) when to stop receiving new measurements and decide to make a prediction and (ii) how to make the prediction about the class of the incoming time series $\mathbf{x}_{t}$.
In the separable approach, these questions are solved using two separate components. The classification one deals with making a prediction: $\mathbf{x}_{t}\mapsto\hat{y}$, while the trigger function decides when to predict. Within this perspective, the classification component is learned independently of the trigger one, while the latter uses the results of the classification component in order to trigger a decision. A simple triggering strategy is to decide it is time to make a prediction as soon as the classification component is sufficiently sure of its prediction. We formalize separable approaches by: $s(\mathbf{x}_{t})=(g\circ h)(\mathbf{x}_{t})$ where $g$ is the decision or trigger function, and $h$ is the prediction function.
In the end-to-end approaches, a single component decides when to make a prediction and what that prediction is. Thus, the function $s$, defined in Equation 2, is responsible both for choosing the time $\hat{t}$ for making the prediction, and for the prediction itself $\hat{y}$.
The question that naturally arises is which type of architecture (i.e. end-to-end or separable) performs best. On the one hand, in separable approaches, the classification component is trained independently of the triggering one, which can be detrimental. On the other hand, separating the ECTS problem into two inherently simpler sub-problems could be an advantage. In this paper, we do not delve any further into this question, which we leave for future work.
2.2 Choice of the optimizing criterion during training
This section covers optimization criteria existing in the literature and used to train ECTS functions. In the following, these criteria are listed by level of cost awareness, and we differentiate between the following situations:
  1. 1.
    The first one, we call cost-informed at training time.
    $\mathbb{P}_{\mathcal{X}}$ being unknown, instead of using Equation 3 describing the true risk, one tries to minimize the empirical risk, also called average cost in the ECTS literature, for a training set of $M$ time series:
    $AvgCost\;=\frac{1}{M}\sum_{i=1}^{M}\mathcal{L}(\hat{y}_{i},y_{i},\hat{t})\;=\;\frac{1}{M}\sum_{i=1}^{M}\mathrm{C}_{m}(\hat{y}_{i}|y_{i})+\mathrm{C}_{d}(\hat{t}_{i})$ (4)
    AvgCost is the most appropriate criterion to both train and evaluate ECTS approches. The following presents proxy measures of AvgCost.
  2. 2.
    The second situation is a sub-case of cost-informed, and can be qualified as cost-proxy at training time. There, while the accuracy and earliness of prediction are taken into account, the optimization criterion combines them in a proxy which is not the AvgCost. Classical proxys include:
    • The Harmonic Mean (see ?):
      $HM=\frac{2\times Accuracy\times(1-Earliness)}{Accuracy+(1-Earliness)}$ (5)
      with:
      $Accuracy$ $=\frac{1}{M}\sum_{i=1}^{M}\mathbb{1}(\hat{y}_{i}=y_{i})$ (6)
      $Earliness$ $=\frac{1}{M\times T}\sum_{i=1}^{M}\hat{t}_{i}$ (7)
      where $\hat{t}_{i}$ is the time that the ECTS function decides to make prediction for the time series $\mathbf{x}^{i}$.
    • The CF (i.e. Cost Function) criterion with $0\leq\alpha\leq 1$ (see ?, ?, ?):
      $CF=\alpha\times(1-Accuracy)+(1-\alpha)\times Earliness$ (8)
      When the cost of misclassification $\mathrm{C}_{m}(\hat{y},y)=\mathbb{1}(\hat{y}=y)$, and the delay cost is $\mathrm{C}_{d}(t)=\hat{t}/T$ and $\alpha=0.5$, CF becomes a particular case of AvgCost.
    The choice of using costs or approximating them by a proxy is a technical issue, which may, for example, be relevant to making the loss function differentiable by approximation. For this reason, in the remainder of this paper, cost-informed/cost-proxys at training time situations are grouped together under the term cost-informed at training time, this difference not being essential.
  3. 3.
    The third situation is qualified as cost-uninformed-train. This is the case of methods that only set the threshold on the confidence of the prediction in order to trigger the prediction, regardless of the costs incurred on the training set. Another example is methods that use rigid rules, such as “decide as early as the first measurement is available”.
It is a question whether any of these approaches fare better than the others. To measure this, we must use the Average Cost (see Equation 4) measured on a test set, which represents the ground truth of the ECTS problem. It should be noted that many of the proposed methods have been evaluated on the basis of other criteria in the literature, with the result of not allowing a rigorous comparison between them. We will come to this problem in Section 4.
The rest of this section is specific to separable ECTS approaches, which represent a large part of the literature.
2.3 Information used by the trigger function during inference
The trigger function can draw on different types of information. In the simplest case, it can decide irrespectively of the incoming time series $\mathbf{x}_{t}$. This is the case, for example, of the rule that would say: “wait until half of the measurements are available, then make a prediction”. The corresponding trigger function can be said to be blind.
Apart from this extreme case, it is interesting to distinguish two dimensions. First, the trigger function may or may not take the misclassification and delay costs into account. Second, it can also make its decision at each time step solely on the basis of past observations, or it can anticipate possible futures to help its decision.
For instance, confidence-based methods (seen Section 2.3.1) do not explicitly take costs into account in the triggering decision.
2.3.1 Confidence-based approaches
Confidence-based approaches are widely used in the literature. The simplest trigger model of this kind consists in monitoring a quantity related to the confidence of the prediction over time and triggering class prediction as soon as a threshold value is exceeded. The confidence metric monitored can take different forms. For example, a baseline approach, referred to as Proba Threshold [^3] Baseline implemented in the aeon (?) library : https://urlz.fr/qmWl in the remainder of this paper, involves monitoring $\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})$ the highest conditional probability estimated by the classifier. This baseline example is qualified as instant-based method, since it takes as input only the last confidence score available at time $t$. Another type of approaches, qualified as sequence-based, monitors the entire sequence of past confidence scores, and triggering a prediction is made conditionally on a particular property of this sequence. Accordingly, trigger functions can either take as input a scalar value, e.g. $g(\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t}))$, in the case of instant based approaches, or a sequence of scalar values, e.g. $g(\{\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{\tau})\}_{1\leq\tau\leq t})$, in the case of sequence-based approaches (see Section 3.1).
2.3.2 cost-informed at testing time
Given that an ECTS approach will ultimately be evaluated on the average cost of using it (see Equation 4), it seems natural to exploit the cost values at testing time, in order to trigger predictions at optimal moments. Methods such as Economy (?) and 2step/NoCluster (?) do that. They can thus be qualified as “cost-informed at testing time”.
Other approaches use instead trigger functions that, once learned, do not take into account the cost at testing time, but rely on other measures such as, for instance, the confidence of the prediction. This is the case of the SR approach (?). Therefore, these approaches can be qualified as “cost-uninformed at testing time”.
Notice that some approaches are cost-informed during training but not during inference. This is the case with the SR approach, which is cost-informed at training time since it uses costs to optimize its parameters, and cost-uninformed at testing time since the resulting trigger function does not use cost values during inference. Table 1 shows these two different properties for each approach in the literature.
2.3.3 Blind, Myopic and Anticipation-based decisions
Some separable approaches consider the output of the classifier $h$ at time $t$ to decide whether this is the right time to make a prediction. For instance, stopping the process when the confidence in the classification $h_{t}(\mathbf{x}_{t})$ is above some threshold, or when the difference of confidence between the best two predictions exceeds some value. These methods can be described as myopic since they only look at the current time step $t$, without trying to guess the future.
But there is a another possibility. As was first noted by ? (?), the ECTS problem can be cast as a LUPI (Learning Using Privileged Information) problem (?).
In this scenario, the learner can benefit at the training time of privileged information that will not be available at test time. Formally, the training set can be expressed as $\mathcal{T}=\{(\mathbf{x}_{i},\mathbf{x}_{i}^{\star},y_{i})\}$, where $\mathbf{x}_{i}$ is what is observable and $\mathbf{x}_{i}^{\star}$ is some additional information not available when the prediction must be made. This is exactly what happens in the ECTS problem. Whereas at test time, only $\mathbf{x}_{t}$ is available, during training the complete time series are known. This brings the possibility to learn what are the likely future of an incoming time series $\mathbf{x}_{t}$ provided it comes from the same distribution. Hence, it becomes also possible to guess the cost to be optimized for all future time steps, and therefore to wait until the moment seems the best. This type of approach can be said anticipation-based (also called non-myopic in the literature). Because more information from the training set is exploited, it can be expected that these methods outperform myopic and blind ones.
Is this confirmed by experience? Are there situations where the advantage is significant? Our experiments in Section 4 provide answers to these questions.
2.4 Choice of the classification component
One source of difficulty when devising an ECTS method in the separable setting is that inputs differ from one time step to another. The number of measurements, and hence the input dimension, varies. Two approaches have been used to deal with the problem.
  1. 1.
    A set of classifiers $\{h_{t}\}_{t\in[1,T]}$ is learned, each dedicated to a given time step $t$, and thus a given input dimension. In practice, authors often chose a limited subset of timestamps, usually a set of twenty (one measurement every 5% of the length of the time series), to restrict the number of classifiers to learn and therefore the associated computational cost.
  2. 2.
    A single classifier $h$ is used for all possible incoming time series $\mathbf{x}_{t}$. One way of doing this is to “project” an input $\mathbf{x}_{t}$ of dimension $t\times d$, if $d$ is the dimension of an observation at time $t$ (i.e. multi-valued time series), into a fixed dimensional vector whatever $t$ and $d$. This may simply be the mean value and standard deviation of the available measurements (multiplied by the dimension $d$) or the result of a more sophisticated feature engineering as tested by ? (?). Deep learning architectures can also be used to learn an encoding of the time series in an intermediate layer. For instance, ? (?) use a CNN architecture, and ? (?) a FCN one. ? (?) show that using deep neural architectures often perform well for time series classification.
Both approaches have their own limitations. On the one hand, using a set of classifiers, each independently dedicated to a time step, does not exploit information sharing. On the other hand, using a single classifier seems to be a more difficult task, as the representation of $\mathbf{x}_{t}$ can be different at times $t$ and $t+1$ and all further time steps which can lead to additional difficulty for the classifier while moreover requiring a more demanding feature engineering step. Therefore, here also, it is interesting to measure experimentally whether one dominates the other. This will be the subject of future work.
3 State of the art
In this section, we position various existing proposed methods along the dimensions of the taxonomy presented in Section 2 (see Table 1). Then, we examine the methods and how they exemplify solutions to the general questions raised in the taxonomy. Thus, while each method combines solutions to all questions, in the following, we underline how each one brings an original solution to one specific problem. For instance, the Economy approach is separable, anticipation-based, and cost-informed at testing time. However, we’re focusing here on the anticipation-based dimension, since this is one of the first methods to have emphasized it and introduced an original solution for this aspect.
References Classifier(s) (collection ✓) End2end Confidence Anticipation Cost informed
EDSC (?) Shapelet
ECTS’ (?) 1NN
Reject (?) SVM
RelClass (?) QDA, Linear SVM
[…]
2step/NoCluster (?) Linear SVM (✓) train & test
ECDIRE (?) Gausian Process (✓)
Stopping Rule (?) Gaussian Process (✓) train
[…]
ECEC (?) WEASEL (✓) train
DDQN (?) MLP train
TEASER (?) WEASEL (✓) train
[…]
DETSCNet (?) TCN train
CALIMERA (?) MiniROCKET (✓) train
ELECTS (?) LSTM train
[…]
EarlyStop-RL (?) MLP train
Table 1: Table of published methods for the ECTS problem with their properties along dimensions underlined in the taxonomy.
3.1 Confidence-based approaches
Most ECTS methods to date are separable, confidence-based, cost-informed at training time, and are not anticipation-based. They implement separately the prediction and the triggering components, they learn them using the costs, hence they are cost-informed at training time, but they decide to trigger a decision based on the information available at the current time step $t$ without trying to anticipate the likely future, and they base their decision upon the confidence of the predictions made by the classifier.
There exist two families of confidence-based approaches. In the first one, only the last time step is considered, a score based on confidence estimations is monitored at each time step and a class prediction is triggered as soon as a threshold on this score is exceeded. By contrast, in the second, a sequence of estimated scores is monitored, and the condition to trigger a decision depends upon some property of this sequence.
3.1.1 Instant-based decision criterion
$\bullet$ One basic method is to monitor $\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})$, the highest conditional probability estimated by the classifier, which is a simple measure of classifier confidence over time. As soon as it exceeds a value, which is an hyper parameter of the method, a prediction is made. We call this method Proba Threshold and use it as baseline for comparison later in our experiments.
$\bullet$ The Reject method (?) uses ensemble consensus as a confidence measure. For each time step, first (i), a pool of classifiers is trained by varying their hyperparameters (i.e. SVMs); then (ii), the most accurate of these are selected; and (iii) the pair of classifiers minimizing their agreement in predictions is chosen to form the ensemble. Finally, the prediction is triggered as soon as both classifiers in the ensemble predict the same class value. In this case, the monitored confidence measure is binary (agreement or disagreement), there is no trigger threshold and thus this trigger model is free of hyper-parameters [^4] The Reject approach involves choosing the number of classifiers trained in step (i) and selected in step (ii) that could be considered as hyper-parameters of the monitored confidence measure..
$\bullet$ Hidden Markov Models (HMMs) are naturally suited to the classification of online sequences. An HMM is learned for each class, and at each time step $t$, the class to be preferred is the one with the highest a posteriori probability given $\mathbf{x}_{t}$. However, the decision to make a prediction now or to postpone it must then involve a threshold so that the prediction is only made if the a posteriori probability of the best HMM is sufficiently high or is greater than that of the second-best. In reaction to this, ? (?) propose to replace the standard HMM with imprecise HMMs based on the concept of credal classifier. This eliminates the need to choose a threshold, since a decision is made when one classification “dominates” (according to a criterion based on probability intervals) all the others.
$\bullet$ Rather than considering only the largest value predicted by the classifier, it is appealing to consider also the difference with the second largest value, since a large difference points to the fact that there is no tie between predictions to expect.
This is one dimension used in the Stopping Rule (SR) approach (?). Specifically, the output of the system is defined as:
$g(h({\mathbf{x}}_{t}))=\begin{cases}\emptyset&\mbox{if extra measures are queried;}\\ \hat{y}=\operatorname*{arg\,max}_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})&\mbox{when }\gamma_{1}\,p_{1}+\gamma_{2}\,p_{2}+\gamma_{3}\,\frac{t}{T}>0\end{cases}$ (9)
where $p_{1}$ is the largest posterior probability $p(y|{\mathbf{x}}_{t})$ estimated by the classifier $h$, $p_{2}$ is the difference between the two largest posterior probabilities, and $\frac{t}{T}$ represents the proportion of the incoming time series at time $t$. The parameters $\gamma_{1}$, $\gamma_{2}$ and $\gamma_{3}$ are learned from the training set.
$\bullet$ Using the same notations as SR, the Early Classification framework based on class DIscriminativeness and RELiability (Ecdire) (?) finds the earliest timestamp for which a threshold applied on $p_{1}$ is reached (defined as in Equation 9). Then, the quantity $p_{2}$ is monitored, and a second threshold is applied to trigger the prediction.
$\bullet$ A different class of methods relies on searching telltale representations of subsequences, such that if the incoming time sequence $\mathbf{x}_{t}$ matches one or more of these representations, then its class can be predicted. Typically, these representations take the form of shapelets that discriminate well one class from the others (?). For instance, the Early Distinctive Shapelet Classification (Edsc) method learns a distance threshold for each shapelet, based on the computation of the Euclidean distance between the considered subsequence and all others valid subsequences in the training set (?). It selects a subset of them, based on a utility measure that combines precision and recall, weighted by the earliness. A prediction is made as soon as $\mathbf{x}_{t}$ matches one of these shapelets well enough. Because this family of methods is computationally expensive, extensions have been developed to reduce the computational load (?, ?). Other extensions aimed at improving the reliability of the predictions (?, ?), and tackling multivariate time series (?, ?, ?, ?).
3.1.2 Sequence-based decision criterion
Other approaches propose sequence-based confidence measures specifically designed for the ECTS problem.
$\bullet$ The Effective Confidence-based Early Classification (Ecec) (?) proposes a confidence measure based on the sequence of predicted class values, from the first one observed to the current timestamp. At each time step, this approach exploits the precision of the classifier to estimate the probability for each possible class value $y\in{\cal Y}$ of being correct if predicted. Then, assuming that successive class predictions are independent, the proposed confidence measure represents the probability that the last class prediction is correct given the sequence of predicted class values. The proposed confidence measure is monitored over time, and prediction is triggered if this measure exceeds a certain threshold $\gamma$ tuned as the single hyper-parameter.
$\bullet$ The Teaser (Two-tier Early and Accurate Series classifiER) (?) approach considers the problem of whether or not a prediction should be triggered as a classification task, the aim of which is to discriminate between correct and bad class predictions. As the authors point out, the balance of this classification task varies according to the time step considered $t\in[1,T]$. Indeed, assuming there is an information gain over time, there are fewer and fewer bad decisions as new measurements are received (or even no bad decisions after a while, i.e. $\forall\;t>t^{\prime}\;(0<t^{\prime}\leq T)$ for some datasets). To exploit this idea, a collection of one-class SVMs is used, learning hyper-spheres around the correct predictions for each time step. A prediction is triggered when it falls within these hyper-spheres for $\nu$ consecutive time steps ($\nu$ being a parameter of the method).
$\bullet$ The Second-Order Confidence Network approach (Socn) (?) considers, as does Teaser, the same classification task aiming to discriminate between correct and bad predictions. To learn this task, a transformer (?) is used, taking as input the complete sequence of conditional probabilities estimated by the classifier $h$, from the first time step, up to the current time step. A confidence threshold $\nu$ is learned by minimizing the same cost function as (?), above which the prediction is considered reliable and therefore triggered.
3.2 Anticipation-based methods
One way of designing approaches that anticipate future measurements is to achieve classification of an incomplete time series while guaranteeing a minimum probability threshold according to which the same decision would be made on the complete series. This is the case of the Reliability Classification (RelClass) approach (?). Assuming that the measurements are i.i.d. and generated by a Gaussian process, this approach estimates $p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})$ the conditional probability of the entire time series ${\mathbf{x}}_{T}$ given an incomplete realization ${\mathbf{x}}_{t}$ and thus derives guarantees of the form:
[…]
where ${\mathbf{x}}_{T}$ is a random variable associated with the complete times series, $\gamma$ is a confidence threshold, and $h_{T}$ is the classifier learned over complete times series. At each time step $t$, $p(h_{T}({\mathbf{x}}_{T})=y|{\mathbf{x}}_{t})$ is evaluated and a prediction is triggered if this term becomes greater than the threshold $\gamma$, which is the only hyper-parameter to be tuned.
Another way of implementing anticipation-based approaches is to exploit the continuations of training time series, which are full-length. One of the first method for ECTS (?) has been derived into such an anticipation-based approach (?). The first, called Early Classification on Time Series (Ects) (?), exploits the concept of Minimum Prediction Length (MPL), defined as the earliest time step for which the predicted label should not change for the incoming time series $\mathbf{x}_{t}$ from $t$ to $T$. This is estimated by looking for the 1NN of $\mathbf{x}_{t}$ in the training set, and check whether from $t$ onward, its predicted label did not change. To be more robust, the MPL is defined based on clusters computed on full-length training time series to estimate the best decision time. The approach has been extended later on to speed up the learning stage (?). This method looks in its own way at the likely future of $\mathbf{x}_{t}$ i.e. an incomplete time series belongs to a cluster whose continuations are known - and thus can be considered as an anticipation-based method.
? (?) present a method that claims explicitly to be “non myopic” in that a decision is taken at time $t$ only insofar as it seems that no better time for prediction is to be expected in the future. In order to do this, the family of Economy methods estimates the future cost expectation based on the incoming time series $\mathbf{x}_{t}$. This can be done since the training data consists of full-length time series and therefore a Learning Using Privileged Information (LUPI) (?) is possible.
More formaly, the objective is to trigger a decision when $\mathbb{E}_{y,\hat{y}}[\mathcal{L}(\hat{y},y,t)|\mathbf{x}_{t}]$ is minimal, with:
$\mathbb{E}_{y,\hat{y}}[\mathcal{L}(\hat{y},y,t)|\mathbf{x}_{t}]=\sum_{y\in{\cal Y}}P(y|\mathbf{x}_{t})\,\sum_{\hat{y}\in{\cal Y}}P(\hat{y}|y,\mathbf{x}_{t})\,C_{m}(\hat{y}|y)\;+\;C_{d}(t)$ (10)
A tractable version of Equation 10 has been proposed by introducing an additional random variable which is the membership of $\mathbf{x}_{t}$ to the groups of a partition ${\cal G}$:
$\mathbb{E}_{y,\hat{y}}[\mathcal{L}(\hat{y},y,t)|\mathbf{x}_{t}]=\sum_{g_{k}\in{\cal G}}P(g_{k}|\mathbf{x}_{t})\sum_{y\in{\cal Y}}P(y|g_{k})\sum_{\hat{y}\in{\cal Y}}P(\hat{y}|y,g_{k})C_{m}(\hat{y}|y)+C_{d}(t)$ (11)
In technical terms, training approaches from the Economy framework involve estimating the three probability terms of Equation 11, for the current time step $t$, as well as for future time steps $t+\tau\in[t+1,T]$, with:
  • $P(g_{k}|\mathbf{x}_{t})$ the probability of $\mathbf{x}_{t}$ belonging to the groups $g_{k}\in{\cal G}$,
  • $P(y|g_{k})$ the prior probability of classes in each group,
  • $P(\hat{y}|y,g_{k})$ the probability of predicting $\hat{y}$ when the true class is $y$ within the group $g_{k}$.
A key challenge in this framework is to design approaches achieving the most useful partition for predicting decision costs expectation. In the first article which presents this framework (?), a method, called Economy-$K$, is designed as follows. (i) A partition of training examples is first performed by a K-means algorithm ; (ii) then a simple model uses the Euclidean distance as a proxy of the probability that $\mathbf{x}_{t}$ belongs to each group; (iii) the continuation of training time series within each group is exploited to predict the cost expectation for future time steps.
In order to avoid the clustering step with the associated choice of hyper parameters, (?) presented a variant called NoCluster which uses the 1 nearest neighbor in the training set in order to guess the likely future of $\mathbf{x}_{t}$.
Then, Economy-$\gamma$ was introduced in (?) which relies on a supervised method to define a confidence-based partition of training time series. The algorithm, dedicated to binary classification problems, is designed as follows: (i) a collection of partitions is constructed by discretizing the output of each classifier $\{h_{i}\}_{i\in[1,T]}$ into equal-frequency intervals, the groups thus formed correspond to confidence levels for each time step; (ii) at the current time $t$, the incoming time series $\mathbf{x}_{t}$ belongs to only one group, since the output of the classifier $h_{t}$ falls within a particular confidence level; (iii) then, a Markov chain model is trained to estimate the probabilities of the future time step confidence levels. Economy-$\gamma$-max (?) generalizes this approach to multi-class problems, aggregating the multiple conditional probabilities in the classifiers’ output by using only the most probable class value.
Finally, Calimera (?) uses anticipation about the future from another perspective. Instead of trying to guess the likely continuation of $\mathbf{x}_{t}$ which allows one to compute expected future costs, and therefore to wait until there seems no better time to make a prediction, their method is based on predicting directly the difference in cost between predicting the class now or wait at least one more time step. If this difference is positive, then it is better to postpone the prediction. They advocate furthermore, that a calibration step should intervene above the regression in order to make a decision.
3.3 Reinforcement Learning based methods
To learn an ECTS function, it is possible to use an agent that learns what to do by exploring the outcomes associated with different decision strategies as it interacts with the world. The ECTS problem can therefore be recast as a reinforcement learning (RL) problem.
In RL, an agent must learn to associate an action $a$ with each observable state $s$ so that the expected gain be maximized. Let us suppose that the agent-environment interactions break naturally into subsequences which we call episodes
At each time step $t$, the agent perceives the environment’s state $s_{t}$ (i.e. $\mathbf{x}_{t}$), choses an action $a_{t}$ (i.e. either make a prediction now and measure the gain $G_{t}=-\mathrm{C}_{m}(\hat{y}|y)-\mathrm{C}_{d}(\hat{t})$, or postpone the decision and receive the next measurement $x_{t+1}$) according to its current policy $\pi$, where $\pi(a|s)=p(a_{t}=a|s_{t}=s)$ is the probability that action $a$ is taken given the state $s$.
The goal is for the agent to learn an optimal policy $\pi^{\star}$ from sequences of interactions with its environment $\langle s_{t},a_{t},s_{t+1}\rangle$. This can be done by computing the utility function $Q(s,a)$ defined for all (state, action) pairs. By definition:
$Q_{\pi}(s,a)\;=\;\operatornamewithlimits{\mathbb{E}_{\>\pi}^{\>}}\left[G_{t}|S_{t}=s,A_{t}=a\right]$ (12)
and the optimal policy can be derived from:
$Q^{\star}(s,a)\;=\;\operatornamewithlimits{max}_{\pi}Q_{\pi}(s,a)$ (13)
It suffices at each observed state $s$ to chose action $a^{\star}$ such as:
$a^{\star}\;=\;\operatorname*{arg\,max}_{a}Q^{\star}(s,a)$ (14)
One way to learn the function $Q^{\star}$ is using the Q-learning algorithm (?, ?) [^5] Note that other approaches are possible in the reinforcement learning scenario, like learning the utility function $V(s)$ and using TD-learning, or even learning the policy $\pi$ directly. The Q-learning approach is however widely used. and its variants for continuous definition of environment’s states, such as Deep Q-learning (?).
In End-to-end approaches, only one function is responsible for both stopping and making a prediction about the class of $\mathbf{x}_{t}$ whereas, in separable approaches involves two combined functions, respectively dedicated to triggering the prediction and to the classification itself. In the case of Reinforcement Learning based approaches, this distinction takes the following form:
  1. 1.
    The separable approach. RL can be used to learn the Trigger function only, once the classifiers $h_{t}$ have been learnt. In that case, the set of actions $\mathcal{A}_{t}$ at each time step $t$ is restricted to two elements: {‘decide now’, ‘postpone decision’}. The $Q$ function evaluates for each state the expected gain for each of the two possibilities, allowing one to decide what to do at each time step.
  2. 2.
    The End-to-End approach. RL can also be used to learn at once both when to trigger a prediction and what prediction to make. In principle, it suffices to extend the set of actions to $\mathcal{A}_{t}=\{`\textit{postpone decision'},c_{1},\ldots,c_{N}\}$, where there are $N$ classes. Either, the agent postpones the decision, or it predicts a class for the incoming time series $\mathbf{x}_{t}$.
In the literature, Reinforcement Learning-based ECTS approaches frequently include Deep Learning modules:
Separable RL approaches:
The Earliest (Early and Adaptive Recurrent Label ESTimator) uses a RNN architecture (?) to make the prediction and a Reinforcement Learning agent trained jointly using policy gradient, to trigger prediction or not. If a prediction is triggered, the hidden representation given by the RNN is sent to a Discriminator, whose role is to predict a class, given this representation. The model has been adapted to deal with irregularly sampled time series (?). ? (?) extend the ECTS framework to channel filtering, using here also Reinforcement Learning.
End-to-end RL approaches:
? (?, ?) use a Deep Q-Network (?), alongside a specifically designed reward signal, encouraging the agent to find a good trade-off between earliness and accuracy. Those types of approaches also naturally extend to online settings where time series are not of fixed length. (?) introduce EarlytStop-RL, in which model-free RL is used to address the problem of early diagnosis of lung cancer.
3.4 Deep Learning based methods
Alongside the Reinforcement Learning based approaches, there exist deep learning methods that do not use RL.
The Decouple ETSC Network (Detscnet) (?) architecture leverages a gradient projection technique in order to jointly learn two sub-modules: one for variable-length series classification, the other for the early exiting task.
The End-to-end Learned Early Classification of Time Series method (Elects) leverages a LSTM architecture, adding a stopping prediction head to the network and adapting the loss function to promote good early predictions (?).
4 Experiments & Results
This section presents the extensive set of experiments carried out in order to provide a consistent and fair evaluation of a wide range of the existing literature’s methods. We first describe the experimental protocol used. We then turn to the experiments and their results. Figure 4 provides a synthetic view of the organization of these experiments.
  • Section 4.1 introduces the experimental protocol as well as the global evaluation methodologies.
  • In Section 4.2, the main state-of-the-art and the three baseline methods are evaluated using a widely used cost setting, i.e. with a a binary balanced misclassification cost and a linear delay cost.
  • In Section 4.3, methods are tested in an anomaly detection scenario where the misclassification cost matrix is severely imbalanced, with false negatives being much more costly than false positives, and where the delay cost is no longer linear with time but increases exponentially with time.
  • Finally, Section 4.4 briefly describes a set of other experimental setups, derived from either standard setting or the anomaly detection one, including for instance testing the impact of z-normalization. Complementary results can be found in Appendix B.
[…]
Figure 2: Experiments diagram: each line corresponds to one (or two) full benchmark runs, representing in total twelve full benchmarks. While this Section mainly discusses results about the cost settings in Subsections 4.2 and 4.3, many other alternative experiments are briefly analyzed in Subsection 4.4 and are more detailed in the Appendix B.
The experiments presented here aim to evaluate the effects of design choices on method performance and thus to provide answers to the questions:
  • Do anticipation-based methods perform better than blind or myopic ones?
  • Do methods that are cost-informed for their decision (i.e. explicitly estimating costs) perform better than methods that are cost-uninformed (e.g. confidence-based) (see Section 2.3.2).
  • How the various methods fare when modifying the form of the delay cost and/or the misclassification cost matrix?
4.1 Experimental Protocol
This section covers the shared part of the experimental protocol for all experiments irrespective of the choice of the cost functions (see Sections 4.2 and 4.3 for this).
4.1.1 Evaluation of the performance
The ECTS problem explicitly balances the pressure to make a correct prediction and a time pressure. The correctness of the prediction is measured by the misclassification cost $\mathrm{C}_{m}(\hat{y}|y)$ where $\hat{y}$ is the prediction and $y$ is the true class. The time pressure is sanctioned by a delay cost $\mathrm{C}_{d}(t)$ that is assumed to be positive and, in most applications, an increasing function of time. Sometimes, the two can be combined when the misclassification cost is function of the time : $\mathrm{C}_{md}(\hat{y}|y,t)$.
Therefore, for each test times series $\mathbf{x}^{i}$, an ECTS method incurs a cost assumed to be of the following additive form: $\mathrm{C}_{m}(\hat{y}_{i}|y_{i})\,+\,\mathrm{C}_{d}(\hat{t})$, where $\hat{t}$ is the time when the system decided to make a prediction, this prediction being $\hat{y}_{i}$.
For a test set of $M$ time series, the average cost of a method is:
$AvgCost_{\text{test}}\;=\;\frac{1}{M}\sum_{i=1}^{M}C_{m}(\hat{y}_{i}|y_{i})+C_{d}(\hat{t}_{i})$ (15)
This is accordingly the criterion with which we evaluated the methods in the experiments reported in this paper.
In addition, in order to assess how the methods adapt to various balances between the misclassification and the delay costs, we vary the settings of these costs by weighting them during training and testing. The performance of the methods is therefore evaluated using the weighted average cost, as defined in Equation 16, for different values of the costs balance $\alpha$, ranging from 00 to $1$, with a $0.1$ step:
$AvgCost_{\alpha}=\frac{1}{N}\sum_{i=0}^{N}\alpha\times C_{m}(\hat{y}_{i}|y_{i})+(1-\alpha)\times C_{d}(\hat{t}_{i})$ (16)
Small values of $\alpha$ correspond to a high delay cost and a small misclassification cost ; inversely, large values of $\alpha$ give more weight to the misclassification cost with a lower delay cost.
4.1.2 Optimization of the parameters of the methods
[…]
4.1.3 Comparing the trigger methods
In order to carry out a fair comparison between the tested methods, we isolated as far as possible the triggering function, responsible for deciding when to stop receiving measurements, from the prediction one, responsible for predicting a label to the incoming time series. As advocated by ? (?), we have chosen the MiniROCKET algorithm (?) to be the base classifier for all methods. It is indeed recognized as among the best performing classifier in the Time Series classification literature as well as one of the fastest one. Of course, distinguishing the decision component from the prediction one is only possible for the “separable” methods. In our experiments, we chose to not evaluate “end-to-end” methods, leaving that for future work.
Trigger models: Nine trigger models were selected from the literature based on their usage and their performances [^6] The EDSC algorithm (?), even though available in the provided library, is not included in the following experiments, due to high space and time complexity (which hinders fair comparisons).
  • Economy-$\gamma$-Max (?): triggers a decision if the predicted cost expectation is the lowest at time $t$ when compared with the expected cost for all future time steps (cf. Section 3.2, Anticipation-based).
  • Calimera (?): triggers a decision when a regressor model which predicts the difference between the current observed cost and the minimum cost in the future is negative (cf. Section 3.2, Anticipation-based).
  • Stopping Rule (?): uses a trigger function based on a linear combination of confidence estimates and a delay measure linear on time (cf. Section 3.1, Confidence-based).
[…]
  • Ecec (?): defines a confidence measure, based on the aggregated confidence of the predictions up to time $t$, and triggers a prediction if it exceeds a threshold, tuned by explicit grid-search (cf. Section 3.1, Confidence-based).
  • Ecdire (?): determines “safe” timestamps, based on classifier performance, from which predictions about possible classes can be made. Predictions cannot be triggered if those timestamps have not been reached. In addition, the difference between the two highest predicted probabilities must also exceed a certain threshold. (cf. Section 3.1, Confidence-based).
  • Ects (?): computes the first time $t$ for which nearest neighbors of the incoming time series $\mathbf{x}_{t}$ in the training set were given a label that did not change by the classifier (cf. Section 3.2, Anticipation-based).
All these methods have been re-implemented using Python, reproducing results close to the published ones. Except the code for the Ects implementation, which has been taken from ? (?). Hyper-parameters are the ones chosen in the original published methods. Code to reproduce the experiments is available publicly at https://github.com/ML-EDM/papers_experiments/tree/main/ects_reloading.
Baselines: Furthermore, in order to evaluate the benefits, if any, of the various methods, it is telltale to compare them with simple ones. We chose three such baselines:
  • Asap (As Soon As Possible) always triggers a prediction at the first possible timestep.
  • Alap (As Late As Possible) always waits the complete series to trigger the prediction.
  • Proba Threshold is a natural, confidence-based, cost-informed at training time, baseline: it triggers a prediction if the estimated probability of the likeliest prediction exceeds some threshold, found by grid search (cf. Section 3.1, Confidence-based).
4.1.4 Calibration of the classifications
Like ? (?), we add a calibration step when learning the classifiers, i.e. Platt’s scaling (?). Indeed, as we are dealing with collections of independently trained classifiers, the prediction scores may not remain consistent with one another over the time dimension. However, the trigger methods usually have their parameters set with the same values for all time steps. This is the case, for example with the Proba Threshold approach. In addition, some approaches such as Calimera and Economy-$\gamma$-Max exploit the estimated posterior probabilities $\{p(y|\textbf{x}_{t})\}_{y\in\mathcal{Y}}$ to estimate the future cost expectation. It is therefore highly desirable for all classifiers, at all times, to have their output calibrated.
4.1.5 Datasets and training protocol
Datasets [^7] All original datasets of the paper can be downloaded, already prepared and splitted, from https://urlz.fr/qRqu: In order to be able to directly compare our results to past experiments, we first use the usual TSC datasets from the UCR Archive (?) with the default split. In total, we have used 77 datasets from the UCR Archive, i.e. the ones with enough training samples to satisfy our experimental protocol end-to-end (blue cylinder in Figure 4). In this way, most of the datasets used by either ? (?) and ? (?) or by ? (?) are contained in our experiments.
A second collection of non z-normalized data sets is also provided. In this way, the associated potential information leakage is avoided (see Section 1). Any difference with the performance obtained on the z-normalized data sets can thus signal the danger of z-normalization with firm evidence. Considering the limited amount of non z-normalized datasets within the UCR archive (?), we have decided to look for complementary new datasets so as to provide another collection of datasets. To this end, the Monash archive for extrinsic regression (?), provided 20 new time series datasets, for which we have discretized the the numerical target variable into binary classes based on a threshold value. For instance, if this threshold is equal to the median value of the regression target, the resulting classification datasets will be balanced in term of classes (as in Section 4.2). Note that this threshold can be chosen differently to get imbalanced datasets (as in Section 4.3.2), several thresholds could also be used to increase number of classes. As a result, we get a new set of classification tasks, as has recently been done by ? (?). Finally, 35 datasets have been gathered: 15 from the original archive and 20 from the Monash extrinsic regression archive. (orange cylinder in Figure 4).
Splitting strategy:
When not using predefined splits, the train sets are split into two distinct sets in a stratified fashion: a first one to train the different classifiers, corresponding to 60% of the training set and another one to train the trigger model, trained over the 40% left. The set used to train the classifiers, is itself split into two different sets in order to train calibrators, using 30% of the given data. Because of this procedure, we have been led to exclude some of the datasets, due to their limited training set size.
All the experiments have been performed using a linux operating machine, with an Intel Xeon E5-2650 2.20GHz (24-cores) and 252GB of RAM. Proceeding all datasets (including both blue and orange cylinders) over all competing approaches takes between 9-10 days, using MiniROCKET classifier, which is the most efficient tested.
4.2 Experiments with balanced misclassification and linear delay costs
This first setting is the one most widely used in the literature to date.
4.2.1 Cost definition
The misclassification cost is symmetrical and balanced, the delay cost is linear. They can be defined as follows:
$C_{m}(\hat{y}|y)$ $=\mathbb{1}(\hat{y}\neq y)$
$C_{d}(t)$ $=\frac{t}{T}$
Thus, for each dataset, the AvgCost is bounded between 0 and 1.
4.2.2 Results and analysis
For comparability reasons, this first set of experiments is analyzed over the classical ECTS benchmark used in the literature so far (blue cylinder in Figure 4). Results over the new, non z-normalized, datasets can be found in Appendix B.
(a) Evolution of the mean ranks, for every $\alpha$, based on the AvgCost metric. Shaded areas correspond to 90% confidence intervals.
(b) Alpha is now fixed to $\alpha=0.5$. Wilcoxon signed-rank test labeled with mean AvgCost.
Figure 3: The ranking plot (a) shows that, across all values of $\alpha$, a top group of four approaches distinguishes itself. The significance of this result is supported by statistical tests. Specifically, we report this for $\alpha=0.5$ as shown in (b).
Figure 3(b) provides a global views about the relative performances of the tested methods. The Wilcoxon-Holm Ranked test provides an overall statistical analysis. It examines the critical difference among all techniques to plot the method’s average rank in a horizontal bar. Lower ranks denote better performance, and the methods connected by a horizontal bar are similar in terms of statistical significance. When evaluated by their average rank on all data sets with respect to the average cost (Equation 16), here for $\alpha=0.5$, four methods significantly outperform the others:
        Methods Confidence Anticipation Cost informed
Stopping Rule train
Proba Threshold train
Economy-$\gamma$-max train & test
Calimera train
Figure 3(a) allows a closer look, this time varying the relative costs of misclassification and delaying prediction using Equation 16, where a small value of $\alpha$ means that delay cost is paramount. $90\%$ level confidence intervals have been computed using bootstrap [^8] Resample with replacement has been done a large number of times (10.000 $\times$) and are reported as shaded colors on the figure. The statistic of interest is studied, here the mean, by examining the bootstrap distribution at the desired confidence level.. Again, the same four methods top the others for almost every values of $\alpha$. Not surprisingly, the baseline Asap (predict as soon as possible) is very good when the delay cost is very high, while Alap (predict at time $T$) is very good when there is no cost associated with delaying decision.
It is remarkable that, in this cost setting, the simple Proba Threshold method exhibits a strong performance for almost all values of $\alpha$. It is therefore worth including in the evaluation of new methods. However, while Figures 3(a), 3(b) are useful for general analysis, they do not provide insights about how the Accuracy vs Earliness trade-off is optimized for each of the competitor. Figure 4 provides some explanation for this.
Figure 4: Pareto front, displaying for each $\alpha$ the Accuracy on the $y$ axis and $Earliness$ on the $x$ axis. Best approaches are located on the top left corner. In zoomed boxes, on the right of the Figure, points corresponding to a single $\alpha$ are highlighted, while other points are smaller and gray. Each of the trigger model is optimizing the trade-off in its own way, resulting in many different approaches having points in the Pareto dominant set.
In this figure, the two evaluation measures: ‘Accuracy’ and ‘Earliness’, are considered as dimensions in conflict. The Pareto front is the set of points for which no other point dominates with respect to both Accuracy and Earliness. It is drawn here when varying their relative importance using $\alpha$ (in the set $\{0,0.1,0.2,\ldots,1.0\}$).
[…]
Each of the leading methods Stopping Rule, Proba Threshold, Economy and Calimera have at least one point on the Pareto front and generally exhibits a combined performance very close to it. A closer look reveals how each approach optimizes the earliness vs. accuracy trade-off differently for a fixed cost. If we consider $\alpha=0.8$, for example, it appears that Economy takes its decision earlier than Proba Threshold, itself being more precocious than Ecec. Because this is also an area of problems where the delay cost is low, by doing so, Economy prevents itself from benefiting from waiting for more measurements and increasing its performance. Hence its slight downward slope on Figure 3(a) for high values of $\alpha$.
It is worth noting that the two naive baselines Asap and Alap perform better than the majority of approaches on seven $\alpha$ values out of ten. This is specially the case when the delay cost is large, i.e. for $\alpha\in[0.1,0.3]$, for which the Asap baseline is as competitive as top performers. Globally, the performance of Proba Threshold is remarkable in this cost setting. Even though, it is simply based on a single threshold on the confidence in the current prediction, its performance makes it one of the best methods.
The results computed over the proposed datasets ensemble (i.e. orange cylinder) are displayed in Figure 8 of Appendix B. No significant changes can be observed in the ranking of competing approaches.
One question is how such simple method, like Proba Threshold, can adapt to scenarios where the misclassification and delay costs, not being symmetrical for the misclassification cost, and not linear for the delay one, reflect other application settings.
4.3 Experiments with unbalanced misclassification and non-linear delay costs
While the previous section has provided a first assessment of how the various methods adapt to different respective weights for the misclassification and the delay costs, it nonetheless assumed that the misclassification costs were balanced (e.g. 0 if correctly classified and 1 otherwise) and that the delay cost was a linear function of time.
There are however applications where these assumptions do not hold, for instance predictive maintenance or hospital emergency services, are characterized by (i) imbalanced misclassification costs (e.g. it is more costly to have to repair a machine than to carry out a maintenance operation that turns out not being necessary) and by (ii) non linear delay cost (e.g. usually, the later the surgical operation is decided, the costlier it is to organize it and the larger the risk for the patient). In the following, we call all applications presenting these characteristics “anomaly detection” applications.
The question arises as to how the various ECTS algorithms behave in this case, depending on their level of cost awareness and whether or not they are anticipation-based. This is what is investigated in the series of experiments reported in this section.
(a) Exponential delay cost:
$C_{d}(t)=(1-\alpha)\exp(\frac{t}{T}*\log(100))$
(b) Misclassification cost matrix
(three class problem)
Figure 5: Representative delay cost (a) and misclassification ones (b) for an anomaly detection scenario. In our experiments, $\alpha\in[0,1]$.
4.3.1 Cost definition for anomaly detection
In order to study the behavior of the various algorithms on scenarios corresponding to anomaly detection, we set the unbalanced misclassification cost matrix such that a false negative (i.e. missing an anomaly) was 100 times costlier than a false positive (i.e. wrongly predicting an anomaly) (see Figure 5(b)). For this last situation, the cost was arbitrarily set to 1. The delay cost is defined as an exponential function of time. In order to have a cost commensurable with the misclassification one, we decided that waiting for the entire time series to be seen, at $T$, would cost 100 (see Figure 5(a)), starting at 1 for $t=0$ and reaching 100 when $t=T$.
4.3.2 Results and analysis
In this part, as a new cost setting is explored, there is no need to produce comparable results from previous works. Thus, we choose to use the new non z-normalized datasets collection (orange cylinder in Figure 4). In order for the imbalanced misclassification cost to make sense, those datasets have been altered so that the minority class represents 20% of all labels. As explained in Section 4.1, some extrinsic regression datasets are turn into classification ones. In these cases, the threshold value has been set to the second decile of the regression target. For the original classification datasets, the minority class has been sub-sampled when necessary.
[…]
figurec
(a) Evolution of the mean ranks, for every $\alpha$, based on the AvgCost metric. Shaded areas correspond to 90% confidence intervals.
(b) Alpha is now fixed to $\alpha=0.5$. Wilcoxon signed-rank test labeled with mean AvgCost.
Figure 6: The ranking plot (a) shows that, across all $\alpha$, a top group composed by three approaches distinguish. This result is significant as supported by statistical tests. Specifically, for $\alpha=0.5$ as shown in (b).
Results from the Wilcoxon-Holm Ranked test (both regarding the average rank and the value for AvgCost) (see Figure 6(b)) and from the AvgCost plot (see Figure 6(a)) with varying values of $\alpha$ (in Equation 16) show that now the best method overall is Economy which is both cost-informed at training and testing time, beside being anticipating-based. However, Stopping-Rule is a very strong contender while being cost-informed at training time but not at testing time and confidence-based. There is a reason for it. When Stopping Rule equals or overpasses Economy, this is for high values of $\alpha$ when the delay cost loses its importance, therefore leaving the misclassification cost to reign and confidence-based methods to be good.
It may come as a surprise that Calimera lags behind Economy for $\alpha\in[0,0.4]$, despite being similarly based on the estimation of future cost expectation. One reason for this is that the cost expectation is achieved by considering only the predicted class. This poor estimate of the cost expectancy becomes critical when the delay cost is important.
Similarly, Proba Threshold is surprisingly good in this scenario, even if it is no longer in the top tier. Looking solely at prediction confidence, we might expect it to be blind to the rapid increase in delay cost in the anomaly detection scenario. However, it is noticeable that the cost of delay only increases sharply after around 60% of the complete time series has been observed, which is generally sufficient to exceed the confidence threshold. Hence, Proba Threshold does not suffer from high delay costs that are to come, and exhibits a good performance here.
Figure 7 plots the Pareto front considering two axes based on decision costs. The horizontal axis corresponds to the average delay cost incurred for each example, normalized by the worst delay cost paid at $t=T$. It is better to be on the left of the $x$-axis. The vertical axis corresponds to 1 minus the misclassification cost incurred for each example, normalized by the worst prediction cost. It is better to be high on the $y$-axis.
We observe that the Pareto front is composed almost exclusively of points corresponding to the Economy method. This is consistent with the evaluation based on the AvgCost metric. This figure highlights the fact that the design of approaches capable of handling arbitrarily parameterized decision costs requires a cost-informed application framework.
Figure 7: Pareto front, displaying for each $\alpha$, the normalized version of the AvgCost, decomposed over delay and misclassification cost on $x$-axis and $y$-axis respectively. Best approaches are located on the top left corner. Due to the exponential shape of the delay cost, the $x$-axis is on log scale.
4.4 Other experiments: ablation and substitution studies
In this section, complementary experiments, namely ablation studies as well as sanity checks are briefly discussed. For the sake of brevity, the figures supporting the analysis are reported in Appendix B.
Impact of removing calibration
? (?) assert that calibration of the classifiers is paramount for the performance of ECTS algorithms. In order to test this claim, we have repeated the experiments removing the calibration step. The examples used for calibration have been removed as well during training, so that all else remains the same as before.
The results of Figure 11 show that indeed Calimera suffers greatly if no calibration is done. Indeed, this approach relies on estimating the expectation of future costs via a regression problem, and a miscalibration may have a negative impact on the built targets. For its part, Proba Threshold suffers somewhat mildly. This is no surprise as they rely on a single threshold on the confidence of the prediction for all time steps.
Impact of the choice of base classifier
All methods have been compared using the same classifier: MiniROCKET so that only the decision components differ. However, the choice of the base classifiers could induce a bias favoring or hampering some methods. In order to clarify this, we have repeated the experiments replacing MiniROCKET with two base classifiers: WEASEL 2.0 (?), and the XGBoost classifier (?) using features produced by tsfresh (?). Both of these classifiers have already been tested within the ECTS literature by ? (?, ?) and ? (?) respectively. Figure 12 and 13 in Appendix B report the results respectively with these two classification methods. One can observe that the results are not significantly altered with the same overall ordering of the methods when varying the value of $\alpha$. Furthermore, our results on AvgCost show that performances tends to be better for all methods using MiniROCKET. It is thus to be preferred given its simplicity and good performances.
Impact of z-normalization
Considering the newly proposed ensemble of datasets, we were not able to identify any problems of information leakage over time. This inconclusive result simply indicates that the variance of the time series measurements is not informative for theses datasets, which still could be the case considering past published results. For further details, please refer to Section B.5 of Appendix B.
5 Conclusion
The first contribution of this research work is the coding of all the methods tested and making the codes available in a repository open to everyone. In this way, the experiments reported can be duplicated and further studies carried out. Furthermore, the deposited datasets and the experimental framework provide a ground for fair comparisons bewteen competing methods. We claim that the AvgCost is the appropriate measure by which to evaluate the performance of the methods. This is indeed what will be “paid” at the end of the day by a practitioner using a method. We have accordingly characterized a number of methods from the literature.
Our extensive experiments have shown that:
  • It is worthwhile to resort to dedicated ECTS methods, and more so in scenarios like anomaly detection with asymmetrical misclassification cost and exponential delay cost.
  • However, it is noticeable that Proba Threshold, a baseline method, is surprisingly good overall in the standard setting with symmetrical misclassification cost and linear delay cost, exhibiting comparable performance to confidence-based myopic methods such as Stopping-Rule and anticipation-based cost-informed ones such as Calimera and Economy.
  • Calibration of the classifiers has a large impact on some methods such as Calimera in particular, less so on other methods like Proba Threshold and Ecdire.
In this paper, we have proposed a reading guide to highlight the main characteristics of ECTS methods, namely (i) the importance of the two components: decision and prediction which are distinct in the “separable” architecture and not in the “end-to-end” one, (ii) the distinction between anticipation-based and myopic methods, and (iii) between cost-informed and cost-uninformed techniques. On the basis of these dimensions, it becomes easy to imagine new methods that combine them in original ways, which could lead to new properties and better performance for solving the problem of early classification of time series, which, present in many applications has potentially great impacts.
To go a step further, future work could be carried out to study the literature’s approaches applied in as yet unexplored cost settings. For example, in many applications, the delay cost depends on the true class and the predicted one, and thus a single cost function integrating misclassification and delay costs should then be used. The use of this general cost form requires the adaptation of some state-of-the-art methods and has not yet been studied.
In addition, in real ECTS applications, it is up to the business expert to define the costs, which is not an easy task in practice. Future work could study the impact of: (i) noisy costs, (ii) and cost drift between training and testing stages. This would make it possible to identify the most resilient literature approaches.
Finally, in the case of existing separable approaches, the misclassification cost is not exploited for training the classification function. Future work could investigate the interest of using cost-sensitive classifiers in the case of ECTS.
A Data description
A.1 UCR Time Series Classification datasets
Table 2: UCR TSC datasets : 77 datasets from the UCR archive have been retained to run the experiments over the 128 contained in the full archive. Those are the ones with fixed length, without missing values and with enough training samples to execute our experiments pipeline end-to-end. Italic datasets are not included in experiments using default split for this reason.
Train Test Length Class Type
Data
ACSF1 100 100 1460 10 Device
Adiac 390 391 176 37 Image
Beef 30 30 470 5 Spectro
BeetleFly 20 20 512 2 Image
BME 30 150 128 3 Simulated
Car 60 60 577 4 Sensor
CBF 30 900 128 3 Simulated
Chinatown 20 345 24 2 Traffic
ChlorineConcentration 467 3840 166 3 Sensor
CinCECGTorso 40 1380 1639 4 Sensor
Coffee 28 28 286 2 Spectro
Computers 250 250 720 2 Device
CricketX 390 390 300 12 Motion
CricketY 390 390 300 12 Motion
CricketZ 390 390 300 12 Motion
Crop 7200 16800 46 24 Image
DiatomSizeReduction 16 306 345 4 Image
DistalPhalanxOutlineCorrect 600 276 80 2 Image
Earthquakes 322 139 512 2 Sensor
ECG200 100 100 96 2 ECG
ECG5000 500 4500 140 5 ECG
ECGFiveDays 23 861 136 2 ECG
ElectricDevices 8926 7711 96 7 Device
EOGVerticalSignal 362 362 1250 12 EOG
EthanolLevel 504 500 1751 4 Spectro
FaceAll 560 1690 131 14 Image
FaceFour 24 88 350 4 Image
FacesUCR 200 2050 131 14 Image
FiftyWords 450 455 270 50 Image
Fish 175 175 463 7 Image
FordA 3601 1320 500 2 Sensor
FreezerRegularTrain 150 2850 301 2 Sensor
GunPoint 50 150 150 2 Motion
Ham 109 105 431 2 Spectro
HandOutlines 1000 370 2709 2 Image
Haptics 155 308 1092 5 Motion
Herring 64 64 512 2 Image
HouseTwenty 34 101 3000 2 Device
InlineSkate 100 550 1882 7 Motion
InsectEPGRegularTrain 62 249 601 3 EPG
InsectWingbeatSound 220 1980 256 11 Sensor
ItalyPowerDemand 67 1029 24 2 Sensor
LargeKitchenAppliances 375 375 720 3 Device
Lightning2 60 61 637 2 Sensor
Lightning7 70 73 319 7 Sensor
Mallat 55 2345 1024 8 Simulated
Meat 60 60 448 3 Spectro
MedicalImages 381 760 99 10 Image
MelbournePedestrian 1200 2450 24 10 Traffic
MixedShapesRegularTrain 500 2425 1024 5 Image
MoteStrain 20 1252 84 2 Sensor
NonInvasiveFetalECGThorax1 1800 1965 750 42 ECG
NonInvasiveFetalECGThorax2 1800 1965 750 42 ECG
OSULeaf 200 242 427 6 Image
OliveOil 30 30 570 4 Spectro
PhalangesOutlinesCorrect 1800 858 80 2 Image
Plane 105 105 144 7 Sensor
PowerCons 180 180 144 2 Power
ProximalPhalanxOutlineCorrect 600 291 80 2 Image
RefrigerationDevices 375 375 720 3 Device
Rock 20 50 2844 4 Spectrum
ScreenType 375 375 720 3 Device
SemgHandGenderCh2 300 600 1500 2 Spectrum
ShapesAll 600 600 512 60 Image
SmoothSubspace 150 150 15 3 Simulated
SonyAIBORobotSurface1 20 601 70 2 Sensor
SonyAIBORobotSurface2 27 953 65 2 Sensor
StarLightCurves 1000 8236 1024 3 Sensor
Strawberry 613 370 235 2 Spectro
SwedishLeaf 500 625 128 15 Image
Symbols 25 995 398 6 Image
SyntheticControl 300 300 60 6 Simulated
ToeSegmentation1 40 228 277 2 Motion
Trace 100 100 275 4 Sensor
TwoLeadECG 23 1139 82 2 ECG
TwoPatterns 1000 4000 128 4 Simulated
UMD 36 144 150 3 Simulated
UWaveGestureLibraryX 896 3582 315 8 Motion
UWaveGestureLibraryY 896 3582 315 8 Motion
UWaveGestureLibraryZ 896 3582 315 8 Motion
Wafer 1000 6164 152 2 Sensor
Wine 57 54 234 2 Spectro
WordSynonyms 267 638 270 25 Image
Worms 181 77 900 5 Motion
Yoga 300 3000 426 2 Image
A.2 Proposed, non z-normalized, datasets
Train Test Length Class Type
Data
[…]
Chinatown 20 345 24 2 Traffic
Crop 7200 16800 46 24 Image
DodgerLoopDay 78 80 288 7 Sensor
EOGVerticalSignal 362 362 1250 12 EOG
GestureMidAirD1 208 130 360 26 Trajectory
GunPointAgeSpan 135 316 150 2 Motion
HouseTwenty 34 101 3000 2 Device
InsectEPGRegularTrain 62 249 601 3 EPG
MelbournePedestrian 1200 2450 24 10 Traffic
PLAID 537 537 Vary 11 Device
Rock 20 50 2844 4 Spectrum
SemgHandGenderCh2 300 600 1500 2 Spectrum
SmoothSubspace 150 150 15 3 Simulated
UMD 36 144 150 3 Simulated
[…]
AluminiumConcentration 440 189 2542 2 Environnement
BitcoinSentiment 232 100 24 2 Sentiment
ChilledWaterPredictor 321 138 168 2 Energie
CopperConcentration 440 189 2542 2 Environnement
Covid19Andalusia 142 62 91 2 Santé
DailyOilGasPrices 133 58 30 2 Économie
DhakaHourlyAirQuality 1447 621 24 2 Environnement
ElectricityPredictor 567 243 168 2 Energie
FloodModeling3 429 184 266 2 Environnement
HouseholdPowerConsumption1 1001 430 1440 2 Energie
HotwaterPredictor 245 106 168 2 Energie
MadridPM10Quality_nmv 4845 2077 168 2 Environnement
ParkingBirmingham_eq 1391 597 14 2 Environnement
PrecipitationAndalusia_nmv 470 202 365 2 Environnement
SierraNevadaMountainsSnow 350 150 30 2 Environnement
SolarRadiationAndalusia_nmv 470 202 365 2 Energie
SteamPredictor 210 90 168 2 Energie
TetuanEnergyConsumption 254 110 144 2 Energie
WindTurbinePower 596 256 144 2 Energie
Table 3: New datasets collection : 35 datasets from both the UCR archive (dashed line) and the Monash UEA extrinsic regression archive. When missing values and/or varying lengths, replace missing values with 0 and pad series to maximum length with 0. All of the datasets are not $z$-normalized and thus don’t suffer from potential data leakage. Italic datasets are not included when classes are imbalanced as problems become too difficult for the chosen classifiers.
B Supplementary results
B.1 Additional figures : Standard cost setting
Figure 8: Standard cost setting, non $z$-normalized proposed datasets (orange cylinder). Compared to Figure 3(a), the global ranking is not altered much. One can observe that for $\alpha\in[0.5,0.7]$ the top group is now more populated, gathering the first six approaches, probably due to the limited amount of datasets available in this case.
B.2 Additional figures : Anomaly detection cost setting
Figure 9: Anomaly detection cost setting, original UCR datasets (blue cylinder). Compared to Figure 6(a), one can see that Calimera is now clearly dominating all other methods for all $\alpha$. The global ranking remains globally stable otherwise.
B.3 Removing calibration
Figure 10: Standard cost setting, original UCR datasets (blue cylinder). The calibration step is now removed, i.e. the outputs from the decision function is now simply passed through a softmax function. Both Calimera and Proba Threshold suffer heavily from using uncalibrated scores.
Figure 11: Pairwise comparison (?), calibration ($calib$) vs no calibration ($no\>\>calib$), $\alpha=0.8$. Square colors are indexed on the mean AvgCost difference. For example, Calimera has a lower mean AvgCost when trained over calibrated scores: it appears in dark blue. The Wilcoxon p-value is equal to 0.0288, which is lower than significance level equal to 0.05. Thus, Calimera statistically under-performs when using uncalibrated scores. This is also the case for the Proba Threshold method.
B.4 Changing the base classifier
Figure 12: Standard cost setting, original UCR datasets (blue cylinder). The base classifier is now Weasel 2.0 (?). Results are very close to those exposed in Figure 3(a).
Figure 13: Standard cost setting, original UCR datasets (blue cylinder). The base classifier is now a pipeline including features extraction with tsfresh (?) and classification using XGBoost (?). Results are a bit noisier than those exposed in Figure 3(a).
B.5 Impact of z-normalization
Clearly, using z-normalized datasets is not applicable in practice, as it would require knowledge of the entire incoming time series. In a research context, previous work has used such training sets to test the proposed algorithms. Our goal here, is to assess whether this could have a large impact on the performances. For example, when a normalized time series has a low variance at the beginning, we can expect a high variance in the rest of the series since the mean variance is 1. There is therefore an information leakage that can be exploited by a ECTS algorithm, while this is not representative of what happens in real applications. A proposal such as the one presented in (?), where the z-normalization of available time series is repeated at each time step, has its own problems. In particular, it means that if a single classifier is used for all time steps, the representation of $\mathbf{x}_{t}$ can be different at times $t$ and $t+1$ and all further time steps which can induce confusion for the classifier.
On the one hand, z-normalization induces an information leakage that could help methods to unduly exploit knowledge about the future of incoming time series. On the other hand, any normalization rescale the signal and therefore, potentially, hinder the recognition of telltale features. So, does z-normalization affect the performance of ECTS methods? And if yes, in which way?
In order to answer this question, we took the new datasets collection described in Section 4.1. They are indeed not z-normalized originally. We duplicate and z-normalized them to get a second collection. As explained in Section 4.1, some extrinsic regression datasets has been converted into classification one. Here, the threshold value chosen to discretize the output into binary classes has been set the median of the regression target. In this way, classes within those datasets are equally populated.
In these experiments, the delay cost is linear as in Section 4.2 and as in most of the literature. Figure 14 reports pairwise comparisons done on the 35 datasets. We look at $\alpha=0.8$, as this is the only value for which significant differences are observed. One can see that most of the trigger models do not actually benefit from the z-normalization. Quite the opposite: out of nine trigger models, only one, i.e. Ecdire, actually has a better mean AvgCost when being trained on z-normalized data. Regarding the remaining methods, both Stopping Rule and TeaserAvg performs significantly worse when operating on z-normalized data. Those trends are quite similar for other $\alpha$ values, without any significance on the statistical tests though. Thus, while z-normalization has some impact, since privileged information from the future can be leaked, our experiments, for the proposed datasets collection at least, show that this does not alter the overall results reported in the literature, and are globally in accordance with the results presented in Section 4.
Figure 14: Pairwise comparison (?), $z$-normalization ($Z$) vs no $z$-normalization ($\bar{Z}$), $\alpha=0.8$. Square colors are indexed on the mean AvgCost difference. For example, Calimera has a lower mean AvgCost when trained over non $z$-normalized datasets by 1.3e-2 and appears in light blue. It beats the $z$-normalized version over 25 datasets, loses over 9 and are tied on 1. The Wilcoxon p-value is equal to 0.2224, which is higher than significance level equal to 0.05. Thus, no statistical difference can be observed for the considered approach.
Figure 15: Multi-comparison-matrices (?). The upper triangle, with dark blue contours, displays the comparison of the competitive methods trained over non z-normalized dataset (orange cylinder). The values within this triangle has to be read by lines, i.e. for a considered line, red shades indicate better performances, blue shades weaker performances. The lower triangle, with dark red contours, is the comparison of the methods trained over the same datasets, z-normalized (chocolate cylinder). The values within this triangle has to be read by columns, i.e. for a considered column, red shades indicate better performances, blue shades weaker performances. The complete figure being symmetrical indicates that z-normalization does not impact much relative ranking between methods.
B.6 Anomaly detection cost setting : an ablation study
Exponential delay cost only
Figure 16: Exponential delay cost, symmetric binary misclassification cost, non z-normalized proposed imbalanced datasets (orange cylinder with a whole).
Imbalanced misclassification cost only
Figure 17: Linear delay cost, non symmetric imbalanced misclassification cost, non z-normalized proposed imbalanced datasets (orange cylinder with a whole).
Standard cost setting, imbalanced datasets
Figure 18: Linear delay cost, symmetric binary misclassification cost, non z-normalized proposed imbalanced datasets (orange cylinder with a whole).

### v6

@@
-Early Classification of Time Series: Taxonomy of Methods and Extensive Benchmark
+Early Classification of Time Series: A Survey and Benchmark
 Abstract
@@
 Although it has been the subject of a growing body of literature, there is still a lack of a systematic, shared evaluation protocol to compare the relative merits of the various existing methods.
+In this paper, we highlight the two components of an ECTS system: decision and prediction, and focus on the approaches that separate them.
 This document begins by situating these methods within a principle-based taxonomy.
-It defines dimensions for organizing their evaluation, and then reports the results of a very extensive set of experiments along these dimensions involving nine state-of-the art ECTS algorithms.
+It defines dimensions for organizing their evaluation and then reports the results of a very extensive set of experiments along these dimensions involving nine state-of-the-art ECTS algorithms.
 In addition, these and other experiments can be carried out using an open-source library in which most of the existing ECTS algorithms have been implemented (see https://github.com/ML-EDM/ml_edm).
 1 Introduction
-In hospital emergency rooms (?), in the control rooms of national or international power grids (?), in government councils assessing critical situations, there is a time pressure to make early decisions.
+In hospital emergency rooms (Mathukia et al., 2015), in the control rooms of national or international power grids (Dachraoui et al., 2015), in government councils assessing critical situations, in many situations there is a time pressure to make early decisions.
 On the one hand, the longer a decision is delayed, the lower the risk of making the wrong decision, as knowledge of the problem increases with time.
 On the other hand, late decisions are generally more costly, if only because early decisions allow one to be better prepared.
-For example, a cyber-attack that is not detected quickly enough gives hackers time to exploit the security flaw found.
-A number of applications involve making decisions that optimizes a trade-off between accuracy of the prediction and its earliness.
+A number of applications thus involve making decisions that optimize a trade-off between the accuracy of the prediction and its earliness.
 The problem is that favoring one usually works against the other.
 Greater accuracy comes at the price of waiting for more data.
-Such a compromise between the Earliness and the Accuracy of decisions has been particularly studied in the field of Early Classification of Time Series (ECTS) (?), and introduced by ?
-(?).
+Such a compromise between the Earliness and the Accuracy of decisions has been particularly studied in the field of Early Classification of Time Series (ECTS), and introduced by Xing et al.
+(2008).
 ECTS consists in finding the optimal time to trigger the class prediction of an input time series observed over time.
-Successive measurements provide more and more information about the incoming time series, and ECTS algorithms aim to optimize online the trade-off between two conflicting objectives, namely, the earliness and accuracy of class predictions.
-More formally, we have the following problem.
-Problem statement:
-In the ECTS problem, measurements of an input time series are observed over time.
-At time $t$, the incomplete time series ${\mathbf{x}}_{t}\,=\,\langle{x_{1}},\ldots,{x_{t}}\rangle$ is available where ${x_{i}}_{(1\leq i\leq t)}$ denotes the time indexed measurements.
+As pointed out by Bondu et al.
+(2022), the ECTS problem is rooted in optimal stopping, of which it is a special case (Shepp, 1969; Ferguson, 1989), where the decision to be made concerns both: (i) when to stop receiving new measurements in order to (ii) predict the class of the incoming time series.
+To the best of our knowledge, Alonso González & Diez (2004) are the earliest explicitly mentioning “classification when only part of the series are presented to the classifier”.
+Since then, several researchers have continued their efforts in this direction and have published a large number of research articles (Xing et al., 2009; Anderson et al., 2012; Dachraoui et al., 2015; Mori et al., 2017a; Schäfer & Leser, 2020; Lv & Hu, 2022; Ebihara et al., 2025).
+However, despite the growing interest in ECTS over the last twenty years (Gupta et al., 2020; Akasiadis et al., 2024; Santos & Kern, 2016), there still remains a need for a shared taxonomy of approaches and an agreed well-grounded evaluation methodology (see Table 3 in Appendix A).
+Guidelines that we feel are important for a fair and informative comparison of existing ECTS approaches are listed below:
+  1. 1.
+    Costs taken into account for evaluating the performance of the proposed method should be explicitly stated.
+It seems natural to distinguish between the misclassification costs, and the delay cost, and to add them in order to define the cost of making a decision at time $t$.
+The delay cost may also depend on the true class $y$ and the predicted one $\hat{y}$, and a single cost function integrating misclassification and delay costs should then be used.
+For the sake of clarity, we keep the simple notation that distinguishes both cost functions in the rest of this paper.
+  2. 2.
+    The performance of the proposed methods should be evaluated against a range of possible types of cost functions.
+It is usual to evaluate “by default” the methods using a $\ell_{0-1}$ loss function that penalizes a wrong classification by a unity cost, and to consider a linear delay cost function.
+However, lots of applications rather involve unbalanced misclassification costs, and possibly also non-linear delay costs.
+  3. 3.
+    The contributions of the various components of an ECTS algorithm should be as clearly delineated as possible.
+The predominant approach to ECTS is to have a decision component which is in charge of evaluating the best moment to make the prediction about the class of the incoming time series, and a classifier one which makes the prediction itself.
+We call these methods “separable methods”.
+In order to fairly compare the triggering methods at play, which are at the heart of ECTS, the classifier used should be the same for all methods in the experiments.
+    Recent approaches mix the decision and the prediction components.
+This is the case of “end-to-end” methods using neural networks.
+These methods do not allow the evaluation of the merits of the decision component by itself, independently of the classification component, and they are accordingly not considered here.
+  4. 4.
+    Performance obtained should be compared with that of “baseline” algorithms.
+Failing to do this weakens any claim about the value of the proposed method.
+In the case of ECTS tasks, two naive baselines are: (1) make a prediction as soon as it is allowed, and (2) make a prediction after the entire time series has been observed.
+In our experiments reported in Section 4, we have added a third baseline less simple than the two aforementioned, which, to our knowledge has never been published as an original method.
+This is a confidence-based method where a decision is triggered as soon as the confidence for the likeliest class given $\mathbf{x}_{t}$ is greater than a threshold.
+  5. 5.
+    The datasets used for training and testing should remove the biases that are specific to the ECTS task and that may result in erroneous evaluations.
+A case in point, concerns the normalization often used in time series datasets.
+Dau et al.
+(2019) have reported that 71% of the reference time series classification datasets used to evaluate ECTS methods are made up of z-normalized time series, Not only, this setting is not applicable in practice, as z-normalization would require knowledge of the entire incoming time series, but it could also introduce a leakage of information from the future of the time series to the previous time steps which is detrimental to a fair assessment of the methods (Wu et al., 2021).
+In this article, we aim to present a methodology for a fair and informative comparison between separable ECTS methods in the literature. Specifically, this paper makes the following contributions:
+  • A taxonomy is proposed in Section 2, classifying separable approaches in the literature according to their design choices.
+  • Extensive experiments have been performed, taking into account the obstacles mentioned above.
+(1) The experimental protocol in Section 4.1 explicitly defines the costs used during training and evaluation, and varies the balance between misclassification and delay costs by using a large range of cost values.
+(2) Experiments are performed repeatedly for several types of cost function, i.e.
+balanced or unbalanced misclassification costs, and linear or exponential delay costs (see Sections 4.2 and 4.3) and many intermediate results are available in the supplementary materials.
+(3) Ablation and substitution studies are conducted in Section 4.4 with the aim of evaluating the impact of methodological choices, such as the choice of classifier, its calibration, or even z-normalization of training time series, as well as the non-myopia property of some trigger functions.
+(4) The experiments include three baseline approaches, rarely considered in the literature, which often prove to be efficient.
+(5) In addition to the reference data used in the ECTS field, a collection of some thirty non-z-normalized datasets is proposed and provided to the community.
+  • An open source library is made available111(see https://github.com/ML-EDM/ml_edm) to enable reproducible experiments, as well as to facilitate the scientific community’s development of future approaches.
+Particular care has been taken to ensure the quality of the code, so that this library may be used to develop real-life applications (see Appendix D).
+The scope of this paper is a survey and a benchmark of separable ECTS approaches, in line with our guideline #3.
+Interested readers can refer to Appendix C for complementary results on end-to-end recent approaches.
+The rest of this paper is organized as follows: Section 2 proposes and describes an ECTS taxonomy in order to situate separable approaches in relation to each other.
+It underlines the different choices to be made in a well-founded way when designing an ECTS method.
+Section 3 presents a review of state-of-the-art methods for ECTS organized along the dimensions of the suggested taxonomy.
+In Section 4, we present the pipeline developed to conduct extensive experimentation, and we report the main results obtained for different cost settings.
+Finally, Section 5 concludes this paper.
+Position with respect to other literature surveys
+In reaction to the abundant scientific activity around ECTS, Gupta et al.
+(2020) wrote a survey, however with the following characteristics.
+First, their perspective on ECTS is confidence-based as they declare “A primary task of an early classification approach is to classify an incomplete time series as soon as possible with some desired level of accuracy” (italics are ours) and it is not centered on optimizing the tradeoff between predictive accuracy and earliness.
+Second, they organize the covered approaches by application domains and methods for representing time series, viz.
+prefix-based, shapelet-based, model-based, and miscellaneous.
+A perspective that, in a way, is tangential to the problem of selecting the right decision time.
+Another paper (Akasiadis et al., 2024) intends to propose a framework to evaluate ECTS systems.
+Noticeably, it also adopts a confidence-based approach: “The objective is to find the earliest time-point of a time-series at which a reliable prediction can be made”.
+Alongside the survey, a benchmark is proposed whose characteristics are detailed in Table 3.
+The earliest survey published in 2016 by Santos & Kern (2016) deals mostly with classification methods for complete time series. The final section describes four ECTS algorithms with no experiments.
+2 ECTS: concepts and taxonomy
+The aim of this section is (i) to formally describe the ECTS problem, and (ii) to outline in a principled way the various choices that need to be made when designing one ECTS method.
+Problem statement
+In the ECTS problem, an input time series of size $T$ is progressively observed over time.
+At time $t\leq T$, the incomplete time series ${\mathbf{x}}_{t}\,=\,\langle{x_{1}},\ldots,{x_{t}}\rangle$ is available where ${x_{i}}_{(1\leq i\leq t)}$ denotes the time-indexed measurements.
 These measurements can be single or multi-valued.
 The input time series belongs to an unknown class $y\in{\cal Y}$.
-The task is to make a prediction $\hat{y}\in{\cal Y}$ about the class of the incoming time series, at a time $\hat{t}\in[1,T]$ before the deadline $T$ which corresponds to the length of the full time series.
-A misclassification cost is incurred when a prediction is made, denoted by $\mathrm{C}_{m}(\hat{y}|y)$.
-Furthermore, there exists a delay cost $\mathrm{C}_{d}(\hat{t})$ that expresses the time pressure and encourages early decisions (defined in Section 2.2).
-The choice of the best trigger time should optimize a compromise between these costs that move in opposite direction.
-The choice of the best triggering time must optimize a compromise between the two costs, which are moving in opposite directions.
-To the best of our knowledge, ?
-(?) are the earliest explicitly mentioning “classification when only part of the series are presented to the classifier”.
-Since then, several researchers have continued their efforts in this direction and have published a large number of research articles.
-A recent and extensive review of the ETCS approaches can be found in the paper written by ?
-(?), including the applications that motivated the researchers to work in this area, and covering about fifty relevant papers selected from the $213$ papers found by search engines at the time of this writing.
-As pointed out by ?
-(?), the ECTS problem is a special case of optimal stopping (?, ?), where the decision to be made concerns both: (i) when to stop receiving new measurements in order to (ii) predict the class of the incoming time series.
-In the same paper, the ECTS problem has been extended into a more general one, which consists in optimizing the decision times of Machine Learning (ML) models in a wide range of settings where data is collected over time.
-The authors proposed a set of open research questions to the community, in order to widen the range of applications that are amenable to the ECTS framework (i.e.
-dealing with other learning tasks, other types of data, other application contexts etc.).
-However, despite the growing interest in ECTS over the last twenty years, there still remains a need for a shared taxonomy of approaches and an agreed well-grounded evaluation methodology.
-Here, in particular, we list limits that hamper a fair comparison of ECTS methods and algorithms:
-  1. 1.
-    Costs taken into account for evaluating the performance of the proposed method are not always clearly stated.
-It seems natural to distinguish between the misclassification costs $\mathrm{C}_{m}(\hat{y}|y)$, and the delay cost $\mathrm{C}_{d}(t)$, and to add them in order to define the cost of making a decision at time $t$.
-More generally, the delay cost may depend on the true class $y$ and the predicted one $\hat{y}$, and a single cost function $\mathrm{C}_{m}(\hat{y}|y,t)$ integrating misclassification and delay costs should then be used.
-For the sake of clarity, we keep the simple notation which distinguishes both cost functions in the rest of this paper.
-But in all cases, it is essential to state the framework used and the associated evaluation metric.
-  2. 2.
-    The performance of the proposed methods should be evaluated against a range of possible types of cost functions.
-It is usual to evaluate “by default” the methods using a $\ell_{0-1}$ loss function that penalizes a wrong classification by a unity cost, and to consider a linear delay cost function: $\mathrm{C}_{d}(t)=\lambda\,t$, for a value $\lambda>0$.
-However, lots of applications rather involve unbalanced miss-classification costs, and possibly also non linear delay costs.
-This is the case, for instance, in maintenance applications where wrongly not recognizing a critical situation is much more costly than wrongly predicting a problem and taking steps to fix it, and where delay cost may rise as a an exponential function of time: $\mathrm{C}_{d}(t)=\lambda\,e^{t}$ It is therefore quite important to assess the adaptability of the methods to various representative problem settings.
-  3. 3.
-    The contributions of the various components of a ECTS algorithm should be clearly delineated.
-The predominant approach to ECTS is to have a decision component which is in charge of evaluating the best moment to make the prediction about the class of the incoming times series, and a classifier one which makes the prediction itself.
-In order to fairly compare the triggering methods, which are at the heart of ECTS, the classifier used should be the same.
-We call these methods “separable methods”.
-    An alternative approach relies on having a system that classifies the incoming time series $\mathbf{x}_{t}$ at each time with a prediction $\hat{y}$ in the set {‘postpone decision’, $y_{1},\ldots,y_{N}$} where $N$ is the number of classes.
-Therefore, within this approach, a single system decides either to wait at least one more time step, or to predict a class and stops the process.
-In this case, no distinction can be made between a decision component and a classifier one, and the whole system is evaluated as such.
-In the spirit of deep neural networks, we call this type of methods “end-to-end” to underline the fact that a single learning system is in charge of all operations, here decision and classification.
-Of course, this precludes a comparison involving only the choice of the decision component with other methods.
-  4. 4.
-    As with other supervised learning tasks, performance should be compared with that of “baseline” algorithms.
-In the case of ECTS tasks, two naive baselines are: (1) make a prediction as soon as it is allowed, and (2) make a prediction at $T$, after the entire time series has been observed.
-In our experiments reported in Section 4, we have added a third baseline, less simple than the two afore-mentioned ones, but still too obvious so that, to our knowledge, it has never been published as an original method.
-This is a confidence-based method where a decision is triggered as soon as the confidence for the likeliest class given $\mathbf{x}_{t}$ is greater than a threshold.
-(Formally, let $\hat{y}=\operatorname*{arg\,max}_{y\in\mathcal{Y}}p(y|\mathbf{x}_{t})$, then a prediction is made (i.e.
-$\hat{y}$) as soon as $p(\hat{y}|\mathbf{x}_{t})\geq\varepsilon$, for some threshold $\varepsilon\in[0,1]$.)
-  5. 5.
-    Precautions should be taken when using datasets of training time series to ensure that no bias enters unwillingly into the training and evaluation process.
-A case in point, concerns the normalization often used in time series datasets.
-?
-(?) have reported that 71% of the reference time series classification datasets used to evaluate ECTS methods are made up of z-normalized time series, i.e.
-with measurements independently modified on each complete series to obtain a mean of zero and a standard deviation equal to 1.
-Clearly, this setting is not applicable in practice, as z-normalization would require knowledge of the entire incoming time series.
-In a research context, previous work has used such training sets to test the proposed algorithms.
-As ?
-(?, ?) note, this preprocessing is irreversible and can generate a problem for ECTS by introducing a temporal leakage of information.
-In order to assess its impact, we report in Section B.5 of Appendix B a comparison of results for z-normalized and non-normalized time series.
-Up until now, it has been difficult to conduct fair comparisons between competing methods.
-Often, published performances are based on choices concerning data sets, the precise performance measure used, hyperparameter values and evaluation protocols (e.g., the split between training and test sets) that are not entirely explicit or, in any case, are difficult to reproduce.
-This is why we have recoded all the methods, specified a shared evaluation protocol with variants that can be employed by everyone, and searched for a collection of data sets that can be widely used to test and compare new as well as existing methods.
-We hope this will be a useful resource for the scientific community working in this field.
-As this community shows a growing interest in the ECTS problem, granted by the increasing number of applications that fall in its range, it is timely (1) to propose a framework into which to cast the various approaches and thus indicate avenues for future research, and (2) a well-grounded evaluation methodology.
-Specifically, this paper makes the following contributions:
-  • A taxonomy is proposed in Section 2, classifying approaches in the literature according to their design choices.
-  • Extensive experiments have been performed, meeting the above-mentioned shortcomings.
-(1) The experimental protocol in Section 4.1 explicitly defines the costs used during training and evaluation, and varies the balance between misclassification and delay costs by using a large range of cost values.
-(2) Experiments are performed repeatedly for several types of cost function, i.e.
-balanced or unbalanced misclassification cost, and linear or exponential delay cost (see Sections 4.2 and 4.3) and many intermediate results are available in the supplementary materials.
-(3) Ablation and substitution studies are conducted in Section 4.4 with the aim of evaluating the impact of methodological choices, such as the choice of classifier, its calibration, or even z-normalization of training time series.
-(4) The experiments include three baseline approaches, rarely considered in the literature, which often give surprisingly good results.
-(5) In addition to the reference data used in the ECTS field, a collection of some thirty non-z-normalized datasets is proposed and provided to the community.
-  • An open source library is being made available [^1] https://github.com/ML-EDM/ml_edm in order to enable reproducible experiments, as well as facilitate the scientific community’s development of future approaches.
-Particular care has been taken to ensure the quality of the code, so that this library may be used to develop real-life applications.
-The rest of this paper is organised as follows : The section 2 proposes and describes a new ECTS taxonomy, the different choices to be made in a well-founded way when designing an ECTS method and a set of four questions that need to be answered to make these choices.
-Section 3 presents a comprehensive view of the ECTS field, along the suggested taxonomy and the four questions raised in Section 2.
-In Section 4 we present the pipeline developed in order to realize extensive experimentation and we report the main results obtained for different cost setting.
-This benchmark is supported by a library released as Open Source for dissemination and used in the ECTS research community.
-Finally Section 5 concludes this paper.
-Appendix A lists the datasets used for the experiments, and complementary results are provided in Appendix B.
-2 Organizing the ECTS approaches: a taxonomy
-The aim of this section is to outline in a principled way the various choices that need to be made when designing one ECTS method.
-General form of an ECTS model
+The task is to make a prediction $\hat{y}\in{\cal Y}$ about the class of the incoming time series, at a time $\hat{t}\in[1,T]$ before the deadline $T$.
 An ECTS approach aims at optimizing a trade-off between accuracy and earliness of the prediction, and thus must be evaluated on this ground.
@@
 We thus consider:
-  • $\mathrm{C}_{m}(\hat{y}|y):{\cal Y}\times{\cal Y}\rightarrow\mathbb{R}$, that corresponds to the misclassification cost of predicting $\hat{y}$ when the true class is $y$.
+  • $\mathrm{C}_{m}(\hat{y}|y):{\cal Y}\times{\cal Y}\rightarrow\mathbb{R}$, that corresponds to the cost of predicting $\hat{y}$ when the true class is $y$.
   • $\mathrm{C}_{d}(t):\mathbb{R}^{+}\rightarrow\mathbb{R}$, the delay cost that, usually, is a non-decreasing function over time.
@@
 The best decision time $t^{*}$ is given by:
-$t^{\star}=\operatorname*{arg\,min}_{t\in[1,T]}\mathcal{L}(\hat{y}({\mathbf{x}}_{t}),t,y).$ (1)
+$t^{\star}=\operatorname*{arg\,min}_{t\in[1,T]}\mathcal{L}(\hat{y}({\mathbf{x}}_{t}),y,t).$ (1)
 Let $s^{\star}\in\mathcal{S}$ an optimal ECTS function belonging to a class of functions $\mathcal{S}$, whose output at time $t$ when receiving $\mathbf{x}_{t}$ is:
-$s^{\star}({\mathbf{x}}_{t})=\left\{\begin{array}[]{ll}\emptyset&\mbox{if extrameasures are queried;}\\ y^{\star}=\hat{y}(x_{t^{\star}})&\mbox{when prediction is triggered at $t=t^{\star}$;}\end{array}\right.$ (2)
-ECTS is however an online optimization problem, where at each time step $t$ a function $s({\mathbf{x}}_{t})$ must decide to make a prediction or not.
-Equation 1 is thus no longer operational since it requires a complete knowledge of the time series.
-In practice, the function $s({\mathbf{x}}_{t})$ triggers a decision at $\hat{t}$, based on a partial description ${\mathbf{x}}_{\hat{t}}$ of the incoming time series ${\mathbf{x}}_{T}$ (with $t\leq T$).
-The goal of an ECTS system is to chose a triggering time $\hat{t}$ as close as possible to the optimal one $t^{*}$, at least in terms of cost, minimizing $\mathcal{L}(\hat{y}({\mathbf{x}}_{\hat{t}}),\hat{t},y)-\mathcal{L}(\hat{y}({\mathbf{x}}_{t^{\star}}),t^{\star},y)$ as much as possible.
-From a machine learning point of view, the goal is to find a function $s\in{\cal S}$ that best optimizes the loss function $\mathcal{L}$, minimizing the true risk over all time series distributed according to the distribution222 Notice that the notation $\mathcal{X}$ is an abuse that we use use to simplify our purpose.
+[…]
+ECTS is however an online optimization problem, where at each time step $t$ a function $s({\mathbf{x}}_{t})$ must decide whether to make a prediction or not.
+Equation 1 is thus no longer operational since it requires complete knowledge of the time series.
+In practice, the function $s({\mathbf{x}}_{t})$ triggers a decision at $\hat{t}$, based on a partial description $^{\mathbf{x}}_{\hat{t}}$ of the incoming time series ${\mathbf{x}}_{T}$ (with $\hat{t}\leq T$).
+The goal of an ECTS system is to choose a triggering time $\hat{t}$ as close as possible to the optimal one $t^{*}$, at least in terms of cost, minimizing $\mathcal{L}(\hat{y}({\mathbf{x}}_{\hat{t}}),y,\hat{t})-\mathcal{L}(\hat{y}({\mathbf{x}}_{t^{\star}}),y,t^{\star})$ as much as possible.
+From a machine learning point of view, the goal is to find a function $s\in{\cal S}$ that best optimizes the loss function $\mathcal{L}$, minimizing the true risk over all time series distributed according to the distribution [^2] Notice that the notation $\mathcal{X}$ is an abuse that we use use to simplify our purpose.
 In all mathematical rigor, the measurements observed successively constitute a family of time-indexed random variables $\mathbf{x}=(\mathbf{x}_{t})_{t\in[1,T]}$.
-This stochastic process $\mathbf{x}$ is not generated as commonly by a distribution, but by a filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[1,T]}$ which is defined as a collection of nested $\sigma$-algebras (?) allowing to consider time dependencies.
+This stochastic process $\mathbf{x}$ is not generated as commonly by a distribution, but by a filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[1,T]}$ which is defined as a collection of nested $\sigma$-algebras (Klenke, 2013) allowing to consider time dependencies.
 Therefore, the distribution $\mathcal{X}$ should also be re-written as a filtration.
-$\mathbb{P}_{\mathcal{X}}$ that governs the time series in the application:
-$\small\normalsize\operatorname*{arg\,min}_{s\in\mathcal{S}}\operatornamewithlimits{\mathbb{E}_{\>\mathbf{x}\sim\mathbb{P}_{\mathcal{X}}}^{\>}}\left[\mathcal{L}(\hat{y}({\mathbf{x}}_{\hat{t}}),\hat{t},y)\right]$ (3)
-The questions then are:
-  1. 1.
-    Which form can take the function $s(\cdot)$? We will distinguish end-to-end architecture from separable ones.
-  2. 2.
-    How the criterion to be optimized accounts for the trade-off between accuracy and earliness?
-We will see that the costs implied, about misclassification and delay, are variously explicit in the existing methods.
-  3. 3.
-    How the when question of the stopping problem can be approached?
-This will lead us to distinguish between cost-informed and cost-uninformed ones on one hand, and between anticipation-based methods versus myopic ones, on the other hand.
-  4. 4.
-    How the prediction problem itself can be solved given that $\mathbf{x}_{t}$ belongs to a different input spaces at each time step $t$?
-This set of questions and possible choices for their solution are illustrated in Figure 1. We turn successively to each one in the following.
-Figure 1: Proposed ECTS taxonomy
-2.1 The different forms of the function $s(\cdot)$
+$\mathbb{P}_{(\mathcal{X}\times\mathcal{Y})}$ that governs the time series in the application:
+$\operatorname*{arg\,min}_{s\in\mathcal{S}}\mathbb{E}_{(\mathbf{x},y)\sim\mathbb{P}_{(\mathcal{X}\times\mathcal{Y})}}\left[{\mathcal{L}(\hat{y}({\mathbf{x}}_{\hat{t}}),y,\hat{t})}\right$] (3)
+In order to solve the ECTS problem, a training set composed of $M$ labeled time series, denoted by $({\mathbf{x}}^{i}_{T},y^{i})_{i\in[0,M]}\in(\mathcal{X}\times\mathcal{Y})$, where each series ${\mathbf{x}}_{T}=\,\langle{x_{1}},\ldots,{x_{T}}\rangle$ is complete and of the same size $T$, and associated with its label $y\in\mathcal{Y}$, is used to learn how to predict the class of an incoming time series ${\mathbf{x}}_{t}\,=\,\langle{x_{1}},\ldots,{x_{t}}\rangle$, and to learn what can be expected in the future given ${\mathbf{x}}_{t}$.
+Consequently, model training and deployment are of different natures.
+The training stage is carried out as a supervised batch process, with access to the full labeled time series.
+When it comes to testing, on the other hand, decision-triggering is an online process which stops at time $\hat{t}$, and at the latest, when the deadline $T$ is reached.
+2.1 Separable vs. end-to-end approaches
 An ECTS function must solve both the question of (i) when to stop receiving new measurements and decide to make a prediction and (ii) how to make the prediction about the class of the incoming time series $\mathbf{x}_{t}$.
@@
 Within this perspective, the classification component is learned independently of the trigger one, while the latter uses the results of the classification component in order to trigger a decision.
-A simple triggering strategy is to decide it is time to make a prediction as soon as the classification component is sufficiently sure of its prediction.
-We formalize separable approaches by: $s(\mathbf{x}_{t})=(g\circ h)(\mathbf{x}_{t})$ where $g$ is the decision or trigger function, and $h$ is the prediction function.
+We formalize separable approaches by:
+$s(\mathbf{x}_{t})=(g\circ h)(\mathbf{x}_{t})$ (4)
+where $g$ is the decision or trigger function, and $h$ is the prediction function.
 In the end-to-end approaches, a single component decides when to make a prediction and what that prediction is.
-Thus, the function $s$, defined in Equation 2, is responsible both for choosing the time $\hat{t}$ for making the prediction, and for the prediction itself $\hat{y}$.
+Thus, the function $s$, defined in Equation 2, is responsible both for choosing the time $\hat{t}$ for making the predictions, and for the prediction itself $\hat{y}$.
 The question that naturally arises is which type of architecture (i.e.
 end-to-end or separable) performs best.
-On the one hand, in separable approaches, the classification component is trained independently of the triggering one, which can be detrimental.
-On the other hand, separating the ECTS problem into two inherently simpler sub-problems could be an advantage.
-In this paper, we do not delve any further into this question, which we leave for future work.
-2.2 Choice of the optimizing criterion during training
-This section covers optimization criteria existing in the literature and used to train ECTS functions.
-In the following, these criteria are listed by level of cost awareness, and we differentiate between the following situations:
-  1. 1.
-    The first one, we call cost-informed at training time.
-    $\mathbb{P}_{\mathcal{X}}$ being unknown, instead of using Equation 3 describing the true risk, one tries to minimize the empirical risk, also called average cost in the ECTS literature, for a training set of $M$ time series:
-    $AvgCost\;=\frac{1}{M}\sum_{i=1}^{M}\mathcal{L}(\hat{y}_{i},y_{i},\hat{t})\;=\;\frac{1}{M}\sum_{i=1}^{M}\mathrm{C}_{m}(\hat{y}_{i}|y_{i})+\mathrm{C}_{d}(\hat{t}_{i})$ (4)
-    AvgCost is the most appropriate criterion to both train and evaluate ECTS approches. The following presents proxy measures of AvgCost.
-  2. 2.
-    The second situation is a sub-case of cost-informed, and can be qualified as cost-proxy at training time.
-There, while the accuracy and earliness of prediction are taken into account, the optimization criterion combines them in a proxy which is not the AvgCost.
-Classical proxys include:
-    • The Harmonic Mean (see ?):
-      $HM=\frac{2\times Accuracy\times(1-Earliness)}{Accuracy+(1-Earliness)}$ (5)
-      with:
-      $Accuracy$ $=\frac{1}{M}\sum_{i=1}^{M}\mathbb{1}(\hat{y}_{i}=y_{i})$ (6)
-      $Earliness$ $=\frac{1}{M\times T}\sum_{i=1}^{M}\hat{t}_{i}$ (7)
-      where $\hat{t}_{i}$ is the time that the ECTS function decides to make prediction for the time series $\mathbf{x}^{i}$.
-    • The CF (i.e. Cost Function) criterion with $0\leq\alpha\leq 1$ (see ?, ?, ?):
-      $CF=\alpha\times(1-Accuracy)+(1-\alpha)\times Earliness$ (8)
-      When the cost of misclassification $\mathrm{C}_{m}(\hat{y},y)=\mathbb{1}(\hat{y}=y)$, and the delay cost is $\mathrm{C}_{d}(t)=\hat{t}/T$ and $\alpha=0.5$, CF becomes a particular case of AvgCost.
-    The choice of using costs or approximating them by a proxy is a technical issue, which may, for example, be relevant to making the loss function differentiable by approximation.
-For this reason, in the remainder of this paper, cost-informed/cost-proxys at training time situations are grouped together under the term cost-informed at training time, this difference not being essential.
-  3. 3.
-    The third situation is qualified as cost-uninformed-train.
-This is the case of methods that only set the threshold on the confidence of the prediction in order to trigger the prediction, regardless of the costs incurred on the training set.
-Another example is methods that use rigid rules, such as “decide as early as the first measurement is available”.
-It is a question whether any of these approaches fare better than the others.
-To measure this, we must use the Average Cost (see Equation 4) measured on a test set, which represents the ground truth of the ECTS problem.
-It should be noted that many of the proposed methods have been evaluated on the basis of other criteria in the literature, with the result of not allowing a rigorous comparison between them.
-We will come to this problem in Section 4.
+On the one hand, in separable approaches, the classification component is trained independently of the triggering one, which can be detrimental, for example, by propagating errors from one module to another.
+On the other hand, separating the ECTS problem into two inherently simpler sub-problems could be an advantage, for example, in terms of convergence during training.
+Additionally, the separable framework allows one to directly leverage the latest advances in the Time Series Classification (TSC) literature (Bagnall et al., 2017).
+In this paper, we do not delve any further into the question of which type of architecture is best, which we leave for future work.
 The rest of this section is specific to separable ECTS approaches, which represent a large part of the literature.
-2.3 Information used by the trigger function during inference
-The trigger function can draw on different types of information.
-In the simplest case, it can decide irrespectively of the incoming time series $\mathbf{x}_{t}$.
-This is the case, for example, of the rule that would say: “wait until half of the measurements are available, then make a prediction”.
-The corresponding trigger function can be said to be blind.
-Apart from this extreme case, it is interesting to distinguish two dimensions.
-First, the trigger function may or may not take the misclassification and delay costs into account.
-Second, it can also make its decision at each time step solely on the basis of past observations, or it can anticipate possible futures to help its decision.
-For instance, confidence-based methods (seen Section 2.3.1) do not explicitly take costs into account in the triggering decision.
+In the following, we first examine the decision component (i.e.
+the trigger function) and then the prediction one (i.e.
+the classifier).
+Figure 1 gives a general view of the proposed taxonomy.
+2.2 Trigger function’s properties
+[…]
+Figure 1: Proposed taxonomy for separable ECTS approaches, which can be viewed as a function composition of trigger and classification functions.
+The taxonomy does not follow a tree structure as the trigger function may have several properties, that are not mutually exclusive.
+For instance, a trigger function can be both confidence-based and anticipation based.
+2.2.1 Cost-informed or cost-uniformed
+The first property is whether or not the trigger function actually takes as input any cost functions (or eventually some proxy of them) during training.
+We call these approaches cost-informed.
+On the contrary, cost-uninformed approaches are usually based on some hard pre-defined rules to trigger and are not always easily expandable to a generic cost-setting framework.
+Additionally, when being cost-informed, trigger functions can leverage cost information even more, by being adaptable at testing time, i.e.
+during inference, trigger decisions directly depend on cost functions.
 2.2.2 Confidence-based approaches
-Confidence-based approaches are widely used in the literature.
 The simplest trigger model of this kind consists in monitoring a quantity related to the confidence of the prediction over time and triggering class prediction as soon as a threshold value is exceeded.
 The confidence metric monitored can take different forms.
-For example, a baseline approach, referred to as Proba Threshold [^3] Baseline implemented in the aeon (?) library : https://urlz.fr/qmWl in the remainder of this paper, involves monitoring $\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})$ the highest conditional probability estimated by the classifier.
+For example, a baseline approach, referred to as Proba Threshold [^3] Baseline implemented in the aeon (Middlehurst et al., 2024a) library : https://urlz.fr/qmWl in the remainder of this paper, involves monitoring $\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})$ the highest conditional probability estimated by the classifier.
 This baseline example is qualified as instant-based method, since it takes as input only the last confidence score available at time $t$.
-Another type of approaches, qualified as sequence-based, monitors the entire sequence of past confidence scores, and triggering a prediction is made conditionally on a particular property of this sequence.
+Another type of approaches, qualified as sequence-based, monitors the entire sequence of past confidence scores, and a triggering prediction is made conditionally on a particular property of this sequence.
 Accordingly, trigger functions can either take as input a scalar value, e.g.
-$g(\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t}))$, in the case of instant based approaches, or a sequence of scalar values, e.g.
+$g(\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t}))$, in the case of instant-based approaches, or a sequence of scalar values, e.g.
 $g(\{\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{\tau})\}_{1\leq\tau\leq t})$, in the case of sequence-based approaches (see Section 3.1).
-2.3.2 cost-informed at testing time
-Given that an ECTS approach will ultimately be evaluated on the average cost of using it (see Equation 4), it seems natural to exploit the cost values at testing time, in order to trigger predictions at optimal moments.
-Methods such as Economy (?) and 2step/NoCluster (?) do that.
-They can thus be qualified as “cost-informed at testing time”.
-Other approaches use instead trigger functions that, once learned, do not take into account the cost at testing time, but rely on other measures such as, for instance, the confidence of the prediction.
-This is the case of the SR approach (?).
-Therefore, these approaches can be qualified as “cost-uninformed at testing time”.
-Notice that some approaches are cost-informed during training but not during inference.
-This is the case with the SR approach, which is cost-informed at training time since it uses costs to optimize its parameters, and cost-uninformed at testing time since the resulting trigger function does not use cost values during inference.
-Table 1 shows these two different properties for each approach in the literature.
-2.3.3 Blind, Myopic and Anticipation-based decisions
-Some separable approaches consider the output of the classifier $h$ at time $t$ to decide whether this is the right time to make a prediction.
-For instance, stopping the process when the confidence in the classification $h_{t}(\mathbf{x}_{t})$ is above some threshold, or when the difference of confidence between the best two predictions exceeds some value.
-These methods can be described as myopic since they only look at the current time step $t$, without trying to guess the future.
-But there is a another possibility. As was first noted by ? (?), the ECTS problem can be cast as a LUPI (Learning Using Privileged Information) problem (?).
-In this scenario, the learner can benefit at the training time of privileged information that will not be available at test time.
+Those kind of examples can be described as myopic since they only look at the current time step $t$ (or past ones as well in the sequence-based case), without trying to anticipate the likely future, i.e.
+not directly having a decision made based on a forecast of some form.
+The anticipation-based approaches do this.
+2.2.3 Anticipation-based decisions
+As was first noted by Achenchabe et al.
+(2021), the ECTS problem can be cast as a LUPI (Learning Using Privileged Information) problem (Vapnik & Vashist, 2009).
+In this scenario, the learner can benefit at the training time from privileged information that will not be available at test time.
 Formally, the training set can be expressed as $\mathcal{T}=\{(\mathbf{x}_{i},\mathbf{x}_{i}^{\star},y_{i})\}$, where $\mathbf{x}_{i}$ is what is observable and $\mathbf{x}_{i}^{\star}$ is some additional information not available when the prediction must be made.
 This is exactly what happens in the ECTS problem.
-Whereas at test time, only $\mathbf{x}_{t}$ is available, during training the complete time series are known.
-This brings the possibility to learn what are the likely future of an incoming time series $\mathbf{x}_{t}$ provided it comes from the same distribution.
+Whereas at test time, only $\mathbf{x}_{t}$ is available, during training the complete time series is known.
+This brings the possibility to learn what are the likely futures of an incoming time series $\mathbf{x}_{t}$ provided it comes from the same distribution.
 Hence, it becomes also possible to guess the cost to be optimized for all future time steps, and therefore to wait until the moment seems the best.
 This type of approach can be said anticipation-based (also called non-myopic in the literature).
-Because more information from the training set is exploited, it can be expected that these methods outperform myopic and blind ones.
+They can come with many flavors, as long as their decisions are based on some kind of anticipation (see Section 3.2).
+Because more information from the training set is exploited, it can be expected that these methods outperform myopic ones.
 Is this confirmed by experience? Are there situations where the advantage is significant? Our experiments in Section 4 provide answers to these questions.
 2.3 Choice of the classification component
-One source of difficulty when devising an ECTS method in the separable setting is that inputs differ from one time step to another.
-The number of measurements, and hence the input dimension, varies.
+[…]
+One source of difficulty when devising an ECTS method in the separable setting is that, at testing time, inputs differ from one time step to another.
+When an incoming time series is progressively observed, the number of measurements, and hence the input dimension, varies.
 Two approaches have been used to deal with the problem.
@@
     A set of classifiers $\{h_{t}\}_{t\in[1,T]}$ is learned, each dedicated to a given time step $t$, and thus a given input dimension.
-In practice, authors often chose a limited subset of timestamps, usually a set of twenty (one measurement every 5% of the length of the time series), to restrict the number of classifiers to learn and therefore the associated computational cost.
+In practice, authors often choose a limited subset of timestamps, usually a set of twenty (one measurement every 5% of the length of the time series), to restrict the number of classifiers to learn and therefore the associated computational cost.
   2. 2.
@@
 multi-valued time series), into a fixed dimensional vector whatever $t$ and $d$.
-This may simply be the mean value and standard deviation of the available measurements (multiplied by the dimension $d$) or the result of a more sophisticated feature engineering as tested by ?
-(?).
-Deep learning architectures can also be used to learn an encoding of the time series in an intermediate layer.
-For instance, ?
-(?) use a CNN architecture, and ?
-(?) a FCN one.
-?
-(?) show that using deep neural architectures often perform well for time series classification.
+This may simply be the mean value and standard deviation of the available measurements (multiplied by the dimension $d$) or the result of a more sophisticated feature engineering as tested by Skakun et al.
+(2017).
+Deep learning architectures can also be used to learn an encoding of the time series in an intermediate layer (Wang et al., 2016; Sawada et al., 2022).
 Both approaches have their own limitations.
@@
 This will be the subject of future work.
-3 State of the art
-In this section, we position various existing proposed methods along the dimensions of the taxonomy presented in Section 2 (see Table 1).
-Then, we examine the methods and how they exemplify solutions to the general questions raised in the taxonomy.
-Thus, while each method combines solutions to all questions, in the following, we underline how each one brings an original solution to one specific problem.
-For instance, the Economy approach is separable, anticipation-based, and cost-informed at testing time.
-However, we’re focusing here on the anticipation-based dimension, since this is one of the first methods to have emphasized it and introduced an original solution for this aspect.
-References Classifier(s) (collection ✓) End2end Confidence Anticipation Cost informed
-EDSC (?) Shapelet
-ECTS’ (?) 1NN
-Reject (?) SVM
-RelClass (?) QDA, Linear SVM
-[…]
-2step/NoCluster (?) Linear SVM (✓) train & test
-ECDIRE (?) Gausian Process (✓)
-Stopping Rule (?) Gaussian Process (✓) train
-[…]
-ECEC (?) WEASEL (✓) train
-DDQN (?) MLP train
-TEASER (?) WEASEL (✓) train
-[…]
-DETSCNet (?) TCN train
-CALIMERA (?) MiniROCKET (✓) train
-ELECTS (?) LSTM train
-[…]
-EarlyStop-RL (?) MLP train
-Table 1: Table of published methods for the ECTS problem with their properties along dimensions underlined in the taxonomy.
+3 An organized state-of-the-art of separable methods
+In this section, approaches from the literature are considered and organized around two key notions from the introduced taxonomy: confidence-based and anticipation-based.
+Subsection 3.1 presents methods whose decisions are triggered in a myopic way, based on some confidence measure.
+Subsection 3.2 describes approaches using a non-myopic decision criterion, which attempts to anticipate likely continuations.
+Table 1 shows how various approaches described in the literature can be organized using the characteristics underlined in the proposed taxonomy.
+Table 1: Table of published separable methods for the ECTS problem with their properties along dimensions underlined in the taxonomy.
+Note that void values indicate that the corresponding property is not present in the referred system.
+References Classifier(s) (collection ✓) Confidence Anticipation Cost informed
+Reject (Hatami & Chira, 2013) SVM instant
+iHMM (Antonucci et al., 2015) HMM instant
+ECDIRE (Mori et al., 2017b) Gaussian Process (✓) instant
+Stopping Rule (Mori et al., 2017a) Gaussian Process (✓) instant
+ECEC (Lv et al., 2019) WEASEL (✓) sequence
+TEASER (Schäfer & Leser, 2020) WEASEL (✓) sequence
+SOCN (Lv et al., 2023) FCN sequence
+ECTS (Xing et al., 2012) 1NN instant
+RelClass (Parrish et al., 2013) QDA, Linear SVM instant
+2step/NoCluster (Tavenard & Malinowski, 2016) Linear SVM (✓) ✓(test)
+ECONOMY-$\gamma$-max (Zafar et al., 2021) XGBoost ++ tsfel (✓) ✓(test)
+CALIMERA (Bilski & Jastrzebska, 2023) MiniROCKET (✓)
+FIRMBOUND (Ebihara et al., 2025) LSTM
 3.1 Confidence-based approaches
-Most ECTS methods to date are separable, confidence-based, cost-informed at training time, and are not anticipation-based.
-They implement separately the prediction and the triggering components, they learn them using the costs, hence they are cost-informed at training time, but they decide to trigger a decision based on the information available at the current time step $t$ without trying to anticipate the likely future, and they base their decision upon the confidence of the predictions made by the classifier.
 There exist two families of confidence-based approaches.
@@
 $\bullet$ One basic method is to monitor $\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})$, the highest conditional probability estimated by the classifier, which is a simple measure of classifier confidence over time.
-As soon as it exceeds a value, which is an hyper parameter of the method, a prediction is made.
-We call this method Proba Threshold and use it as baseline for comparison later in our experiments.
-$\bullet$ The Reject method (?) uses ensemble consensus as a confidence measure.
+As soon as it exceeds a value, which is a hyperparameter of the method, a prediction is made.
+We call this method Proba Threshold and use it as a baseline for comparison later in our experiments.
+$\bullet$ The Reject method (Hatami & Chira, 2013) uses ensemble consensus as a confidence measure.
 For each time step, first (i), a pool of classifiers is trained by varying their hyperparameters (i.e.
@@
 Finally, the prediction is triggered as soon as both classifiers in the ensemble predict the same class value.
-In this case, the monitored confidence measure is binary (agreement or disagreement), there is no trigger threshold and thus this trigger model is free of hyper-parameters [^4] The Reject approach involves choosing the number of classifiers trained in step (i) and selected in step (ii) that could be considered as hyper-parameters of the monitored confidence measure..
 $\bullet$ Hidden Markov Models (HMMs) are naturally suited to the classification of online sequences.
@@
 However, the decision to make a prediction now or to postpone it must then involve a threshold so that the prediction is only made if the a posteriori probability of the best HMM is sufficiently high or is greater than that of the second-best.
-In reaction to this, ?
-(?) propose to replace the standard HMM with imprecise HMMs based on the concept of credal classifier.
+In reaction to this, Antonucci et al.
+(2015) propose to replace the standard HMM with imprecise HMMs based on the concept of credal classification.
 This eliminates the need to choose a threshold, since a decision is made when one classification “dominates” (according to a criterion based on probability intervals) all the others.
 $\bullet$ Rather than considering only the largest value predicted by the classifier, it is appealing to consider also the difference with the second largest value, since a large difference points to the fact that there is no tie between predictions to expect.
-This is one dimension used in the Stopping Rule (SR) approach (?). Specifically, the output of the system is defined as:
-$g(h({\mathbf{x}}_{t}))=\begin{cases}\emptyset&\mbox{if extra measures are queried;}\\ \hat{y}=\operatorname*{arg\,max}_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})&\mbox{when }\gamma_{1}\,p_{1}+\gamma_{2}\,p_{2}+\gamma_{3}\,\frac{t}{T}>0\end{cases}$ (9)
+This is one dimension used in the Stopping Rule (SR) approach (Mori et al., 2017a). Specifically, the output of the system is defined as:
+[…]
 where $p_{1}$ is the largest posterior probability $p(y|{\mathbf{x}}_{t})$ estimated by the classifier $h$, $p_{2}$ is the difference between the two largest posterior probabilities, and $\frac{t}{T}$ represents the proportion of the incoming time series at time $t$.
-The parameters $\gamma_{1}$, $\gamma_{2}$ and $\gamma_{3}$ are learned from the training set.
-$\bullet$ Using the same notations as SR, the Early Classification framework based on class DIscriminativeness and RELiability (Ecdire) (?) finds the earliest timestamp for which a threshold applied on $p_{1}$ is reached (defined as in Equation 9).
+The parameters $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$ are learned from the training set, using genetic algorithm.
+$\bullet$ Using the same notations as SR, the Early Classification framework based on class DIscriminativeness and RELiability (Ecdire) (Mori et al., 2017b) finds the earliest timestamp for which a threshold applied on $p_{1}$ is reached (defined as in Equation 5).
 Then, the quantity $p_{2}$ is monitored, and a second threshold is applied to trigger the prediction.
-$\bullet$ A different class of methods relies on searching telltale representations of subsequences, such that if the incoming time sequence $\mathbf{x}_{t}$ matches one or more of these representations, then its class can be predicted.
-Typically, these representations take the form of shapelets that discriminate well one class from the others (?).
-For instance, the Early Distinctive Shapelet Classification (Edsc) method learns a distance threshold for each shapelet, based on the computation of the Euclidean distance between the considered subsequence and all others valid subsequences in the training set (?).
-It selects a subset of them, based on a utility measure that combines precision and recall, weighted by the earliness.
-A prediction is made as soon as $\mathbf{x}_{t}$ matches one of these shapelets well enough.
-Because this family of methods is computationally expensive, extensions have been developed to reduce the computational load (?, ?).
-Other extensions aimed at improving the reliability of the predictions (?, ?), and tackling multivariate time series (?, ?, ?, ?).
+$\bullet$ Ringel et al.
+(2024) use the Learning Then Test (LTT) (Angelopoulos et al., 2021) calibration framework to address ECTS.
+In practice, the proposed approach greedily computes thresholds at each time step, in order to monitor some conditional control risk measure, given a pre-defined error rate.
+This paper investigates text applications.
+$\bullet$ Historically, a related scenario predates the ECTS problem but is different.
+In the sequential decision making and optimal statistical decisions frameworks (DeGroot, 2005; Berger, 1985), the successive measurements are supposed to be independently and identically distributed (i.i.d.) according to a distribution of unknown “parameter” $\theta$.
+The problem is to determine as soon as possible whether the measurements have been generated by a distribution of parameter $\theta_{0}$ (hypothesis $H_{0}$) or of parameter $\theta_{1}$ (hypothesis $H_{1}$) with $\theta_{0}\neq\theta_{1}$.
+In the Wald’s Sequential Probability Ratio Test (Wald & Wolfowitz, 1948; Ghosh & Sen, 1991), the log-likelihood ratio $R_{t}\;=\;\log\frac{P(\langle x_{1}^{i},\ldots,x_{t}^{i}\rangle\;|\;y=-1)}{P(\langle x_{1}^{i},\ldots,x_{t}^{i}\rangle\;|\;y=+1)}$ is computed and compared with two thresholds that are set according to the required error of the first kind $\alpha$ (false positive error) and error of the second kind $\beta$ (false negative error).
+This beautiful setting allows one to get optimal decision times at the cost of being able to compute the log-likelihood.
+However, it differs from the ECTS problem, where successive observations are dependent.
+The i.i.d.
+assumption being not valid for the ECTS problem, a generalization to the non-i.i.d.
+case was proposed by Tartakovsky et al.
+(2014), providing guarantees for the asymptotic case (with $T\rightarrow\infty$).
+Despite this latter limitation, Ebihara et al.
+(2025) has recently applied this type of approach to ECTS with finite time horizons.
+The authors propose practical ways of both estimating $R_{t}$ (Ebihara et al., 2023) and triggering times by solving a backward induction problem (Tartakovsky et al., 2014).
 3.1.2 Sequence-based decision criterion
 Other approaches propose sequence-based confidence measures specifically designed for the ECTS problem.
-$\bullet$ The Effective Confidence-based Early Classification (Ecec) (?) proposes a confidence measure based on the sequence of predicted class values, from the first one observed to the current timestamp.
+$\bullet$ The Effective Confidence-based Early Classification (Ecec) (Lv et al., 2019) proposes a confidence measure based on the sequence of predicted class values, from the first one observed to the current timestamp.
 At each time step, this approach exploits the precision of the classifier to estimate the probability for each possible class value $y\in{\cal Y}$ of being correct if predicted.
 Then, assuming that successive class predictions are independent, the proposed confidence measure represents the probability that the last class prediction is correct given the sequence of predicted class values.
-The proposed confidence measure is monitored over time, and prediction is triggered if this measure exceeds a certain threshold $\gamma$ tuned as the single hyper-parameter.
-$\bullet$ The Teaser (Two-tier Early and Accurate Series classifiER) (?) approach considers the problem of whether or not a prediction should be triggered as a classification task, the aim of which is to discriminate between correct and bad class predictions.
+The proposed confidence measure is monitored over time, and prediction is triggered if this measure exceeds a certain threshold $\gamma$ tuned as the single hyperparameter.
+$\bullet$ The Teaser (Two-tier Early and Accurate Series classifiER) (Schäfer & Leser, 2020) approach considers the problem of whether or not a prediction should be triggered as a classification task, the aim of which is to discriminate between correct and bad class predictions.
 As the authors point out, the balance of this classification task varies according to the time step considered $t\in[1,T]$.
@@
 To exploit this idea, a collection of one-class SVMs is used, learning hyper-spheres around the correct predictions for each time step.
-A prediction is triggered when it falls within these hyper-spheres for $\nu$ consecutive time steps ($\nu$ being a parameter of the method).
-$\bullet$ The Second-Order Confidence Network approach (Socn) (?) considers, as does Teaser, the same classification task aiming to discriminate between correct and bad predictions.
-To learn this task, a transformer (?) is used, taking as input the complete sequence of conditional probabilities estimated by the classifier $h$, from the first time step, up to the current time step.
-A confidence threshold $\nu$ is learned by minimizing the same cost function as (?), above which the prediction is considered reliable and therefore triggered.
+A prediction is triggered when it falls within these hyper-spheres for $\nu$ consecutive time steps.
+$\bullet$ The Second-Order Confidence Network approach (Socn) (Lv et al., 2023) considers, as does Teaser, the same classification task aiming to discriminate between correct and bad predictions.
+To learn this task, a transformer (Vaswani et al., 2017) is used, taking as input the complete sequence of conditional probabilities estimated by the classifier $h$, from the first time step, up to the current time step.
+A confidence threshold $\nu$ is learned by minimizing the same cost function as Lv et al.
+(2019) do, above which the prediction is considered reliable and therefore triggered.
 3.2 Anticipation-based methods
-One way of designing approaches that anticipate future measurements is to achieve classification of an incomplete time series while guaranteeing a minimum probability threshold according to which the same decision would be made on the complete series.
-This is the case of the Reliability Classification (RelClass) approach (?).
+$\bullet$ One way of designing approaches that anticipate future measurements is to achieve classification of an incomplete time series while guaranteeing a minimum probability threshold, according to which the same decision would be made on the complete series.
+This is the case of the Reliability Classification (RelClass) approach (Parrish et al., 2013).
 Assuming that the measurements are i.i.d.
 and generated by a Gaussian process, this approach estimates $p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})$ the conditional probability of the entire time series ${\mathbf{x}}_{T}$ given an incomplete realization ${\mathbf{x}}_{t}$ and thus derives guarantees of the form:
-[…]
-where ${\mathbf{x}}_{T}$ is a random variable associated with the complete times series, $\gamma$ is a confidence threshold, and $h_{T}$ is the classifier learned over complete times series.
-At each time step $t$, $p(h_{T}({\mathbf{x}}_{T})=y|{\mathbf{x}}_{t})$ is evaluated and a prediction is triggered if this term becomes greater than the threshold $\gamma$, which is the only hyper-parameter to be tuned.
-Another way of implementing anticipation-based approaches is to exploit the continuations of training time series, which are full-length.
-One of the first method for ECTS (?) has been derived into such an anticipation-based approach (?).
-The first, called Early Classification on Time Series (Ects) (?), exploits the concept of Minimum Prediction Length (MPL), defined as the earliest time step for which the predicted label should not change for the incoming time series $\mathbf{x}_{t}$ from $t$ to $T$.
-This is estimated by looking for the 1NN of $\mathbf{x}_{t}$ in the training set, and check whether from $t$ onward, its predicted label did not change.
+$p\bigl(h_{T}({\mathbf{x}}_{T})=y|{\mathbf{x}}_{t}\bigr)\,=\,\int_{{\mathbf{x}}_{T}\text{ s.t. }h_{T}({\mathbf{x}}_{T})=y}\,p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})\,d{\mathbf{x}}_{T}\,\geq\,\gamma$
+where ${\mathbf{x}}_{T}$ is a random variable associated with the complete time series, $\gamma$ is a confidence threshold, and $h_{T}$ is the classifier learned over the complete time series.
+At each time step $t$, $p(h_{T}({\mathbf{x}}_{T})=y|{\mathbf{x}}_{t})$ is evaluated and a prediction is triggered if this term becomes greater than the threshold $\gamma$, which is the only hyperparameter to be tuned.
+$\bullet$ Another way of implementing anticipation-based approaches is to exploit the continuations of training time series, which are full-length.
+One of the first methods for ECTS has been derived into such an anticipation-based approach.
+The first, called Early Classification on Time Series (Ects) (Xing et al., 2009), exploits the concept of Minimum Prediction Length (MPL), defined as the earliest time step for which the predicted label should not change for the incoming time series $\mathbf{x}_{t}$ from $t$ to $T$.
+This is estimated by looking for the 1NN of $\mathbf{x}_{t}$ in the training set, and checks whether from $t$ onward, its predicted label does not change.
 To be more robust, the MPL is defined based on clusters computed on full-length training time series to estimate the best decision time.
-The approach has been extended later on to speed up the learning stage (?).
-This method looks in its own way at the likely future of $\mathbf{x}_{t}$ i.e.
+The approach has been extended later on to speed up the learning stage (Xing et al., 2012).
+This method looks in its own way at the likely future of $\mathbf{x}_{t}$ - i.e.
 an incomplete time series belongs to a cluster whose continuations are known - and thus can be considered as an anticipation-based method.
-?
-(?) present a method that claims explicitly to be “non myopic” in that a decision is taken at time $t$ only insofar as it seems that no better time for prediction is to be expected in the future.
+$\bullet$ Dachraoui et al.
+(2015) present a method that claims explicitly to be “non-myopic” in that a decision is taken at time $t$ only insofar as it seems that no better time for prediction is to be expected in the future.
 In order to do this, the family of Economy methods estimates the future cost expectation based on the incoming time series $\mathbf{x}_{t}$.
-This can be done since the training data consists of full-length time series and therefore a Learning Using Privileged Information (LUPI) (?) is possible.
-More formaly, the objective is to trigger a decision when $\mathbb{E}_{y,\hat{y}}[\mathcal{L}(\hat{y},y,t)|\mathbf{x}_{t}]$ is minimal, with:
+This can be done since the training data consists of full-length time series and therefore a Learning Using Privileged Information (LUPI) (Vapnik & Vashist, 2009) is possible.
+More formally, the objective is to trigger a decision when $\mathbb{E}_{y,\hat{y}}[\mathcal{L}(\hat{y},y,t)|\mathbf{x}_{t}]$ is minimal, with:
 $\mathbb{E}_{y,\hat{y}}[\mathcal{L}(\hat{y},y,t)|\mathbf{x}_{t}]=\sum_{y\in{\cal Y}}P(y|\mathbf{x}_{t})\,\sum_{\hat{y}\in{\cal Y}}P(\hat{y}|y,\mathbf{x}_{t})\,C_{m}(\hat{y}|y)\;+\;C_{d}(t)$ (6)
@@
 A key challenge in this framework is to design approaches achieving the most useful partition for predicting decision costs expectation.
-In the first article which presents this framework (?), a method, called Economy-$K$, is designed as follows.
+In the first article which presents this framework Dachraoui et al.
+(2015), a method, called Economy-$K$, is designed as follows.
 (i) A partition of training examples is first performed by a K-means algorithm ; (ii) then a simple model uses the Euclidean distance as a proxy of the probability that $\mathbf{x}_{t}$ belongs to each group; (iii) the continuation of training time series within each group is exploited to predict the cost expectation for future time steps.
-In order to avoid the clustering step with the associated choice of hyper parameters, (?) presented a variant called NoCluster which uses the 1 nearest neighbor in the training set in order to guess the likely future of $\mathbf{x}_{t}$.
-Then, Economy-$\gamma$ was introduced in (?) which relies on a supervised method to define a confidence-based partition of training time series.
+In order to avoid the clustering step with the associated choice of hyperparameters (Tavenard & Malinowski, 2016) presented a variant called NoCluster which uses the 1-nearest neighbor in the training set in order to guess the likely future of $\mathbf{x}_{t}$.
+Then, Economy-$\gamma$ was introduced by Achenchabe et al.
+(2021) which relies on a supervised method to define a confidence-based partition of training time series.
 The algorithm, dedicated to binary classification problems, is designed as follows: (i) a collection of partitions is constructed by discretizing the output of each classifier $\{h_{i}\}_{i\in[1,T]}$ into equal-frequency intervals, the groups thus formed correspond to confidence levels for each time step; (ii) at the current time $t$, the incoming time series $\mathbf{x}_{t}$ belongs to only one group, since the output of the classifier $h_{t}$ falls within a particular confidence level; (iii) then, a Markov chain model is trained to estimate the probabilities of the future time step confidence levels.
-Economy-$\gamma$-max (?) generalizes this approach to multi-class problems, aggregating the multiple conditional probabilities in the classifiers’ output by using only the most probable class value.
-Finally, Calimera (?) uses anticipation about the future from another perspective.
-Instead of trying to guess the likely continuation of $\mathbf{x}_{t}$ which allows one to compute expected future costs, and therefore to wait until there seems no better time to make a prediction, their method is based on predicting directly the difference in cost between predicting the class now or wait at least one more time step.
+Economy-$\gamma$-max (Zafar et al., 2021) generalizes this approach to multi-class problems, aggregating the multiple conditional probabilities in the classifiers’ output by using only the most probable class value.
+$\bullet$ Calimera (Bilski & Jastrzebska, 2023) uses anticipation about the future from another perspective.
+Instead of trying to guess the likely continuation of $\mathbf{x}_{t}$ which allows one to compute expected future costs, and therefore to wait until there seems no better time to make a prediction, their method is based on predicting directly the difference in cost between predicting the class now and the best reachable cost from next timestep to last one $T$.
 If this difference is positive, then it is better to postpone the prediction.
-They advocate furthermore, that a calibration step should intervene above the regression in order to make a decision.
-3.3 Reinforcement Learning based methods
-To learn an ECTS function, it is possible to use an agent that learns what to do by exploring the outcomes associated with different decision strategies as it interacts with the world.
-The ECTS problem can therefore be recast as a reinforcement learning (RL) problem.
-In RL, an agent must learn to associate an action $a$ with each observable state $s$ so that the expected gain be maximized.
-Let us suppose that the agent-environment interactions break naturally into subsequences which we call episodes
-At each time step $t$, the agent perceives the environment’s state $s_{t}$ (i.e.
-$\mathbf{x}_{t}$), choses an action $a_{t}$ (i.e.
-either make a prediction now and measure the gain $G_{t}=-\mathrm{C}_{m}(\hat{y}|y)-\mathrm{C}_{d}(\hat{t})$, or postpone the decision and receive the next measurement $x_{t+1}$) according to its current policy $\pi$, where $\pi(a|s)=p(a_{t}=a|s_{t}=s)$ is the probability that action $a$ is taken given the state $s$.
-The goal is for the agent to learn an optimal policy $\pi^{\star}$ from sequences of interactions with its environment $\langle s_{t},a_{t},s_{t+1}\rangle$.
-This can be done by computing the utility function $Q(s,a)$ defined for all (state, action) pairs.
-By definition:
-$Q_{\pi}(s,a)\;=\;\operatornamewithlimits{\mathbb{E}_{\>\pi}^{\>}}\left[G_{t}|S_{t}=s,A_{t}=a\right]$ (12)
-and the optimal policy can be derived from:
-$Q^{\star}(s,a)\;=\;\operatornamewithlimits{max}_{\pi}Q_{\pi}(s,a)$ (13)
-It suffices at each observed state $s$ to chose action $a^{\star}$ such as:
-$a^{\star}\;=\;\operatorname*{arg\,max}_{a}Q^{\star}(s,a)$ (14)
-One way to learn the function $Q^{\star}$ is using the Q-learning algorithm (?, ?) [^5] Note that other approaches are possible in the reinforcement learning scenario, like learning the utility function $V(s)$ and using TD-learning, or even learning the policy $\pi$ directly.
-The Q-learning approach is however widely used.
-and its variants for continuous definition of environment’s states, such as Deep Q-learning (?).
-In End-to-end approaches, only one function is responsible for both stopping and making a prediction about the class of $\mathbf{x}_{t}$ whereas, in separable approaches involves two combined functions, respectively dedicated to triggering the prediction and to the classification itself.
-In the case of Reinforcement Learning based approaches, this distinction takes the following form:
-  1. 1.
-    The separable approach.
-RL can be used to learn the Trigger function only, once the classifiers $h_{t}$ have been learnt.
-In that case, the set of actions $\mathcal{A}_{t}$ at each time step $t$ is restricted to two elements: {‘decide now’, ‘postpone decision’}.
-The $Q$ function evaluates for each state the expected gain for each of the two possibilities, allowing one to decide what to do at each time step.
-  2. 2.
-    The End-to-End approach.
-RL can also be used to learn at once both when to trigger a prediction and what prediction to make.
-In principle, it suffices to extend the set of actions to $\mathcal{A}_{t}=\{`\textit{postpone decision'},c_{1},\ldots,c_{N}\}$, where there are $N$ classes.
-Either, the agent postpones the decision, or it predicts a class for the incoming time series $\mathbf{x}_{t}$.
-In the literature, Reinforcement Learning-based ECTS approaches frequently include Deep Learning modules:
-Separable RL approaches:
-The Earliest (Early and Adaptive Recurrent Label ESTimator) uses a RNN architecture (?) to make the prediction and a Reinforcement Learning agent trained jointly using policy gradient, to trigger prediction or not.
-If a prediction is triggered, the hidden representation given by the RNN is sent to a Discriminator, whose role is to predict a class, given this representation.
-The model has been adapted to deal with irregularly sampled time series (?).
-?
-(?) extend the ECTS framework to channel filtering, using here also Reinforcement Learning.
-End-to-end RL approaches:
-?
-(?, ?) use a Deep Q-Network (?), alongside a specifically designed reward signal, encouraging the agent to find a good trade-off between earliness and accuracy.
-Those types of approaches also naturally extend to online settings where time series are not of fixed length.
-(?) introduce EarlytStop-RL, in which model-free RL is used to address the problem of early diagnosis of lung cancer.
-3.4 Deep Learning based methods
-Alongside the Reinforcement Learning based approaches, there exist deep learning methods that do not use RL.
-The Decouple ETSC Network (Detscnet) (?) architecture leverages a gradient projection technique in order to jointly learn two sub-modules: one for variable-length series classification, the other for the early exiting task.
-The End-to-end Learned Early Classification of Time Series method (Elects) leverages a LSTM architecture, adding a stopping prediction head to the network and adapting the loss function to promote good early predictions (?).
+They advocate furthermore, that a calibration step should intervene on the collection of classifiers, in order to build meaningful regression target to learn the trigger model.
 4 Experiments & Results
-This section presents the extensive set of experiments carried out in order to provide a consistent and fair evaluation of a wide range of the existing literature’s methods.
+This section presents the extensive set of experiments carried out in order to provide a consistent and fair evaluation of a wide range of existing literature’s methods.
 We first describe the experimental protocol used.
@@
   • Section 4.1 introduces the experimental protocol as well as the global evaluation methodologies.
-  • In Section 4.2, the main state-of-the-art and the three baseline methods are evaluated using a widely used cost setting, i.e.
-with a a binary balanced misclassification cost and a linear delay cost.
-  • In Section 4.3, methods are tested in an anomaly detection scenario where the misclassification cost matrix is severely imbalanced, with false negatives being much more costly than false positives, and where the delay cost is no longer linear with time but increases exponentially with time.
-  • Finally, Section 4.4 briefly describes a set of other experimental setups, derived from either standard setting or the anomaly detection one, including for instance testing the impact of z-normalization.
-Complementary results can be found in Appendix B.
+  • In Section 4.2, eight state-of-the-art methods and the three baseline ones are evaluated using a widely used cost setting, i.e.
+with a binary balanced misclassification cost and a linear delay cost.
+  • In Section 4.3, methods are tested in an anomaly detection scenario [^4] We consider anomalies to be actual phenomena of interest, such as the failure of a machine (Boniol et al., 2024).
+These anomalies may be accompanied by revealing precursor signals in the time series, which an ECTS system should be able to detect by optimizing both the accuracy of prediction and its earliness.
+where the misclassification cost matrix is severely imbalanced, with false negatives being much more costly than false positives, and where the delay cost is no longer linear with time but increases exponentially with time.
+  • Finally, Section 4.4 briefly describes a set of other experimental setups, derived from either the standard setting or the anomaly detection one, including, for instance, testing the impact of z-normalization.
+Complementary results can be found in Appendix LABEL:app_results.
 […]
-Figure 2: Experiments diagram: each line corresponds to one (or two) full benchmark runs, representing in total twelve full benchmarks.
-While this Section mainly discusses results about the cost settings in Subsections 4.2 and 4.3, many other alternative experiments are briefly analyzed in Subsection 4.4 and are more detailed in the Appendix B.
-The experiments presented here aim to evaluate the effects of design choices on method performance and thus to provide answers to the questions:
-  • Do anticipation-based methods perform better than blind or myopic ones?
-  • Do methods that are cost-informed for their decision (i.e. explicitly estimating costs) perform better than methods that are cost-uninformed (e.g. confidence-based) (see Section 2.3.2).
-  • How the various methods fare when modifying the form of the delay cost and/or the misclassification cost matrix?
+Figure 2: Experiments diagram.
+While this section mainly discusses results about the cost settings in Subsections 4.2 and 4.3, many other alternative experiments are briefly analyzed in Subsection 4.4 and are more detailed in the Appendix LABEL:app_results.
+Details on the used datasets can be found in Appendix E.
+The experiments presented here aim to evaluate the effects of design choices on method performance and thus provide answers to the questions:
+  • Do anticipation-based methods perform better than myopic ones?
+  • Do methods that are cost-informed for their decisions (i.e. explicitly estimating costs) perform better than methods that are cost-uninformed? (see Section 2.2.1)
+  • How do the various methods fare when modifying the form of the delay cost and/or the misclassification cost matrix?
 4.1 Experimental Protocol
-This section covers the shared part of the experimental protocol for all experiments irrespective of the choice of the cost functions (see Sections 4.2 and 4.3 for this).
+This section covers the shared part of the experimental protocol for all the experiments, irrespective of the choice of the cost functions (see Sections 4.2 and 4.3 for this).
+Additional implementation specifications can be found in Appendix B.
 4.1.1 Evaluation of the performance
-The ECTS problem explicitly balances the pressure to make a correct prediction and a time pressure.
-The correctness of the prediction is measured by the misclassification cost $\mathrm{C}_{m}(\hat{y}|y)$ where $\hat{y}$ is the prediction and $y$ is the true class.
-The time pressure is sanctioned by a delay cost $\mathrm{C}_{d}(t)$ that is assumed to be positive and, in most applications, an increasing function of time.
-Sometimes, the two can be combined when the misclassification cost is function of the time : $\mathrm{C}_{md}(\hat{y}|y,t)$.
-Therefore, for each test times series $\mathbf{x}^{i}$, an ECTS method incurs a cost assumed to be of the following additive form: $\mathrm{C}_{m}(\hat{y}_{i}|y_{i})\,+\,\mathrm{C}_{d}(\hat{t})$, where $\hat{t}$ is the time when the system decided to make a prediction, this prediction being $\hat{y}_{i}$.
-For a test set of $M$ time series, the average cost of a method is:
+[…]
+For a test set of $M$ time series, the average cost of a method (Equation 8) is used as the criterion with which to evaluate the methods.
 $AvgCost_{\text{test}}\;=\;\frac{1}{M}\sum_{i=1}^{M}C_{m}(\hat{y}_{i}|y_{i})+C_{d}(\hat{t}_{i})$ (8)
-This is accordingly the criterion with which we evaluated the methods in the experiments reported in this paper.
 In addition, in order to assess how the methods adapt to various balances between the misclassification and the delay costs, we vary the settings of these costs by weighting them during training and testing.
-The performance of the methods is therefore evaluated using the weighted average cost, as defined in Equation 16, for different values of the costs balance $\alpha$, ranging from 00 to $1$, with a $0.1$ step:
-$AvgCost_{\alpha}=\frac{1}{N}\sum_{i=0}^{N}\alpha\times C_{m}(\hat{y}_{i}|y_{i})+(1-\alpha)\times C_{d}(\hat{t}_{i})$ (16)
+The performance of the methods is therefore evaluated using the weighted average cost, as defined in Equation 9, for different values of the costs balance $\alpha$, ranging from 0 to $1$, with a $0.1$ step:
+$AvgCost_{\alpha}=\frac{1}{M}\sum_{i=0}^{M}\alpha\times C_{m}(\hat{y}_{i}|y_{i})+(1-\alpha)\times C_{d}(\hat{t}_{i})$ (9)
 Small values of $\alpha$ correspond to a high delay cost and a small misclassification cost ; inversely, large values of $\alpha$ give more weight to the misclassification cost with a lower delay cost.
-4.1.2 Optimization of the parameters of the methods
+4.1.2 Comparing the trigger methods
+Our experiments aim foremost at comparing the trigger methods used that are responsible for deciding when to make a prediction about the class of the incoming time series.
+To this end, all compared methods in our experiments use the same prediction component.
+As advocated by Bilski & Jastrzebska (2023), we have chosen the MiniROCKET algorithm (Dempster et al., 2021) to be the base classifier for all methods.
+It is indeed recognized as among the best performing classifiers in the time series classification literature as well as one of the fastest ones.
+However, we have carried out additional experiments with two other classifiers (see Section 4.4).
+In our experiments, we have additionally reported results for end-to-end methods in order to situate the performance of these methods in respect to separable ones in Appendix C.2.
 […]
-4.1.3 Comparing the trigger methods
-In order to carry out a fair comparison between the tested methods, we isolated as far as possible the triggering function, responsible for deciding when to stop receiving measurements, from the prediction one, responsible for predicting a label to the incoming time series.
-As advocated by ?
-(?), we have chosen the MiniROCKET algorithm (?) to be the base classifier for all methods.
-It is indeed recognized as among the best performing classifier in the Time Series classification literature as well as one of the fastest one.
-Of course, distinguishing the decision component from the prediction one is only possible for the “separable” methods.
-In our experiments, we chose to not evaluate “end-to-end” methods, leaving that for future work.
-Trigger models: Nine trigger models were selected from the literature based on their usage and their performances [^6] The EDSC algorithm (?), even though available in the provided library, is not included in the following experiments, due to high space and time complexity (which hinders fair comparisons).
-  • Economy-$\gamma$-Max (?): triggers a decision if the predicted cost expectation is the lowest at time $t$ when compared with the expected cost for all future time steps (cf.
-Section 3.2, Anticipation-based).
-  • Calimera (?): triggers a decision when a regressor model which predicts the difference between the current observed cost and the minimum cost in the future is negative (cf.
-Section 3.2, Anticipation-based).
-  • Stopping Rule (?): uses a trigger function based on a linear combination of confidence estimates and a delay measure linear on time (cf. Section 3.1, Confidence-based).
-[…]
-[…]
-  • Ecec (?): defines a confidence measure, based on the aggregated confidence of the predictions up to time $t$, and triggers a prediction if it exceeds a threshold, tuned by explicit grid-search (cf.
-Section 3.1, Confidence-based).
-  • Ecdire (?): determines “safe” timestamps, based on classifier performance, from which predictions about possible classes can be made.
-Predictions cannot be triggered if those timestamps have not been reached.
-In addition, the difference between the two highest predicted probabilities must also exceed a certain threshold.
-(cf.
-Section 3.1, Confidence-based).
-  • Ects (?): computes the first time $t$ for which nearest neighbors of the incoming time series $\mathbf{x}_{t}$ in the training set were given a label that did not change by the classifier (cf.
-Section 3.2, Anticipation-based).
 All these methods have been re-implemented using Python, reproducing results close to the published ones.
-Except the code for the Ects implementation, which has been taken from ?
-(?).
-Hyper-parameters are the ones chosen in the original published methods.
-Code to reproduce the experiments is available publicly at https://github.com/ML-EDM/papers_experiments/tree/main/ects_reloading.
+Except for the code for the Ects implementation, which has been taken from Kladis et al.
+(2021).
+Hyperparameters are the ones chosen in the original published methods.
+Code to reproduce the experiments is available publicly at (see https://github.com/ML-EDM/ECTS_survey) .
 Baselines: Furthermore, in order to evaluate the benefits, if any, of the various methods, it is telltale to compare them with simple ones. We chose three such baselines:
   • Asap (As Soon As Possible) always triggers a prediction at the first possible timestep.
-  • Alap (As Late As Possible) always waits the complete series to trigger the prediction.
-  • Proba Threshold is a natural, confidence-based, cost-informed at training time, baseline: it triggers a prediction if the estimated probability of the likeliest prediction exceeds some threshold, found by grid search (cf.
+  • Alap (As Late As Possible) always waits for the complete series to trigger the prediction.
+  • Proba Threshold is a natural, confidence-based, cost-informed, baseline: it triggers a prediction if the estimated probability of the likeliest prediction exceeds some threshold, found by grid search (cf.
 Section 3.1, Confidence-based).
-4.1.4 Calibration of the classifications
-Like ?
-(?), we add a calibration step when learning the classifiers, i.e.
-Platt’s scaling (?).
-Indeed, as we are dealing with collections of independently trained classifiers, the prediction scores may not remain consistent with one another over the time dimension.
-However, the trigger methods usually have their parameters set with the same values for all time steps.
-This is the case, for example with the Proba Threshold approach.
-In addition, some approaches such as Calimera and Economy-$\gamma$-Max exploit the estimated posterior probabilities $\{p(y|\textbf{x}_{t})\}_{y\in\mathcal{Y}}$ to estimate the future cost expectation.
-It is therefore highly desirable for all classifiers, at all times, to have their output calibrated.
 4.1.3 Datasets and training protocol
-Datasets [^7] All original datasets of the paper can be downloaded, already prepared and splitted, from https://urlz.fr/qRqu: In order to be able to directly compare our results to past experiments, we first use the usual TSC datasets from the UCR Archive (?) with the default split.
+Two collections of datasets [^6] All original datasets of the paper can be downloaded, already prepared and split, from https://urlz.fr/qRqu: In order to be able to directly compare our results to past experiments, we first use the usual TSC datasets from the UCR Archive (Dau et al., 2019) with the default split.
 In total, we have used 77 datasets from the UCR Archive, i.e.
-the ones with enough training samples to satisfy our experimental protocol end-to-end (blue cylinder in Figure 4).
-In this way, most of the datasets used by either ?
-(?) and ?
-(?) or by ?
-(?) are contained in our experiments.
+the ones with enough training samples to satisfy our experimental protocol from the start to the end, i.e.
+with at least one example per class within each of the used disjoint subsets (see Splitting strategy below).
+(blue cylinder in Figure 2).
+In this way, most of the datasets used by either Mori et al.
+(2017a) and Lv et al.
+(2019), or by Achenchabe et al.
+(2021) are contained in our experiments.
 A second collection of non z-normalized datasets is also provided.
 In this way, the associated potential information leakage is avoided (see Section 1).
-Any difference with the performance obtained on the z-normalized data sets can thus signal the danger of z-normalization with firm evidence.
-Considering the limited amount of non z-normalized datasets within the UCR archive (?), we have decided to look for complementary new datasets so as to provide another collection of datasets.
-To this end, the Monash archive for extrinsic regression (?), provided 20 new time series datasets, for which we have discretized the the numerical target variable into binary classes based on a threshold value.
-For instance, if this threshold is equal to the median value of the regression target, the resulting classification datasets will be balanced in term of classes (as in Section 4.2).
-Note that this threshold can be chosen differently to get imbalanced datasets (as in Section 4.3.2), several thresholds could also be used to increase number of classes.
-As a result, we get a new set of classification tasks, as has recently been done by ?
-(?).
-Finally, 35 datasets have been gathered: 15 from the original archive and 20 from the Monash extrinsic regression archive.
+Any difference in the performance obtained on the z-normalized data sets can thus signal the danger of z-normalization with firm evidence.
+Considering the limited amount of non z-normalized datasets within the UCR archive (Dau et al., 2019), we have decided to look for complementary new datasets so as to provide another collection of datasets.
+To this end, the Monash archive for extrinsic regression (Tan et al., 2020), provided 20 new time series datasets, for which we have discretized the numeric target variable into binary classes based on a threshold value.
+For instance, if this threshold is equal to the median value of the regression target, the resulting classification datasets will be balanced in terms of classes (as in Section 4.2).
+Note that this threshold can be chosen differently to get imbalanced datasets (as in Section 4.3.2), several thresholds could also be used to increase the number of classes.
+As a result, we obtain a new set of classification tasks, as has recently been done by Middlehurst et al.
+(2024b).
+In the end, 34 datasets have been gathered: 14 from the original archive and 20 from the Monash extrinsic regression archive.
 (orange cylinder in Figure 2).
+More details about the selection of datasets can be found in Appendix B.4
 Splitting strategy:
-When not using predefined splits, the train sets are split into two distinct sets in a stratified fashion: a first one to train the different classifiers, corresponding to 60% of the training set and another one to train the trigger model, trained over the 40% left.
-The set used to train the classifiers, is itself split into two different sets in order to train calibrators, using 30% of the given data.
+When not using predefined splits, the train sets are split into two distinct sets in a stratified fashion: a first one to train the different classifiers, corresponding to 40% of the training set and another one to train the trigger model, trained over the 60% left.
+The set used to train the classifiers is itself split into two different sets in order to train calibrators, using 30% of the given data.
 Because of this procedure, we have been led to exclude some of the datasets, due to their limited training set size.
-All the experiments have been performed using a linux operating machine, with an Intel Xeon E5-2650 2.20GHz (24-cores) and 252GB of RAM.
+All the experiments have been performed using a Linux operating system, with an Intel Xeon E5-2650 2.20GHz (24-cores) and 252GB of RAM.
 Proceeding all datasets (including both blue and orange cylinders) over all competing approaches takes between 9-10 days, using MiniROCKET classifier, which is the most efficient tested.
@@
 4.2.1 Cost definition
-The misclassification cost is symmetrical and balanced, the delay cost is linear. They can be defined as follows:
+The misclassification cost is symmetrical and balanced; the delay cost is linear. They can be defined as follows:
 $C_{m}(\hat{y}|y)$ $=\mathbb{1}(\hat{y}\neq y)$
 $C_{d}(t)$ $=\frac{t}{T}$
-Thus, for each dataset, the AvgCost is bounded between 0 and 1.
+Thus, for each dataset, the $\textit{AvgCost}_{\alpha}$ is bounded between 0 and 1, as, within this cost definition, the average misclassification (resp.
+temporal) cost, across examples is equivalent to the 1Accuracy1-Accuracy (resp.
+Earliness) measure, with $Accuracy=\frac{1}{M}\sum_{i=1}^{M}\mathbb{1}(\hat{y}_{i}=y_{i})$ and $Earliness=\frac{1}{M\times T}\sum_{i=1}^{M}\hat{t}_{i}$.
 4.2.2 Results and analysis
 For comparability reasons, this first set of experiments is analyzed over the classical ECTS benchmark used in the literature so far (blue cylinder in Figure 2).
-Results over the new, non z-normalized, datasets can be found in Appendix B.
+Results over the new, non z-normalized, datasets can be found in Appendix LABEL:app_results.
 (a) Evolution of the mean ranks, for every $\alpha$, based on the AvgCost metric. Shaded areas correspond to 90% confidence intervals.
@@
 Specifically, we report this for $\alpha=0.5$ as shown in (b).
-Figure 3(b) provides a global views about the relative performances of the tested methods.
+[…]
+When evaluated by their average rank on all data sets with respect to the average cost (Equation 9), here for $\alpha=0.5$, four methods significantly outperform the others:
+Table 2: Leading ECTS methods and their properties along dimensions underlined in the taxonomy.
+Methods Confidence Anticipation Cost-informed
+Stopping Rule
+Proba Threshold
+Economy-$\gamma$-max ✓(test)
+Calimera
+Figure 3(b) provides a view about the relative performances of the tested methods in terms of the average cost induced using the methods for $\alpha=0.5$.
 The Wilcoxon-Holm Ranked test provides an overall statistical analysis.
@@
 Lower ranks denote better performance, and the methods connected by a horizontal bar are similar in terms of statistical significance.
-When evaluated by their average rank on all data sets with respect to the average cost (Equation 16), here for $\alpha=0.5$, four methods significantly outperform the others:
-        Methods Confidence Anticipation Cost informed
-Stopping Rule train
-Proba Threshold train
-Economy-$\gamma$-max train & test
-Calimera train
-Figure 3(a) allows a closer look, this time varying the relative costs of misclassification and delaying prediction using Equation 16, where a small value of $\alpha$ means that delay cost is paramount.
-$90\%$ level confidence intervals have been computed using bootstrap [^8] Resample with replacement has been done a large number of times (10.000 $\times$) and are reported as shaded colors on the figure.
-The statistic of interest is studied, here the mean, by examining the bootstrap distribution at the desired confidence level..
-Again, the same four methods top the others for almost every values of $\alpha$.
-Not surprisingly, the baseline Asap (predict as soon as possible) is very good when the delay cost is very high, while Alap (predict at time $T$) is very good when there is no cost associated with delaying decision.
 It is remarkable that, in this cost setting, the simple Proba Threshold method exhibits a strong performance for almost all values of $\alpha$.
 It is therefore worth including in the evaluation of new methods.
-However, while Figures 3(a), 3(b) are useful for general analysis, they do not provide insights about how the Accuracy vs Earliness trade-off is optimized for each of the competitor.
+However, while Figures 3(a), 3(b) are useful for general analysis, they do not provide insights about how the Accuracy vs.
+Earliness trade-off is optimized for each of the competitors.
 Figure 4 provides some explanation for this.
-Figure 4: Pareto front, displaying for each $\alpha$ the Accuracy on the $y$ axis and $Earliness$ on the $x$ axis.
+Figure 4: Pareto front, displaying for each $\alpha$ the Accuracy on the $y$-axis and EarlinessEarliness on the $x$-axis.
 Best approaches are located on the top left corner.
@@
 Each of the trigger model is optimizing the trade-off in its own way, resulting in many different approaches having points in the Pareto dominant set.
-In this figure, the two evaluation measures: ‘Accuracy’ and ‘Earliness’, are considered as dimensions in conflict.
+In this figure, the two evaluation measures: Accuracy and Earliness, are considered as dimensions in conflict.
 The Pareto front is the set of points for which no other point dominates with respect to both Accuracy and Earliness.
@@
 […]
-Each of the leading methods Stopping Rule, Proba Threshold, Economy and Calimera have at least one point on the Pareto front and generally exhibits a combined performance very close to it.
+Each of the leading methods Stopping Rule, Proba Threshold, Economy and Calimera have at least one point on the Pareto front and generally exhibit combined performance very close to it.
 A closer look reveals how each approach optimizes the Earliness vs.
@@
 It is worth noting that the two naive baselines Asap and Alap perform better than the majority of approaches on seven $\alpha$ values out of ten.
-This is specially the case when the delay cost is large, i.e.
+This is especially the case when the delay cost is large, i.e.
 for $\alpha\in[0.1,0.3]$, for which the Asap baseline is as competitive as top performers.
 Globally, the performance of Proba Threshold is remarkable in this cost setting.
-Even though, it is simply based on a single threshold on the confidence in the current prediction, its performance makes it one of the best methods.
-The results computed over the proposed datasets ensemble (i.e. orange cylinder) are displayed in Figure 8 of Appendix B. No significant changes can be observed in the ranking of competing approaches.
-One question is how such simple method, like Proba Threshold, can adapt to scenarios where the misclassification and delay costs, not being symmetrical for the misclassification cost, and not linear for the delay one, reflect other application settings.
+Even though it is simply based on a single threshold on the confidence in the current prediction, its performance makes it one of the best methods.
+The results computed over the proposed datasets ensemble (i.e.
+orange cylinder) are displayed in Figure LABEL:fig:std_orange_cylinder of Appendix LABEL:app_results.
+No significant changes can be observed in the ranking of competing approaches.
 4.3 Experiments with unbalanced misclassification and non-linear delay costs
@@
 0 if correctly classified and 1 otherwise) and that the delay cost was a linear function of time.
-There are however applications where these assumptions do not hold, for instance predictive maintenance or hospital emergency services, are characterized by (i) imbalanced misclassification costs (e.g.
-it is more costly to have to repair a machine than to carry out a maintenance operation that turns out not being necessary) and by (ii) non linear delay cost (e.g.
+There are, however, applications where these assumptions do not hold, for instance, predictive maintenance or hospital emergency services, are characterized by (i) unbalanced misclassification costs (e.g.
+it is more costly to have to repair a machine than to carry out a maintenance operation that turns out not being necessary) and by (ii) non-linear delay costs (e.g.
 usually, the later the surgical operation is decided, the costlier it is to organize it and the larger the risk for the patient).
@@
 wrongly predicting an anomaly) (see Figure 5(b)).
-For this last situation, the cost was arbitrarily set to 1.
+For this last situation, the delay cost was arbitrarily set to 1.
 The delay cost is defined as an exponential function of time.
-In order to have a cost commensurable with the misclassification one, we decided that waiting for the entire time series to be seen, at $T$, would cost 100 (see Figure 5(a)), starting at 1 for $t=0$ and reaching 100 when $t=T$.
+In order to have a delay cost commensurable with the misclassification one, we decided that waiting for the entire time series to be seen, at $T$, would cost $100\times\alpha$ (see Figure 5(a)), starting at $(1-\alpha)$ for $t=0$ and reaching $100\times\alpha$ when $t=T$.
 4.3.2 Results and analysis
@@
 Thus, we choose to use the new non z-normalized datasets collection (orange cylinder in Figure 2).
-In order for the imbalanced misclassification cost to make sense, those datasets have been altered so that the minority class represents 20% of all labels.
-As explained in Section 4.1, some extrinsic regression datasets are turn into classification ones.
+In order for the imbalanced misclassification cost to make sense, the datasets have been altered so that the minority class represents 20% of all labels.
+As explained in Section 4.1, some extrinsic regression datasets are turned into classification ones.
 In these cases, the threshold value has been set to the second decile of the regression target.
@@
 Specifically, for $\alpha=0.5$ as shown in (b).
-Results from the Wilcoxon-Holm Ranked test (both regarding the average rank and the value for AvgCost) (see Figure 6(b)) and from the AvgCost plot (see Figure 6(a)) with varying values of $\alpha$ (in Equation 16) show that now the best method overall is Economy which is both cost-informed at training and testing time, beside being anticipating-based.
-However, Stopping-Rule is a very strong contender while being cost-informed at training time but not at testing time and confidence-based.
+Results from the Wilcoxon-Holm Ranked test (both regarding the average rank and the value for AvgCost) (see Figure 6(b)) and from the AvgCost plot (see Figure 6(a)) with varying values of $\alpha$ (in Equation 9) show that now, the best method overall is Economy which is cost-informed at testing time, in addition to being anticipation-based.
+However, Stopping-Rule is a very strong contender while being cost-informed but not at testing time and confidence-based.
 There is a reason for it.
-When Stopping Rule equals or overpasses Economy, this is for high values of $\alpha$ when the delay cost loses its importance, therefore leaving the misclassification cost to reign and confidence-based methods to be good.
-It may come as a surprise that Calimera lags behind Economy for $\alpha\in[0,0.4]$, despite being similarly based on the estimation of future cost expectation.
+When Stopping Rule equals or overpasses Economy, this applies to high values of $\alpha$ when the delay cost loses its importance, therefore leaving the misclassification cost to reign and confidence-based methods to be effective.
+It may come as a surprise that Calimera lags behind Economy for $\alpha\in[0,0.4]$, despite being similarly based on the estimation of future cost expectations.
 One reason for this is that the cost expectation is achieved by considering only the predicted class.
@@
 However, it is noticeable that the cost of delay only increases sharply after around 60% of the complete time series has been observed, which is generally sufficient to exceed the confidence threshold.
-Hence, Proba Threshold does not suffer from high delay costs that are to come, and exhibits a good performance here.
-Figure 7 plots the Pareto front considering two axes based on decision costs.
+Hence, Proba Threshold does not suffer from high delay costs that are to come, and exhibits good performance here.
+Figure 7 plots the Pareto front, considering two axes based on decision costs.
 The horizontal axis corresponds to the average delay cost incurred for each example, normalized by the worst delay cost paid at $t=T$.
 It is better to be on the left of the $x$-axis.
-The vertical axis corresponds to 1 minus the misclassification cost incurred for each example, normalized by the worst prediction cost.
+The vertical axis corresponds to one minus the misclassification cost incurred for each example, normalized by the worst prediction cost.
 It is better to be high on the $y$-axis.
-We observe that the Pareto front is composed almost exclusively of points corresponding to the Economy method.
+We observe that the Pareto front is composed almost exclusively of points corresponding to the Economy and Calimera methods.
 This is consistent with the evaluation based on the AvgCost metric.
@@
 In this section, complementary experiments, namely ablation studies as well as sanity checks are briefly discussed.
-For the sake of brevity, the figures supporting the analysis are reported in Appendix B.
+For the sake of brevity, the figures supporting the analysis are reported in Appendix LABEL:app_results.
 Impact of removing calibration
-?
-(?) assert that calibration of the classifiers is paramount for the performance of ECTS algorithms.
-In order to test this claim, we have repeated the experiments removing the calibration step.
-The examples used for calibration have been removed as well during training, so that all else remains the same as before.
-The results of Figure 11 show that indeed Calimera suffers greatly if no calibration is done.
-Indeed, this approach relies on estimating the expectation of future costs via a regression problem, and a miscalibration may have a negative impact on the built targets.
+Bilski & Jastrzebska (2023) assert that calibration of the classifiers is paramount for the performance of ECTS algorithms.
+In order to test this claim, we have repeated the experiment, removing the calibration step.
+The examples used for calibration have also been removed during training, so that all else remains the same as before.
+The results of Figure LABEL:fig:calib_heatbar in Appendix LABEL:calib show that indeed Calimera suffers greatly if no calibration is done.
+Indeed, this approach relies on estimating the expectation of future costs via a regression problem, and miscalibration may have a negative impact on the built targets.
 For its part, Proba Threshold suffers somewhat mildly.
-This is no surprise as they rely on a single threshold on the confidence of the prediction for all time steps.
+This is no surprise, as they rely on a single threshold on the confidence of the prediction for all time steps.
+Impact of removing non-myopia
+Experiments have shown that anticipation-based approaches tend to outperform the myopic competitors; we further validate this by conducting an ablation study over the non-myopia of both Economy and Calimera.
+Figure LABEL:fig:non_myopia in Appendix LABEL:non_myopic_ablation clearly shows that, for weak temporal costs, the anticipation-based property is critical, whereas myopic counterparts significantly fall behind original methods.
+Intuitively, for higher temporal costs, the anticipation-based property is less important as decisions tend to be taken earlier, and it is less necessary for the methods to anticipate the future.
 Impact of the choice of base classifier
 All methods have been compared using the same classifier: MiniROCKET so that only the decision components differ.
-However, the choice of the base classifiers could induce a bias favoring or hampering some methods.
-In order to clarify this, we have repeated the experiments replacing MiniROCKET with two base classifiers: WEASEL 2.0 (?), and the XGBoost classifier (?) using features produced by tsfresh (?).
-Both of these classifiers have already been tested within the ECTS literature by ?
-(?, ?) and ?
-(?) respectively.
-Figure 12 and 13 in Appendix B report the results respectively with these two classification methods.
+However, the choice of the base classifier could induce a bias favoring or hampering some methods.
+In order to clarify this, we have repeated the experiments replacing MiniROCKET with two base classifiers: WEASEL 2.0 (Schäfer & Leser, 2023), and the XGBoost classifier (Chen et al., 2015) using features produced by tsfresh (Christ et al., 2018).
+Both of these classifiers have already been tested within the ECTS literature by Schäfer & Leser (2020); Lv et al.
+(2019) and Achenchabe et al.
+(2021) respectively.
+Figure LABEL:fig:weasel and LABEL:fig:xgb in Appendix LABEL:change_clf report the results respectively with these two classification methods.
 One can observe that the results are not significantly altered with the same overall ordering of the methods when varying the value of $\alpha$.
-Furthermore, our results on AvgCost show that performances tends to be better for all methods using MiniROCKET.
-It is thus to be preferred given its simplicity and good performances.
+Furthermore, our results on AvgCost show that performance tends to be better for all methods using MiniROCKET (see Table LABEL:tab:clf_perfs in Appendix LABEL:change_clf).
+It is thus to be preferred given its simplicity and good performance.
 Impact of z-normalization
 Considering the newly proposed ensemble of datasets, we were not able to identify any problems of information leakage over time.
-This inconclusive result simply indicates that the variance of the time series measurements is not informative for theses datasets, which still could be the case considering past published results.
-For further details, please refer to Section B.5 of Appendix B.
+This inconclusive result simply indicates that the variance of the time series measurements is not informative for these datasets, which still could be the case considering past published results.
+For further details, please refer to Appendix LABEL:z_norm_app.
 5 Conclusion
-The first contribution of this research work is the coding of all the methods tested and making the codes available in a repository open to everyone.
-In this way, the experiments reported can be duplicated and further studies carried out.
-Furthermore, the deposited datasets and the experimental framework provide a ground for fair comparisons bewteen competing methods.
-We claim that the AvgCost is the appropriate measure by which to evaluate the performance of the methods.
-This is indeed what will be “paid” at the end of the day by a practitioner using a method.
-We have accordingly characterized a number of methods from the literature.
-Our extensive experiments have shown that:
-  • It is worthwhile to resort to dedicated ECTS methods, and more so in scenarios like anomaly detection with asymmetrical misclassification cost and exponential delay cost.
-  • However, it is noticeable that Proba Threshold, a baseline method, is surprisingly good overall in the standard setting with symmetrical misclassification cost and linear delay cost, exhibiting comparable performance to confidence-based myopic methods such as Stopping-Rule and anticipation-based cost-informed ones such as Calimera and Economy.
-  • Calibration of the classifiers has a large impact on some methods such as Calimera in particular, less so on other methods like Proba Threshold and Ecdire.
-In this paper, we have proposed a reading guide to highlight the main characteristics of ECTS methods, namely (i) the importance of the two components: decision and prediction which are distinct in the “separable” architecture and not in the “end-to-end” one, (ii) the distinction between anticipation-based and myopic methods, and (iii) between cost-informed and cost-uninformed techniques.
-On the basis of these dimensions, it becomes easy to imagine new methods that combine them in original ways, which could lead to new properties and better performance for solving the problem of early classification of time series, which, present in many applications has potentially great impacts.
-To go a step further, future work could be carried out to study the literature’s approaches applied in as yet unexplored cost settings.
+In this paper, we have proposed a taxonomy that allows to underline the main families of separable approaches for the ECTS problem, pointing out the essential components and the questions that have to be tackled when designing a new method.
+We have thus enlightened (i) the importance of the two components: decision and prediction, (ii) the distinction between anticipation-based and myopic methods, and (iii) between cost-informed and cost-uninformed techniques.
+We have defined a methodology for evaluating and comparing ECTS methods.
+In addition, and we hope this will prove useful to the community, we have built an open source library that includes systematic implementation of the methods tested, the proposed evaluation protocol, as well as a collection of 34 datasets designed to enable informative testing of the methods.
+We have also underlined the importance of considering a variety of cost settings in the evaluation of ECTS methods so as to reflect what can happen in real-life applications.
+The in-depth experiments carried out shed light on design choices.
+Firstly, are anticipation-based methods better than myopic ones?
+We have shown that four of the eleven methods tested perform significantly better.
+However, these methods do not share the same characteristics with regard to this question.
+Secondly, to the question: do cost-informed methods outperform cost-uninformed methods?
+The Pareto front indicates a clear superiority of the former.
+Finally, when we test the robustness of methods to variations in cost functions, anticipation-based methods prove superior and, among them, cost-based methods fare even better.
+Our experiments have also shown that calibration of the classifiers has a large impact on some methods (e.g.
+Calimera) in particular, less so on other methods (e.g.
+Proba Threshold and Ecdire).
+As regards to the z-normalization of time series which would be unwisely used in evaluation studies, our results (see Appendix LABEL:z_norm_app) show that the impact on the performance is limited, thus suggesting that the existing trigger functions do not or only slightly benefit from this information leakage.
+Future work could be carried out to study the literature’s approaches applied in as yet unexplored cost settings.
 For example, in many applications, the delay cost depends on the true class and the predicted one, and thus a single cost function integrating misclassification and delay costs should then be used.
-The use of this general cost form requires the adaptation of some state-of-the-art methods and has not yet been studied.
+This general cost form requires the adaptation of some state-of-the-art methods and has not yet been studied.
 In addition, in real ECTS applications, it is up to the business expert to define the costs, which is not an easy task in practice.
-Future work could study the impact of: (i) noisy costs, (ii) and cost drift between training and testing stages.
-This would make it possible to identify the most resilient literature approaches.
+Among the challenges, applications where the costs actually paid are not deterministic are of key interest (e.g.
+a manufacturing defect on an engine part does not necessarily lead to a failure, but it does increase the probability of paying a higher cost).
+Thus, future work could study the impact of stochastic cost functions.
+Another interesting case is applications where the costs paid are slightly different, or changed, from those defined by the business experts for the training phase (e.g.
+a change in the price of raw materials).
+Those kinds of cost drift between training and testing stages could also be further studied.
 Finally, in the case of existing separable approaches, the misclassification cost is not exploited for training the classification function.
 Future work could investigate the interest of using cost-sensitive classifiers in the case of ECTS.
-A Data description
-A.1 UCR Time Series Classification datasets
+Beyond separable approaches, a more extensive review of end-to-end ECTS methods could be interesting to complete the one presented in this paper.
+In particular, an experimental protocol should be designed to place end-to-end and separable methods on similar ground and to design new, efficient approaches.
+Appendix A Benchmarking: related studies
+Table 3: Table of published methods for the ECTS problem with their benchmarking protocol.
+References No. of datasets No. of methods Stats test Eval. metric(s) Cost balance Code
+Reject (Hatami & Chira, 2013) 1 1 Acc./Earl. fixed
+iHMM (Antonucci et al., 2015) 1 2 $\ell_{0/1}$ fixed
+ECDIRE (Mori et al., 2017b) 45 5 Acc./Earl. fixed
+Stopping Rule (Mori et al., 2017a) 45 5 Acc./Earl. vary
+ECEC (Lv et al., 2019) 45+8 6 Acc./Earl. vary
+TEASER (Schäfer & Leser, 2020) 45 5 Acc./Earl./HM vary
+SOCN (Lv et al., 2023) 45 10 Acc./Earl./HM fixed
+ECTS (Xing et al., 2012) 23 3 Acc./Earl. fixed
+RelClass (Parrish et al., 2013) 15 2 Acc./Earl. vary
+2step/NoCluster (Tavenard & Malinowski, 2016) 76 3 AvgCost/Acc. vary
+ECONOMY-$\gamma$-max (Zafar et al., 2021) 45/34 2 AvgCost vary
+CALIMERA (Bilski & Jastrzebska, 2023) 45/34 8 AvgCost/Acc./Earl./HM fixed
+FIRMBOUND (Ebihara et al., 2025) 7 5 AAPR/SAT vary
+EDSC (Xing et al., 2011) 7 4 Acc./Earl. fixed
+EARLIEST (Hartvigsen et al., 2019) 4 3 Acc./Earl. vary
+DDQN (Martinez et al., 2020) 1 2 Acc./Earl. vary
+DETSCNet (Chen et al., 2022) 12 5 HM fixed
+Benefitter (Shekhar et al., 2023) 14 7 Acc./Earl. vary
+CIS (Cao et al., 2023) 3 3 Acc./Earl. fixed
+ELECTS (Rußwurm et al., 2023) 4 2 Acc./Earl. vary
+EarlyStop-RL (Wang et al., 2024) 1 4 FPR/FNR/Earl./F1 fixed
+Empirical Survey (Akasiadis et al., 2024) 3 5 Acc./Earl./HM/F1 fixed
+Columns are defined as :
+  • No. of datasets: the number of datasets used in the paper.
+  • No. of methods: the number of methods tested in the paper (including the one proposed, if any).
+  • Stats test: whether evaluation has been performed using statistical tests (Rainio et al., 2024).
+  • Eval. metric(s): evaluation metrics used.
+  • Cost balance: whether experiments have been conducted for different cost balances between $C_{m}$ and $C_{d}$.
+  • Code: is the code of the method available.
+Experimental methodology is heterogeneous in the literature so far, making comparisons between methods more difficult.
+When considering datasets used or evaluation metrics, a variety of different approaches can be found, as shown in Table 3.
+Appendix B Experimental protocol
+B.1 Optimization of the parameters of the methods
+[…]
+B.2 Hyperparameters
+Table 4: Hyperparameters’ value. $\Delta$ defines grid’s size when performing grid-search over continuous valued intervals.
+Hyperparameters
+Method Fixed Optimized
+Calimera kernel: "rbf"
+[…]
+Ects support = 0
+[…]
+Proba Threshold $\Delta=40$
+Stopping Rule $\gamma_{1},\>\gamma_{2},\>\gamma_{3}\in[-1,1]$
+$\Delta=10^{3}$
+[…]
+B.3 Calibration of the classifications
+Like Bilski & Jastrzebska (2023), we add a calibration step when learning the classifiers, i.e.
+Platt’s scaling (Platt et al., 1999).
+Indeed, as we are dealing with collections of independently trained classifiers, the prediction scores may not remain consistent with one another over the time dimension.
+However, the trigger methods usually have their parameters set with the same values for all time steps.
+This is the case, for example with the Proba Threshold approach.
+In addition, some approaches such as Calimera and Economy-$\gamma$-Max exploit the estimated posterior probabilities $\{p(y|\textbf{x}_{t})\}_{y\in\mathcal{Y}}$ to estimate the future cost expectation.
+It is therefore highly desirable for all classifiers, at all times, to have their output calibrated and is necessary for a fair comparison.
+B.4 Datasets selection
+When a time series does not bring an increasing level of information over time about its class, classifiers are likely to obtain the same confusion matrix for all time steps, and the trigger function should then choose to make a prediction at the first time step to avoid the delay cost even if the prediction is very uncertain.
+Although this case may occur, it cannot form the basis for comparing the trigger functions.
+This is why we have been careful about the selection of data sets.
+We have chosen data sets where the classification performance tends to increase over time, as measured with the same basis classifier for all methods (see Section 4.1.2).
+We check that the train AUC increases averaged over a time window of size 5, when using either the beginning of the series, or almost complete ones.
+Specifically, the beginning of the series is taken using $5\%,\ldots,25\%$, and the almost complete series are of the following lengths $75\%,\ldots,100\%$.
+When the difference of averaged AUC is strictly positive, the dataset is selected.
+(see Table LABEL:tab:orange_cylinder).
+In addition, all datasets exhibiting z-normalization have been removed.
+Thirty-four datasets passed these demanding criteria (see Appendix E).
+For instance, Figure 8 (left) displays the time series of a synthetic data set where each of the three classes is characterized by a specific pattern over time.
+Class 0 can thus be recognized early on, while class 1 and class 2 can only be discriminated after approximately one third of the length of the time series.
+This is confirmed by the accuracy of the recognition of each class (right).
+Selecting data sets appropriate for the comparison of ECTS methods is a demanding process that must check for all the desirable properties described above.
+We thus hope that the corpus thus built and made available on the open source library will be useful to the scientific community for future performance studies.
+Figure 8: Examples of training time series of the SmoothSubspace dataset.
+Each of the class has a discriminative signal within one third of the serie (see Appendix D for a code snippet reproducing these results, based on the proposed library).
+Appendix C End-to-end ECTS
+C.1 State-of-the-art
+Table 5: Table of published end-to-end methods for the ECTS problem.
+References Architecture RL DL Cost-genericity
+EDSC (Xing et al., 2011) Shapelet
+EARLIEST (Hartvigsen et al., 2019) LSTM
+DDQN (Martinez et al., 2020) MLP
+DETSCNet (Chen et al., 2022) TCN
+Benefitter (Shekhar et al., 2023) LSTM
+CIS (Cao et al., 2023) LSTM
+ELECTS (Rußwurm et al., 2023) LSTM
+EarlyStop-RL (Wang et al., 2024) MLP
+$\bullet$ A different class of methods relies on searching telltale representations of subsequences, such that if the incoming time sequence $\mathbf{x}_{t}$ matches one or more of these representations, then its class can be predicted.
+Typically, these representations take the form of shapelets that discriminate well one class from the others (Ye & Keogh, 2011).
+For instance, the Early Distinctive Shapelet Classification (Edsc) method learns a distance threshold for each shapelet, based on the computation of the Euclidean distance between the considered subsequence and all other valid subsequences in the training set (Xing et al., 2011).
+It selects a subset of them, based on a utility measure that combines precision and recall, weighted by the earliness.
+A prediction is made as soon as $\mathbf{x}_{t}$ matches one of these shapelets well enough.
+Because this family of methods is computationally expensive, extensions have been developed to reduce the computational load (Yan et al., 2020; Zhang & Wan, 2022).
+Other extensions aim at improving the reliability of the predictions (Ghalwash et al., 2014; Yao et al., 2019), and tackling multivariate time series (Ghalwash & Obradovic, 2012; He et al., 2013; 2015; Lin et al., 2015).
+$\bullet$ The Earliest (Early and Adaptive Recurrent Label ESTimator) uses a RNN architecture (Hartvigsen et al., 2019) to make the prediction and a Reinforcement Learning agent trained jointly using policy gradient to trigger prediction or not.
+If a prediction is triggered, the hidden representation given by the RNN is sent to a Discriminator, whose role is to predict a class, given this representation.
+The model has been adapted to deal with irregularly sampled time series (Hartvigsen et al., 2022).
+$\bullet$ Martinez et al.
+(2018; 2020) use a Deep Q-Network (Mnih et al., 2015), alongside a specifically designed reward signal, encouraging the agent to find a good trade-off between earliness and accuracy.
+Those types of approaches also naturally extend to online settings where time series are not of fixed length.
+$\bullet$ The Decouple ETSC Network (Detscnet) (Chen et al., 2022) architecture leverages a gradient projection technique in order to jointly learn two sub-modules: one for variable-length series classification, and the other for the early exiting task.
+$\bullet$ The Benefitter algorithm (Shekhar et al., 2023) is an anticipation-based approach that learns to predict the benefit of triggering a prediction early.
+This quantity is equal to the saving one could make by triggering some decision now minus the cost induced by a wrong prediction.
+A LSTM model is learned to regress the benefit, which thus triggers, at inference time, as soon as the benefit is positive, i.e.
+when savings induced by temporal costs exceed estimated misclassification costs.
+$\bullet$ The Classifier-Induced Stopping (CIS) (Cao et al., 2023) model leverages policy gradient Reinforcement Learning in order to directly predict the optimal stopping time in a supervised way, found a posteriori for the training time series.
+$\bullet$ The End-to-end Learned Early Classification of Time Series method (Elects) leverages an LSTM architecture, adding a stopping prediction head to the network and adapting the loss function to promote good early predictions (Rußwurm et al., 2023).
+$\bullet$ Wang et al. (2024) introduce EarlyStop-RL, in which model-free RL is used to address the problem of early diagnosis of lung cancer.
+C.2 End-to-end experiments
+In this section, we test out two popular ECTS end-to-end methods, i.e.
+EARLIEST (Hartvigsen et al., 2019) and ELECTS (Rußwurm et al., 2023), and directly compare them to the separable baseline Proba Threshold.
+Due to the fact that those method do not directly expand to generic cost settings, we only perform this experiment using the standard cost setting, as described in Section 4.2.
+Figure 9: Standard cost setting, non z-normalized proposed datasets (orange cylinder).
+It is found that end-to-end approaches outperform the separable baseline Proba Threshold for high temporal costs.
+Since end-to-end approaches are not constrained to consider time series only over the 20 timestamps corresponding to every 5% of the time series duration for prediction, they can make their prediction in the interval.
+For instance, before 5% of the time series has been observed, which happens for high delay cost, whereas separable methods, in this experimental setting, must wait until the first time step, resulting in higher average costs.
+However, as the delay cost decreases, end-to-end methods tend to be less stable overall, falling behind simple separable base method, such as Proba Threshold.
+Appendix D Library
+More details are given about the proposed library here.
+It leverages the modularity of separable ECTS approaches and offers the possibility to easily assess each of the component contribution to final performance.
+The main library object is an EarlyClassifier object that needs, at least, 3 inputs:
+  • a ChronologicalClassifier that outputs classification prediction based on varying-size time series,
+  • a TriggerModel that decides to whether accept or not the current classifier’s prediction,
+  • a CostMatrices that defines the cost functions to be optimized.
+As used throughout the paper, one way to implement a classification strategy is to use an ensemble of classifiers, each of which is specialized for a particular timestamp.
+The ClassifiersCollection object offers an interface to do that.
+It takes as input any scikit-learn estimator as well as a set of timestamps for which to learn a classifier.
+For example, the following code uses the same classification module used to produce Figure 8.
+from sklearn.linear_model import RidgeClassifierCV
+from sklearn.preprocessing import SplineTransformer
+from sklearn.calibration import CalibratedClassifierCV
+from sklearn.pipeline import make_pipeline
+from ml_edm.classification.classifiers_collection import ClassifiersCollection
+T # time series length
+valid_timestamps = list(range(1, T+1)
+clf = make_pipeline(
+SplineTransformer(),
+CalibratedClassifierCV(RidgeClassifierCV(), method="sigmoid")
+collection_clf = ClassifiersCollection(
+base_classifier=clf,
+timestamps=valid_timestamps)
+Then, having defined a classification strategy, one can fit a full ECTS model, for example, using the ProbabilityThreshold trigger model.
+By default, when no other argument than timestamp is given to the CostMatrices object, it uses the default cost setting, as defined in Section 4.2.
+from ml_edm.early_classifier import EarlyClassifier
+from ml_edm.trigger import ProbabilityThreshold
+from ml_edm.cost_matrices import CostMatrices
+early_clf = EarlyClassifier(
+chronological_classifiers=collection_clf,
+trigger_model=ProbabilityThreshold(valid_timestamps),
+cost_matrices=CostMatrices(valid_timestamps),
+early_clf.fit(X, y)
+Thus, the library allows to easily pursue research in the separable ECTS field, facilitating the implementation of new approaches (whether in the classification or trigger part) and the systematic comparison with previous ones.
+It also pave the way for end-to-end approaches that can be framed as a trigger model that does not rely on any classifier’s outputs.
+Appendix E Data description
+E.1 UCR Time Series Classification datasets
 Table 6: UCR TSC datasets : 77 datasets from the UCR archive have been retained to run the experiments over the 128 contained in the full archive.
@@
 Yoga 300 3000 426 2 Image
-A.2 Proposed, non z-normalized, datasets
-Train Test Length Class Type
-Data
-[…]
-Chinatown 20 345 24 2 Traffic
-Crop 7200 16800 46 24 Image
-DodgerLoopDay 78 80 288 7 Sensor
-EOGVerticalSignal 362 362 1250 12 EOG
-GestureMidAirD1 208 130 360 26 Trajectory
-GunPointAgeSpan 135 316 150 2 Motion
-HouseTwenty 34 101 3000 2 Device
-InsectEPGRegularTrain 62 249 601 3 EPG
-MelbournePedestrian 1200 2450 24 10 Traffic
-PLAID 537 537 Vary 11 Device
-Rock 20 50 2844 4 Spectrum
-SemgHandGenderCh2 300 600 1500 2 Spectrum
-SmoothSubspace 150 150 15 3 Simulated
-UMD 36 144 150 3 Simulated
-[…]
-AluminiumConcentration 440 189 2542 2 Environnement
-BitcoinSentiment 232 100 24 2 Sentiment
-ChilledWaterPredictor 321 138 168 2 Energie
-CopperConcentration 440 189 2542 2 Environnement
-Covid19Andalusia 142 62 91 2 Santé
-DailyOilGasPrices 133 58 30 2 Économie
-DhakaHourlyAirQuality 1447 621 24 2 Environnement
-ElectricityPredictor 567 243 168 2 Energie
-FloodModeling3 429 184 266 2 Environnement
-HouseholdPowerConsumption1 1001 430 1440 2 Energie
-HotwaterPredictor 245 106 168 2 Energie
-MadridPM10Quality_nmv 4845 2077 168 2 Environnement
-ParkingBirmingham_eq 1391 597 14 2 Environnement
-PrecipitationAndalusia_nmv 470 202 365 2 Environnement
-SierraNevadaMountainsSnow 350 150 30 2 Environnement
-SolarRadiationAndalusia_nmv 470 202 365 2 Energie
-SteamPredictor 210 90 168 2 Energie
-TetuanEnergyConsumption 254 110 144 2 Energie
-WindTurbinePower 596 256 144 2 Energie
-Table 3: New datasets collection : 35 datasets from both the UCR archive (dashed line) and the Monash UEA extrinsic regression archive.
-When missing values and/or varying lengths, replace missing values with 0 and pad series to maximum length with 0.
-All of the datasets are not $z$-normalized and thus don’t suffer from potential data leakage.
-Italic datasets are not included when classes are imbalanced as problems become too difficult for the chosen classifiers.
-B Supplementary results
-B.1 Additional figures : Standard cost setting
-Figure 8: Standard cost setting, non $z$-normalized proposed datasets (orange cylinder).
-Compared to Figure 3(a), the global ranking is not altered much.
-One can observe that for $\alpha\in[0.5,0.7]$ the top group is now more populated, gathering the first six approaches, probably due to the limited amount of datasets available in this case.
-B.2 Additional figures : Anomaly detection cost setting
-Figure 9: Anomaly detection cost setting, original UCR datasets (blue cylinder).
-Compared to Figure 6(a), one can see that Calimera is now clearly dominating all other methods for all $\alpha$.
-The global ranking remains globally stable otherwise.
-B.3 Removing calibration
-Figure 10: Standard cost setting, original UCR datasets (blue cylinder).
-The calibration step is now removed, i.e.
-the outputs from the decision function is now simply passed through a softmax function.
-Both Calimera and Proba Threshold suffer heavily from using uncalibrated scores.
-Figure 11: Pairwise comparison (?), calibration ($calib$) vs no calibration ($no\>\>calib$), $\alpha=0.8$.
-Square colors are indexed on the mean AvgCost difference.
-For example, Calimera has a lower mean AvgCost when trained over calibrated scores: it appears in dark blue.
-The Wilcoxon p-value is equal to 0.0288, which is lower than significance level equal to 0.05.
-Thus, Calimera statistically under-performs when using uncalibrated scores.
-This is also the case for the Proba Threshold method.
-B.4 Changing the base classifier
-Figure 12: Standard cost setting, original UCR datasets (blue cylinder). The base classifier is now Weasel 2.0 (?). Results are very close to those exposed in Figure 3(a).
-Figure 13: Standard cost setting, original UCR datasets (blue cylinder).
-The base classifier is now a pipeline including features extraction with tsfresh (?) and classification using XGBoost (?).
-Results are a bit noisier than those exposed in Figure 3(a).
-B.5 Impact of z-normalization
+E.2 Proposed, non z-normalized, datasets
+Size Length Class Type AUC Gain AUC Gain
+train test
+Data (half/full) (half/full)
+BME 180 128 3 Simulated (7%/7%) (11%/16%)
+Chinatown 365 24 2 Traffic (1%/1%) (0%/0%)
+Crop 24000 46 24 Image (9%/10%) (8%/9%)
+DodgerLoopDay 158 288 7 Sensor (4%/5%) (14%/17%)
+EOGVerticalSignal 724 1250 12 EOG (35%/35%) (43%/45%)
+GestureMidAirD1 338 360 26 Trajectory (11%/12%) (23%/26%)
+GunPointAgeSpan 451 150 2 Motion (7%/7%) (7%/7%)
+HouseTwenty 135 3000 2 Device (1%/1%) (5%/6%)
+MelbournePedestrian 3650 24 10 Traffic (15%/15%) (15%/16%)
+PLAID 1074 Vary 11 Device (0%/0%) (2%/2%)
+Rock 70 2844 4 Spectrum (1%/1%) (12%/12%)
+SemgHandGenderCh2 900 1500 2 Spectrum (1%/2%) (11%/13%)
+SmoothSubspace 300 15 3 Simulated (21%/37%) (20%/38%)
+UMD 180 150 3 Simulated (10%/11%) (22%/28%)
+AcousticContaminationMadrid 138 365 2 Environment (1%/2%) (5%/7%)
+AluminiumConcentration 629 2542 2 Environment (3%/4%) (5%/10%)
+BitcoinSentiment 332 24 2 Sentiment (11%/13%) (-5%/-6%)
+ChilledWaterPredictor 459 168 2 Energy (0%/0%) (8%/11%)
+CopperConcentration 629 2542 2 Environment (4%/5%) (4%/3%)
+Covid19Andalusia 204 91 2 Health (7%/8%) (7%/17%)
+DailyOilGasPrices 188 30 2 Economy (25%/23%) (10%/8%)
+DhakaHourlyAirQuality 2068 24 2 Environment (2%/3%) (0%/2%)
+ElectricityPredictor 810 168 2 Energy (4%/5%) (7%/13%)
+FloodModeling3 613 266 2 Environment (20%/22%) (20%/31%)
+HouseholdPowerConsumption1 1431 1440 2 Energy (4%/7%) (14%/27%)
+HotwaterPredictor 351 168 2 Energy (1%/1%) (4%/5%)
+MadridPM10Quality 6923 168 2 Environment (4%/5%) (11%/16%)
+ParkingBirmingham 1888 14 2 Environment (7%/28%) (5%/22%)
+PrecipitationAndalusia 672 365 2 Environment (0%/1%) (3%/4%)
+SierraNevadaMountainsSnow 500 30 2 Environment (6%/7%) (-2%/4%)
+SolarRadiationAndalusia 672 365 2 Energy (1%/1%) (4%/4%)
+SteamPredictor 300 168 2 Energy (2%/2%) (3%/6%)
+TetuanEnergyConsumption 364 144 2 Energy (5%/5%) (1%/6%)
+WindTurbinePower 852 144 2 Energy (11%/12%) (15%/21%)
+Appendix F Supplementary results
+F.1 Additional figures : Standard cost setting
+F.2 Additional figures : Anomaly detection cost setting
+F.3 Removing calibration
+F.4 Ablation study on non-myopic trigger models
+One of the key component when it comes to trigger functions is the non-myopic ability to anticipate the likely future.
+In this section, we conduct an ablation study over the anticipation-based property by making those type of approaches myopic.
+In particular, Economy and Calimera will be of interest here.
+More precisely, the expected cost horizon has been limit to 1 for Economy.
+Concerning Calimera, the backward cost propagation has been blocked such that the only available information to trigger be the expected cost difference between current and next timestamp.
+F.5 Changing the base classifier
+classifier
+$\alpha$ MiniROCKET Weasel 2.0 tsfresh&XGBoost
+0 12.60% 11.95% 16.36%
+0.1 31.56% 18.44% 37.01%
+0.2 32.86% 18.96% 36.49%
+0.3 36.49% 19.48% 32.34%
+0.4 39.87% 17.53% 31.30%
+0.5 43.12% 19.10% 26.62%
+0.6 42.60% 23.25% 23.51%
+0.7 44.16% 24.94% 20.52%
+0.8 48.44% 25.45% 15.71%
+0.9 48.70% 26.62% 14.29%
+1 42.21% 26.75% 14.68%
+F.6 Impact of z-normalization
 Clearly, using z-normalized datasets is not applicable in practice, as it would require knowledge of the entire incoming time series.
@@
 For example, when a normalized time series has a low variance at the beginning, we can expect a high variance in the rest of the series since the mean variance is 1.
-There is therefore an information leakage that can be exploited by a ECTS algorithm, while this is not representative of what happens in real applications.
-A proposal such as the one presented in (?), where the z-normalization of available time series is repeated at each time step, has its own problems.
+There is therefore an information leakage that can be exploited by an ECTS algorithm, while this is not representative of what happens in real applications.
+A proposal such as the one presented by Schäfer & Leser (2020), where the z-normalization of available time series is repeated at each time step, has its own problems.
 In particular, it means that if a single classifier is used for all time steps, the representation of $\mathbf{x}_{t}$ can be different at times $t$ and $t+1$ and all further time steps which can induce confusion for the classifier.
 On the one hand, z-normalization induces an information leakage that could help methods to unduly exploit knowledge about the future of incoming time series.
-On the other hand, any normalization rescale the signal and therefore, potentially, hinder the recognition of telltale features.
+On the other hand, any normalization rescales the signal and therefore, potentially, hinder the recognition of telltale features.
 So, does z-normalization affect the performance of ECTS methods?
@@
 We duplicate and z-normalized them to get a second collection.
-As explained in Section 4.1, some extrinsic regression datasets has been converted into classification one.
-Here, the threshold value chosen to discretize the output into binary classes has been set the median of the regression target.
+As explained in Section 4.1, some extrinsic regression datasets have been converted into classification ones.
+Here, the threshold value chosen to discretize the output into binary classes has been set to the median of the regression target.
 In this way, classes within those datasets are equally populated.
 In these experiments, the delay cost is linear as in Section 4.2 and as in most of the literature.
-Figure 14 reports pairwise comparisons done on the 35 datasets.
+Figure LABEL:fig:heat_bar reports pairwise comparisons done on the 35 datasets.
 We look at $\alpha=0.8$, as this is the only value for which significant differences are observed.
@@
 Ecdire, actually has a better mean AvgCost when being trained on z-normalized data.
-Regarding the remaining methods, both Stopping Rule and TeaserAvg performs significantly worse when operating on z-normalized data.
+Regarding the remaining methods, both Stopping Rule and TeaserAvg perform significantly worse when operating on z-normalized data.
 Those trends are quite similar for other $\alpha$ values, without any significance on the statistical tests though.
 Thus, while z-normalization has some impact, since privileged information from the future can be leaked, our experiments, for the proposed datasets collection at least, show that this does not alter the overall results reported in the literature, and are globally in accordance with the results presented in Section 4.
-Figure 14: Pairwise comparison (?), $z$-normalization ($Z$) vs no $z$-normalization ($\bar{Z}$), $\alpha=0.8$.
-Square colors are indexed on the mean AvgCost difference.
-For example, Calimera has a lower mean AvgCost when trained over non $z$-normalized datasets by 1.3e-2 and appears in light blue.
-It beats the $z$-normalized version over 25 datasets, loses over 9 and are tied on 1.
-The Wilcoxon p-value is equal to 0.2224, which is higher than significance level equal to 0.05.
-Thus, no statistical difference can be observed for the considered approach.
-Figure 15: Multi-comparison-matrices (?).
-The upper triangle, with dark blue contours, displays the comparison of the competitive methods trained over non z-normalized dataset (orange cylinder).
-The values within this triangle has to be read by lines, i.e.
-for a considered line, red shades indicate better performances, blue shades weaker performances.
-The lower triangle, with dark red contours, is the comparison of the methods trained over the same datasets, z-normalized (chocolate cylinder).
-The values within this triangle has to be read by columns, i.e.
-for a considered column, red shades indicate better performances, blue shades weaker performances.
-The complete figure being symmetrical indicates that z-normalization does not impact much relative ranking between methods.
-B.6 Anomaly detection cost setting : an ablation study
+F.7 Anomaly detection cost setting : an ablation study
 Exponential delay cost only
-Figure 16: Exponential delay cost, symmetric binary misclassification cost, non z-normalized proposed imbalanced datasets (orange cylinder with a whole).
 Imbalanced misclassification cost only
-Figure 17: Linear delay cost, non symmetric imbalanced misclassification cost, non z-normalized proposed imbalanced datasets (orange cylinder with a whole).
 Standard cost setting, imbalanced datasets
-Figure 18: Linear delay cost, symmetric binary misclassification cost, non z-normalized proposed imbalanced datasets (orange cylinder with a whole).